### v3

Abstract
We describe the design, implementation, and evaluation of a low-cost (approximately \$15 per person) and scalable program, called Challenges, aimed at aiding women in Poland transition to technology-sector jobs. This program helps participants develop portfolios demonstrating job-relevant competencies. We conduct two independent evaluations, one of the Challenges program and the other of a traditional mentoring program – Mentoring – where experienced tech professionals work individually with mentees to support them in their job search. Exploiting the fact that both programs were oversubscribed, we randomized admissions and measured their impact on the probability of finding a job in the technology sector. We estimate that Mentoring increases the probability of finding a technology job within four months from 29% to 42% and Challenges from 20% to 29%, and the treatment effects do not attenuate over 12 months. Since both programs are capacity constrained in practice (only 28% of applicants can be accommodated), we evaluate the effectiveness of several alternative prioritization rules based on applicant characteristics. We find that a policy that selects applicants based on their predicted treatment effects increases the average treatment effect across the two programs to 22 percentage points. We further analyze how alternative prioritization rules compare to the selection that mentors used. We find that mentors selected applicants who were more likely to get a tech job even without participating in the program, and the treatment effect for applicants with similar characteristics to those selected by mentors is about half of the effect attainable when participants are prioritized optimally.
1 Introduction
An important concern of labor policymakers is identifying and implementing effective programs to support workers transitioning into occupations and industries that offer attractive jobs. One sector that has experienced rapid growth in recent decades and offers many attractive career opportunities is the technology sector. However, not all societal groups have benefited equally from opportunities in this sector; in particular, women remain underrepresented. Despite a number of programs and initiatives aimed at supporting women in the technology sector, [^1] For example, the United Nations organizes an annual event Girls in ICT Day; The European Commission carries out an annual study: Women in Digital Scoreboard. There are various NGOs dedicated to increasing women’s participation in the technology sector; For instance, Women in Technology International is a community of 160,000 women in the technology sector; Girls Who Code operates a network of 10,000 clubs providing support for women in technology. Finally, many major technology companies have initiatives aimed at supporting women in technology, including Women in Tech from Salesforce, The Women at Microsoft, or Accenture’s Women in TS&A. women held only 30% of technology jobs worldwide in 2021 (United Nations, 2021). The shortage of empirical evidence evaluating the effectiveness of specific strategies to address this issue is one of the hurdles in designing a successful policy response.
In this paper, we ask whether interventions can have an impact on this problem, and if so, which workers benefit most. We conduct randomized evaluations of two distinct programs, both implemented in Poland. The first, Mentoring, involves one-to-one mentoring and is thus not easy to scale. We designed and introduced a new program, Challenges, to be scalable. In Challenges, women work in teams to create portfolios to show to employers. Second, we ask which women benefit most from each program. Using data from the two evaluations, we identify the distinct subgroups with the largest benefits. Third, we evaluate the effectiveness of alternative prioritization rules for admission to the program in the presence of capacity constraints, illustrating that there are large gains from the optimal prioritization. Fourth, we examine the implicit selection rule used by mentors, who each selected two candidates from a pool of applicants. We find that mentors do not select the candidates with highest estimated treatment effects, but rather focus more on candidates who would be likely to succeed even without the program.
The results from our randomized evaluation suggest that the Challenges program increases the likelihood of successfully transitioning to the technology sector by approximately 45% within four months of graduation. [^2] Throughout this paper, we consider technology jobs as all jobs in firms in the technology sector other than those in finance, regulatory, legal, accounting, or human resources or jobs outside of the technology sector that involve software development and testing, IT support, and data analytics. Specifically, all jobs in UX and front-end programming are considered technology jobs. Challenges helps participants develop portfolios demonstrating capability relevant to the jobs they seek. The program was designed in partnership with technology companies, delivered online, and relies on interactions between program participants, which makes it cheap to operate (approximately \$15 per person) and suitable for scaling up. The program is also transferable across domains of the technology sector and languages; the first edition, evaluated in this paper, had specializations in user experience design and front-end programming. The program’s second edition was offered in Ukrainian instead of in Polish and offered additional specializations in software testing, cloud computing, and user interface design.
We designed the Challenges program as a scalable alternative to traditional in-person mentoring. To develop it, we partnered with Dare IT, a Polish organization offering a non-profit mentoring program – Mentoring, since 2018. Dare IT’s mentoring program serves women who, despite having invested time into acquiring IT skills, still have not successfully obtained a technology job. There is a high demand for this program, as evidenced by the fact that it has been repeatedly oversubscribed. However, scaling it up to meet the demand faces several challenges. First, it is difficult to find enough mentors. Mentors operate only on a volunteer basis and tend to be women in mid-seniority technology jobs. Since few women are in these positions, recruiting more mentors is difficult. Second, the one-to-one format of Mentoring makes it operationally demanding. Ensuring satisfaction, remedying conflicts, and tracking everyone are complicated tasks that require regular supervision by Dare IT staff.
The development of Challenges proceeded in two steps. First, together with Dare IT, we set up a series of interviews with HR experts and hiring managers of Polish technology companies to gain their perspectives on ways to support women candidates. Each of our interviewees mentioned the shortage of practical experience or clear signals of practical skills in women applicants’ profiles. Second, along with Dare IT and industry partners, we used the insights from the interviews to create Challenges. The program’s core consists of assignments sent every second week that, if completed, result in a job-relevant skill signal that can be added to participants’ resumes or mentioned during an interview. The program is designed based on three principles: (i.) it is delivered online, which lowers the costs; (ii.) it primarily involves interactions between participants, addressing the shortage of mentors; and (iii.) it is closely based on recent business cases. To operationalize the last principle, industry partners were recruited to help design the assignments and provide domain experts to review and critique selected submissions. We launched the registration to the program in December 2021 with 300 available spots; the program became oversubscribed within four hours.
We conducted two concurrent randomized experiments to evaluate the effectiveness of Challenges and Mentoring. Both experiments exploit the fact that the two programs were oversubscribed, so we were able to randomize access and evaluate the impact of participation in a program on the probability of having a technology job. To evaluate Mentoring, we asked mentors to select two mentee candidates (instead of the usual one), and within these pairs, we randomized into treatment and control. In the case of Challenges, we randomly assigned access to 200 spots in the program. Four months after the end of each program, we collected data on the outcomes. We estimate that participation in Mentoring increased the probability of having a job in technology from 29% to 42%, while the Challenges program increased this probability from 20% to 28%. [^3] The difference in the outcomes of the two control groups is likely because, in Mentoring, mentors selected applicants which were later randomized into the control group. The control group of Challenges is a random draw from the pool of applicants. We document that the outcomes of non-selected applicants are substantially lower than those of the applicants selected by mentors and later randomized into the control group. We repeated the evaluation 12 months after the end of the two programs and found that their impact did not attenuate over time.
We find that the two programs lead to different patterns of treatment effect heterogeneity. Mentoring is particularly beneficial for women from smaller towns that do not have a STEM degree, are mothers, and have been active in the labor market for over five years. Working with mentors with managerial experience and a longer tenure in technology also leads to high treatment effects. In the case of Challenges, participants who have just entered the labor market without a STEM degree exhibit high treatment effects. Different patterns of heterogeneity in treatment effects combined with insights from our discussions with Dare IT, mentors, and industry experts suggest that the two programs support participants’ employability through different mechanisms. Mentoring can positively impact the chances of obtaining employment in the technology sector by reinforcing self-efficacy, improving interview skills, or expanding professional networks. In contrast, creating portfolio items in the Challenges program that constitute signals of specific skills relevant to the job can counter a pessimistic belief (e.g., based on stereotypes) held by hiring managers, recruiters, or applicants themselves.
The core motivation for creating the Challenges program was to extend Dare IT’s impact beyond what is possible via the relatively small Mentoring program. Although Challenges expanded the number of people who receive Dare IT services, the existing capacity constraints in Mentoring and Challenges present a significant hurdle, as the number of interested applicants greatly exceeds the number of available spots in each program. Furthermore, at the time this research was conducted, Dare IT was not using paid advertising to promote its services, so the applicants to Dare IT heard about the program through word-of-mouth recommendations or through an internet search, which underscores that the potential demand for Dare IT services is even higher. In the case of Mentoring, the constraint is caused by the limited number of mentors and the operational difficulties of scaling up the program. For the Challenges program, the number of spots has been intentionally limited to 300 to maintain appropriate participant support. [^4] The first 100 applicants were automatically granted a spot in Challenges and were removed from the analysis conducted in this paper. Even though there are plans to incrementally increase the number of available spots in future Challenges cohorts, Dare IT will continue to grapple with the decisions of which applicants to admit to the Challenges program and how to select participants for the Mentoring program in the foreseeable future.
In the presence of capacity constraints and substantial heterogeneity in treatment effects, admission rules prioritizing applicants with the highest benefits can potentially enhance Dare IT’s impact. Adapting the methodology of Athey and Wager (2021) and Zhou et al. (2023), we estimate the value of counterfactual assignment policies that maximize outcomes across all applicants, given the constraints on the number of spots available in each program. These policies are then evaluated in the test set using doubly-robust estimates obtained through cross-fitting.
We estimate that the existing policy, which assigns 15% of all applicants to Challenges and 13% in Mentoring, leads to the weighted average treatment effect of 12 percentage points increase in the probability of getting a tech job. Considering the case in which the available number of spots in each program does not change, but we change the assignment policy to prioritizes applicants with the highest predicted benefits from each program (we refer to this policy as optimal policy), we show that the average treatment effect increases to 22 percentage points. Thus, we find a 10 percentage points incremental benefit of targeting. Second, we evaluate the optimal policy at various capacity levels. While expanding capacity increases the average outcomes as applicants who previously did not participate are admitted, the incremental value of targeting over random admissions might decrease, as applicants with the highest predicted benefits are allocated first, and the marginal participants have lower predicted benefits. However, we estimate that targeting has substantial value even at high capacity levels. We show that when the Challenges program is expanded to 50% of all applicants, the incremental benefit of targeting is 5 percentage points, while when there are no capacity limits, the incremental benefit of targeting is to increase the probability of finding a job by 3 percentage points. The comparison of random and targeted admissions at the full capacity level, underscores that some participants have higher benefits from one program than from the other; thus, the value of targeting is not only in targeting between programs and out of programs, but also across the two offerings.
Dare IT ensures high-quality matching by asking mentors to choose mentees from a list of pre-screened candidates. When comparing the outcomes of applicants to Mentoring, who were selected by mentors but randomized to the control group, with those not selected, we find that mentors tend to pick promising candidates. These candidates will likely find jobs in the technology sector, even without participating in the program. However, this approach to selecting participants might not be maximizing the potential impact of Mentoring. We use the off-policy approach to evaluate an alternative policy where mentors select candidates with the highest treatment effects. We find that the policy of choosing applicants based on their treatment effects leads to an average probability of getting a job in the technology sector of 23% across all applicants to Mentoring (we exclude applicants to Challenges from this analysis), compared to 20% when selecting the most promising candidates. This finding suggests that focusing on those most likely to benefit from the program rather than those who seem most promising overall could increase the program’s effectiveness.
Finally, a non-parametric assignment policy which requires solving a discrete optimization program, might not be realistic in the context of Dare IT. Therefore, to create viable alternatives which lead to similar benefits as optimal policies but are easier to implement, we propose simple rules that define priority groups based on observed characteristics (e.g., applicants from smaller cities are prioritized for Mentoring). We then evaluate these rules using predicted treatment effects. We find that prioritization can substantially increase the two programs’ effectiveness.
2 Literature review
This paper relates to several strands of literature. First, the paper contributes to the literature on the effectiveness of mentoring in improving career outcomes. Several influential papers focus on the effectiveness of mentoring in the academic context: Dennehy and Dasgupta (2017) employ a randomized experiment to show that assigning a female mentor to a female engineering student increases the sense of belonging in the discipline, self-efficacy, motivation, and retention for women majoring in engineering in college. Resnjanskij et al. (2021), also in a randomized experiment, document a mentoring program’s positive impact on disadvantaged high-school youth’s math grades and career plans. Several papers focus on the impact of mentoring on the careers of academic faculty members. Gardiner et al. (2007) provide empirical evidence that mentored individuals are more likely to stay at a university, receive more grant income, and receive promotions more frequently. In more recent studies, Ginther et al. (2020) and Blau et al. (2010) analyze the effectiveness of a mentorship workshop to support women in research careers. In a randomized access design similar to ours, they show that individuals in the treatment group are more likely to receive tenure at their institution, have more publications and citations, and receive more federal grants. We contribute to this literature by providing evidence from a randomized trial of the efficacy of mentoring outside of the context of education or academic jobs. [^5] The impact of mentoring on career outcomes has also been studied using surveys and other techniques based on observational data. For example, Ragins and Cotton (1999), in a survey of over 600 individuals, find that mentored respondents have better career outcomes than non-mentored respondents. They also find that the gender composition of the mentoring pair influences the effectiveness of mentoring. Similarly, Aryee et al. (1996) survey a group of workers in Hong Kong and found that mentored respondents reported a higher number of promotions and were more satisfied with their work than those not mentored. Kammeyer-Mueller and Judge (2008) carry out a meta-analysis of this literature and concludes that mentoring improves job and career satisfaction.
Second, there is rich evidence evaluating the efficacy of offline labor market training programs. Card et al. (2018) carry out a comprehensive meta-analysis of the estimates from over 200 recent studies. They find that on average, the analyzed training programs have close to zero short-term effects on employment and some small positive impact in 2-3 years after completion. Programs evaluated in Card et al. (2018) generally aim at unemployed individuals looking for blue-collar jobs. [^6] Other influential papers evaluating labor training programs have a similar focus: Barnow (1987); Bloom et al. (1997); Heckman et al. (1999). In this category, programs focused on the development of skills on-the-job tend to be more effective (Sianesi, 2008; Lechner and Gerfin, 2000). On-the-job training programs do not necessarily need to take the form of long apprenticeships to be effective: Biewen et al. (2014) compare a long apprenticeship with short-term on-the-job training, combined with a job search workshop, and find that the short-term program is equally effective or more effective than the much longer program. Fein and Hamadyk (2018) use a randomized experiment to evaluate YearUp, a sectoral training program, and find a 40% increase in quarterly earnings in the medium term. We contribute to this literature by bringing evidence from a randomized trial of a training program that happens online, is focused on individuals joining a highly paid technology sector, and is very cheap to operate. Our program has on-the-job training elements as participants develop content designed by domain experts to closely resemble the content they would produce on the job.
Third, a few studies have analyzed the efficacy of Massive Open Online Courses (MOOCs) in supporting learners’ job search. For instance, Zhenghao et al. (2015) shows that a significant majority of Coursera participants reported career benefits, particularly those from economically- and academically-disadvantaged backgrounds. However, through a difference-in-differences approach, Castaño-Muñoz and Rodrigues (2021) finds no effect on wages or employment among MOOC learners in Spain and a moderate impact on job retention. Hadavand et al. (2018) compare participants in a Data Science MOOC that were just above the threshold to graduate with those just below and argue that graduation from this MOOC leads to a 30% increase in the probability of finding a job. Our study contributes to this literature by using a randomized experiment approach. We also compare an online training program with an in-person Mentoring, identifying diverse patterns of heterogeneity that suggest different priority groups for online versus offline learning.
Fourth, we contribute to the literature on skill signaling (Spence, 1978). Previous research showed that credentials demonstrating capability can significantly boost earnings Tyler et al. (2000); Hadavand et al. (2018). Furthermore, in situations where companies might be reticent about hiring less experienced candidates, establishing trustworthy indicators of ability can bring advantages at the market level Pallais (2014). Our research, through a randomized trial, sheds light on the efficacy of a program aimed at developing labor market signals. Our subjects are women seeking tech roles who can utilize skill signaling to counter possible negative stereotypes. However, since the programs we evaluate incorporate both learning and signaling, we cannot separate the signaling effect.
Fifth, we document a high degree of heterogeneity in treatment effects from both programs, which we use to define treatment assignment policies. We show that targeting based on treatment effects leads to substantially higher average outcomes than random targeting, or targeting based on lowest outcomes in the absence of treatment, or highest outcomes with treatment. In doing so, we contribute to a small but growing literature on targeting treatment based on estimated treatment effects. An early contribution to this area was made by Ascarza (2018), who using randomized experiments, documents that treatment effect-based targeting is more effective in customer retention than churn-based prediction. Similar findings have been noted in marketing and development economics and medical contexts (Hitsch and Misra (2018), Devriendt et al. (2021), Haushofer et al. (2022), Inoue et al. (2023)). Lastly, a comparison of assignment policies for financial aid application nudges by Athey et al. (2023) highlights the superiority of maximizing treatment effects over targeting based on the predicted risk of non-compliance.
Finally, this paper adds to the literature on the gender gap in the tech sector. Various reasons for low women’s participation in the technology sector have been studied. For example, Murciano-Goroff (2022) identify women’s tendency to under-report programming skills, impacting their job prospects. Strategies proposed to counter this include increasing role model visibility (Correll and Mackenzie (2016)) and shifting the perception of tech roles being predominantly male (Cheryan et al. (2013), Del Carpio and Guadalupe (2022)). We contribute to this literature by demonstrating how mentoring and job-training programs can enhance women’s success rates in obtaining tech jobs. Notably, our work focuses on the primary outcome of finding a job.
3 Mentoring and Challenges Programs
Over recent decades, Poland’s technology sector has witnessed significant growth, culminating in a 2021 revenue of roughly USD 51 billion, which accounts for 8% of the national GDP and provides employment for 250,000 individuals (GUS, 2022a). The sector’s growth has contributed to an escalating demand for skilled tech professionals, subsequently driving an approximate 9% wage growth in 2021. This led to the sector’s median monthly wage reaching USD 2,600, substantially outpacing the country’s overall median monthly wage for full-time workers, which stands at approximately USD 950 (GUS, 2022b). A survey of gender demographics within this industry unveils an underrepresentation of women, who make up 30% of the total Polish tech workforce (Gawlowska-Bujok and Bujok, 2021). The European Commission (European Commission, 2021) reports that the tech industry accounts for 1.1% of Polish women active in the labor market, compared to 5.2% of men.
Dare IT is a social impact organization that aims at reducing the gender gap in the tech sector. They do so by helping qualified women transition from other careers. Dare IT offers a variety of services for women looking for jobs in tech or making their first steps in the profession, including a mentoring program, career fairs, networking events, and an online social network.
Mentoring.
Dare IT launched Mentoring in 2018 as a not-for-profit initiative to support women aspiring to tech jobs. Since then, the program has been offered twice per year, and the number of participants has increased in each cohort. The program is funded by grants and donations, which allows participation to be free of charge. In recent years, demand for the program has been far greater than supply. For example, for the spring 2021 cohort, Dare IT received over 1,000 applications for 150 spots.
The program follows several steps. Each cycle begins with Dare IT recruiting potential mentors through past participant referrals, social media engagement, and corporate partnerships. All mentors are women who are active professionals in the tech sector, and they contribute their time and expertise to the program on a pro-bono basis. Mentors are screened and selected to participate based on an application process; Dare IT considers multiple criteria, including candidates’ expertise in domains that attract the most interest from mentees, length of professional experience, and reputation.
Simultaneously, the recruitment of potential mentees occurs primarily via social media and word-of-mouth recommendations. During this study, Dare IT did not utilize paid advertisements for recruitment. Dare IT selects participants based on several criteria. Only fully-completed applications by women not currently engaged in a tech role are considered. The program primarily targets individuals with necessary skills — quantified by self-reported learning hours in a particular domain and manual review of applications by Dare IT employees — who require additional support in their tech job search. Applicants who do not meet a baseline skill level are excluded. After Dare IT finalizes the selection process, mentors receive access to a database of eligible applicants. They can then select their preferred mentee.
Once mentors and mentees are paired, they typically meet on a weekly basis for four months in order to support the mentee in their job search. Dare IT provides mentors with guidelines and best practices to increase the likelihood of a successful mentorship experience, although mentors have the flexibility to modify these guidelines to cater to their mentee’s needs. As a result, mentoring takes various forms, from coaching in job searching and interviewing or resume preparation to reviewing code and developing technical skills. Our discussions with mentors and Dare IT staff suggest four potential channels through which Mentoring increases the employability of mentees: (i.) reinforcement of self-efficacy through the provision of emotional support and encouragement, (ii.) help in the development of a professional network by relaying on connections of mentors [^7] Pallais and Sands (2016) emphasizes how referrals can augment job opportunities and delves into the underlying mechanisms., (iii.) explanation of unwritten norms of professional behavior and conduct, especially those intrinsic to the tech sector and, more specifically, during tech recruitment phases, and (iv.) direct transfer of skills through code review and discussions.
Challenges.
Due to the excess demand for Mentoring, Dare IT explored the possibility of expanding its capacity. However, scaling proved to be a complex issue. First, selecting more mentors without sacrificing the quality of mentoring is difficult due to the scarcity of women occupying suitable roles in tech. Second, implementing a large 1:1 mentoring program presents operational difficulties.
We partnered with Dare IT to develop a more easily scalable approach. Our first step was forming a set of hypotheses on barriers that impede women from securing roles in the Polish tech sector. We conducted interviews with hiring managers and HR experts from a dozen prominent tech companies in Poland. Each of our interviewees mentioned a shortage of practical experience or signals of practical skills on job application materials provided by female candidates. They emphasized that tangible proof of hands-on experience, even a short project, could significantly alter the interview dynamics, typically favoring candidates. They also expressed low trust in skill credentials that are available online. Many of our interviewees described online credentials as needing to be more informative about the capability to carry out real-world business tasks. Motivated by these findings, we developed a new program called Challenges, designed to address these specific concerns.
The Challenges program comprises six bi-weekly mini-projects. Successful completion of all mini-projects results in a portfolio item, e.g., a UX design for a mobile application, which functions as proof of practical experience in the chosen domain (see Figure 1 for an example of wireframes from a completed product). The program strongly emphasizes on-the-job training: assignments are curated by domain experts and are readily applicable in a business setting. Each weekly mini-project undergoes evaluation and feedback from domain practitioners, including existing mentors and industry professionals, facilitating ongoing skill and knowledge enhancement for participants. A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
Figure 1: Challenges final product.
Note: Selected wireframes from a design of a mobile app for pets adoption prepared during Challenges. See: https://www.behance.net/gallery/142047393/Pet-Adoption-App-UXUI for the entire design.
We launched the program in December 2021, setting the start date for the initial cohort in mid-January 2022. Dare IT promoted the program on its social media channels to encourage applications. Enrollment was capped at 300 participants. [^8] The decision to limit program capacity to 300 spots was made to account for potential logistical challenges and to ensure that we could provide personalized assistance to all participants if necessary. As the program progressed, it became evident that only a few participants required hands-on support. Consequently, in subsequent cohorts, Dare IT increased the number of available spots. The first 100 slots were allocated to the initial 100 applicants (excluded from this analysis), with the remaining 200 spots randomly assigned among subsequent applicants. The program was oversubscribed within four hours, at which point we closed the registration form. Applications received post-closure were not included in the randomization. However, we collected data on these late applicants’ outcomes for comparison.
4 Experimental design and data
4.1 Experimental design
Mentoring.
We evaluate the mentoring program using the cohort of Autumn 2021. As in previous cohorts, Dare IT received a significantly higher number of applications than the available number of mentors. Even after the screening process to verify eligibility, the ratio of applicants to mentors remained at 4:1. The eligibility for the Mentoring program underwent manual verification by Dare IT staff. Only applicants who did not currently hold positions in the technology field and who, according to Dare IT’s reviewer assessments, have technical skills at a level typical of a junior employee in a tech company were considered eligible for participation. The surplus of eligible applicants compared to the number of available mentors provided us with an opportunity to implement a randomized access experiment. In order to maintain high-quality mentor-mentee pairings while incorporating randomized access, we asked mentors to select two potential mentees instead of just one, as they did in the previous program’s cohorts. Within these pairs, we then randomly determined which of the two applicants would be granted access to the program (the treatment group) and which would not (the control group). Thus, we created a paired experiment.
In the experiment, we initially had 156 mentor-mentee pairs and an equivalent number of mentor-control pairs. However, four pairs did not start the program because the mentee secured a tech job between applying for the program and its commencement. In these cases, the mentor-selected applicant, who was initially randomized into the control group, participated in the program. [^9] We decided against having idle mentors to ensure broader access to the program, as well as, to respect mentors’ time dedicated to selecting candidates and willingness to volunteer their time in mentoring. Consequently, we excluded these mentors and participants from our analysis. As a result, our analysis included a total of 152 pairs. Additionally, we gathered data on 468 eligible applicants whom mentors did not select for participation.
Challenges.
Regarding the Challenges program, we concluded the recruitment process once we received 500 applications that met the established eligibility criteria. To qualify for participation, applicants were required to declare that they were not currently employed in the technology sector, possess a fluency level in English (B1 level), and be proficient in Figma or React depending on the specialization. The first 100 applicants were granted program spots, while the remaining 400 were randomly assigned to either the treatment or control groups. Randomization was conducted with stratification based on variables obtained from the registration survey. [^10] These stratification variables included specialization (front-end development or UX), age, type of residence, and employment status (student, employed, neither employed nor student). Additionally, we gathered data on 160 applicants who submitted their forms after the registration process had concluded. These late applicants were not included in the primary program evaluation. However, they are considered in the off-policy evaluation, as detailed in Section 6.
Furthermore, it’s worth noting that some applicants randomized into the treatment group did not actively participate in the program (did not submit any assignments). Nevertheless, since these individuals still received all program materials and had the opportunity to engage in webinars discussing projects, we have included all such individuals in our analysis.
4.2 Data
We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted following the program’s conclusion, and information collected from participants’ LinkedIn profiles.
Applicant registration survey.
Applicants in both programs were asked to complete a simple registration survey. Data from the survey included the city of residence, age, employment situation, and time spent, which measures how much time the applicant had spent developing their IT skills. The mentoring application form was richer; first, applicants were asked if they have family or friends in IT and whether they have children; second, the form also had several open-ended questions related to the applicant’s motivation and the possibility of attaching a short introductory video, but we do not use either in this analysis. Additionally, we collected information on participants’ preferred domains: the main categories were UX design, manager (project or product), or developer (front or back-end).
Mentors registration survey.
Mentors were asked a number of questions before the start of the program, which allowed us to construct the following variables: first-time mentor, career changer (which indicates whether the mentor herself had worked outside of IT before moving to the sector), managerial experience (which indicates whether the mentor has any managerial experience or not), and years of experience (years of experience in the tech sector).
Outcomes survey.
We surveyed participants of each program and those in control groups four months after the end of the program. In the survey, we were primarily interested in the salaries in the new jobs: we asked the question whether the respondent found a new job and in which range the salary was. We also inquired about the number of job offers and whether the participant negotiated the proposed offer or accepted the initial one. Finally, we carried out a shorter survey right after the end of Mentoring, the “short survey,” where we asked for salaries for the new jobs. Altogether we conducted three surveys: the main mentoring survey in which we received 162 responses (with 111 responded from treatment group subjects), the initial survey of the mentoring program (with 194 responses), and the Challenges survey, for which we received responses from 68 subjects in the control group and 61 from the treatment group. We analyze the outcomes from surveys in Appendix F.
LinkedIn data.
Our primary source of outcome data is the LinkedIn profiles of the participants. This data was collected at two different points: first, four months, and second, 12 months after the conclusion of the programs. During the initial data collection phase, outcomes for all experimental subjects were gathered. In the second phase, Dare IT was unable to observe the outcome of one Mentoring participant, 17 Challenges participants, and 7 Challenges control group members. The principal cause of the missing data is the change in the LinkedIn profile address. [^11] It’s plausible that the propensity to change LinkedIn’s profile address was impacted by participation in the program. In Appendix ate_bounds we provide robustness checks to our primary results by assuming the most pessimistic and optimistic outcomes for the missing subjects. We also estimate AT bounds following the methodology in Lee and Weidner (2021). The primary outcome metric was whether a participant secured a technology job following a program’s start. As new technology jobs, we consider jobs with a starting date of January 2022 or later for Mentoring and February 2022 or later for Challenges that fall into one of the following two categories:
  1. 1.
    All jobs in technology companies other than positions in finance, regulatory, legal, accounting, and HR, where technology companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online shops).
  2. 2.
    Jobs in non-technology companies that involve software development and testing, IT support, and data analytics. In our context, this category includes jobs in banks and management consulting agencies.
Based on these categories, we define three outcomes: new job, tech job 4 months, and tech job 12 months. New job refers to any job, whether in technology or not, that started between January and May 2022 for Mentoring and between February and June for Challenges. Tech job 4 months denotes a job that aligns with category (1) or (2) and is also classified as a new job. Finally, tech job 12 months is a job, corresponding to either category (1) or (2), that began after January 2022 for Mentoring and February for Challenges and before May 2023. We also collect the date on which subjects got their new jobs. In some cases, we do not observe the month in which the new job started, but only the year; for such observations, we assign the start date as December 2022 or May 2023, depending on whether the job started in 2022 or 2023.
We also leveraged LinkedIn to gather data on the highest education level attained by participants, which was organized into the following categories: bachelor’s degree, master’s degree, postgraduate degree, non-traditional credential, and other. The other category includes profiles lacking education information or information that could not be classified into any of the pre-existing categories. Furthermore, we compiled data on college majors, forming two binary variables: STEM and social science or business. The former includes all STEM degrees, excluding social sciences, while the latter comprises degrees in social science, business, finance, or management. Lastly, we deduced a person’s age from a combination of age reported in the registration survey and the commencement date of their latest educational degree. Finally, we also compute their professional experience length by subtracting the start date of their first job recorded on LinkedIn from 2022. Table 1 summarizes the key variables for both programs.
Table 1: Summary statistics of the main variables: treatment, control, and applicants
Variables Treated group Control group
N Mean St. Dev. Min Max N Mean St. Dev. Min Max
Mentoring: mentor characteristics
First time mentor 152 0.467 0.501 0 1 145 0.434 0.497 0 1
Managerial experience 152 0.520 0.501 0 1 145 0.510 0.502 0 1
Career changer 152 0.691 0.464 0 1 145 0.710 0.455 0. 1
Years of tech experience 152 5.856 3.777 2 35 145 6.021 3.794 2 35
Mentoring: experimental groups
New job 152 0.467 0.501 0 1 147 0.422 0.496 0 1
Tech job 4 months 152 0.421 0.495 0 1 147 0.293 0.456 0 1
Tech job 12 months 151 0.609 0.499 0 1 147 0.456 0.500 0 1
Social science 152 0.132 0.339 0 1 147 0.136 0.344 0 1
STEM 152 0.487 0.501 0 1 147 0.531 0.501 0 1
Family friends IT 152 0.697 0.461 0 1 147 0.701 0.460 0 1
Warsaw 152 0.276 0.449 0 1 147 0.259 0.439 0 1
Mother 152 0.257 0.438 0 1 147 0.259 0.439 0 1
Applied before 152 0.243 0.431 0 1 147 0.231 0.423 0 1
Years of professional experience 152 7.053 5.049 0 31 147 7.204 4.816 0 23
Challenges: experimental groups
New job 183 0.333 0.473 0 1 225 0.293 0.456 0 1
Tech job 4 months 183 0.284 0.452 0 1 225 0.196 0.398 0 1
Tech job 12 months 166 0.380 0.487 0 1 218 0.266 0.443 0 1
Social science 183 0.208 0.407 0 1 225 0.267 0.443 0 1
STEM 183 0.421 0.495 0 1 225 0.391 0.489 0 1
Years of professional experience 183 7.645 4.976 0 31 225 7.671 6.237 0 31
Applicants both programs
Tech job 4 months - - - - - 628 0.152 0.359 0 1
Tech job 12 months - - - - - 628 0.279 0.449 0 1
STEM - - - - - 628 0.196 0.367 0 1
Years of professional experience - - - - - 628 7.474 6.318 0 37
Note: Summary statistics of selected variables. The top panel characteristics of mentors. The second panel characteristics of treatment and control groups in the Mentoring experiment. The third panel characteristics of treatment and control groups in the Challenges experiment. The bottom panel, applicants to both programs that are not in experimental groups; this group includes subsets of applicants to Mentoring that were not selected by mentors and applicants to Challenges that submitted their application past registration closure. Career changer takes the value of one when the mentor reported in the registration survey having changed occupations. The variable years of tech experience is reported by mentors in the survey. STEM excludes social sciences and architecture. Social science includes business, management, finance, and accounting. The variable years of professional experience is measured based on LinkedIn profiles. Appendix B shows the covariate balance between treatment and control.
5 Effectiveness of Dare IT programs
Analytical framework
Let each Dare IT applicant be characterized by $(X_{i},Z_{i},W_{i}^{M},W_{i}^{C},S_{i},Y_{i})$, where $i$ is the applicant’s index $i=1,..,\mathcal{I}$, $X_{i}\in\mathcal{X}$ is a matrix of characteristics of applicants that are observed by mentors and recorded in data, $\zeta_{i}\in\mathcal{Z}$ are characteristic of applicants that are not observed in the data (though some might be observed by mentors prior to treatment assignment), and $S_{i}\in\mathcal{S}$ is a random variable indicating whether the applicant has been selected by a mentor. [^12] To select mentees, mentors received information from the application form; the important part of the application that mentors observed, which we did not, are short introductory videos recorded by the applicants. Finally, $W_{i}^{P}\in\left\{0,1\right\}$ is the treatment assignment into program $P$ and $Y_{i}\in\mathcal{Y}$ is the observed outcome.
Let $Y_{i}(P)$ and $Y_{i}(0)$ be potential outcomes when applicant $i$ participates in program $P$ and when she does not; $P\in\left\{M,C\right\}$, where $M$ stands for Mentoring and $C$ for Challenges. We are primarily interested in estimating the magnitude of the average treatment effect, defined as $\tau^{P}\coloneqq E\left[Y_{i}(P)-Y_{i}(0)\right]$, and testing whether average treatment effects are higher than zero. We are also interested in comparing the treatment outcomes to one another, although our estimates of such contrasts are noisy due to the limited sample size in each experiment.
Selected sample in Mentoring.
Dare IT asked mentors to choose two applicants they wanted to work with. Thus, selected applicants are not a random draw from the pool. Let $S_{i}$ be a random variable indicating whether a mentor selected the applicant $i$ or not; $S_{i}=1$ when the applicant has been selected by a mentor while $S_{i}=0$ otherwise. Only selected applicants were included in the randomization. Thus, we do not observe outcomes under treatment for the non-selected applicants. The average treatment effect of Mentoring that we estimate is defined as
$\tau^{M}\coloneqq E\left[Y_{i}(M)-Y_{i}(0)|S_{i}=1\right].$ (1)
Table 1 provides summary statistics of subjects in the experimental groups of the mentoring program and the non-selected applicants. There are notable differences between the mentoring control group (those selected by mentors ($S_{i}=1$) but not assigned to treatment ($W_{i}^{M}=0$) and the applicant group (where $S_{i}=0$ and thus, by construction, $W_{i}^{M}=0$), most notably in terms of the primary outcome of securing a tech job. In the Challenges program, participants were selected randomly from eligible applicants, where eligibility was determined based on observable characteristics.
5.1 Average treatment effects
In Figure 2, we display the proportions of participants employed in technology jobs over the span of 16 months, post-application to the program. This is illustrated across all experimental groups. [^13] We construct outcomes over time based on the new job’s start date. In some cases, we only observe the year the new job started rather than the specific month. For those observations, we assign December of that year as the starting month. Consequently, the average outcomes at the 4th month after the end of either program are lower than the average of tech job 4 months across the experimental groups. In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT). We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months. In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points.
The difference in average outcomes across the three groups appears to be substantial, particularly in the case of Mentoring. However, the subjects in the Mentoring group were selected by mentors; thus, the program’s impact on the probability of getting a tech job might be compounded by the influence of the non-random selection of the sample. In the right figure, we decompose the effect of selection from the impact of Mentoring. We show the outcomes across applicants selected by mentors and randomized into Mentoring, selected, but randomized into control, and those that were not selected (the non-selected group consists of the non-selected applicants to Mentoring, and the control group, and late applicants to Challenges). We observe that in the 16th month, the probability of having a job in technology differs by over 20 percentage points between selected and not-treated and non-selected applicants, and it is substantially larger than the difference between the two groups that were selected but randomized into treatment or control. On average, mentors selected candidates that were more likely to get a job in technology. This observation highlights that, in this context, the ability to carry out a randomized trial is crucial for obtaining reliable estimates of the program’s effectiveness.
Figure 2: Mean outcome per group over time.
Note: Share of subjects with a tech job per group across time. The month zero is the month of application to Dare IT. In the left figure, we group subjects according to the experimental groups: the treatment group in Mentoring (in green), the treatment group in Challenges (in red), and Out of Dare IT (in blue), where the last group consists of the applicants in control groups, non-selected applicants to Mentoring, and late applicants to Challenges. In the right figure, we group applicants into those selected and those not selected by mentors. "Selected and treated" group (in green) shows the outcomes of the treatment group in Mentoring, "Selected but not treated" group (in blue) is the control group of Mentoring "Not selected" group (in red) includes applicants to Mentoring that were not selected by mentors, both experimental groups of Challenges, and late applicants to Challenges.
In Table 2, we present the estimates of the average treatment effects for both the Mentoring and Challenges programs. Our analysis covers both short-term and longer-term outcomes: securing a job in tech within 4 months and 12 months after the programs, respectively. [^14] Tech job 4 months refers to 4 months after the end of a program which corresponds to approximately 8 months after application; analogously for Tech job 12 months. Additionally, we assess the influence of the programs on the acquisition of any new job within 4 months following the end of a program. For the short-term outcomes, we employ the difference-in-means estimator. [^15] Given the paired design of the mentoring experiment, we estimate variance as $\hat{V}(\hat{\tau}^{M})=\frac{1}{J\times(J-1)}\sum_{g=1}^{J}(\hat{\tau}^{M}_{j}-\hat{\tau}^{M})^{2}.$ (2) Athey and Imbens (2017) show that this is a conservative estimator of variance in a paired experiment. In the case of Challenges, we use the Splawa-Neyman estimator of variance. When estimating the longer-term effect, we utilize the fact that we observe monthly outcomes to estimate a Cox proportional hazard model (Cox, 1972). In Appendix C, we provide an estimate for short-term outcomes using the Augmented Inverse Propensity Weighting estimator (Robins et al., 1994; Athey et al., 2019). Furthermore, Appendix D contains alternative specifications of the Cox proportional hazard model, and the results align with those shown in Table 2. The test results of proportionality of hazard rates can be found in Appendix G.
Table 2: Average treatment effect estimates.
Mentoring Challenges
Difference-in-means Proportional hazard Difference-in-means Proportional hazard
New job Tech job 4 months Tech job 12 months New job Tech job 4 months Tech job 12 months
ATE 0.045 0.129 0.126 0.040 0.089 0.071
(0.05) (0.05) (0.025) (0.05) (0.04) (0.02)
ATE % baseline 10.749 43.941 36.887 13.636 45.306 46.972
(14.23) (17.09) (7.34) (12.00) (14.07) (10.01)
No. of obs. treatment 152 152 2265 183 183 2490
No. of obs. control 147 147 2205 225 225 3330
Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology within 4 months, and within 12 months. The first row shows estimates of the average treatment effect using the difference-in-means estimator in the case of new job and tech job 4 months (see Appendix C for AIPW estimates). The average treatment effect on Tech job 12 months is estimated using Cox proportional hazard model (we present estimates from a model with no covariates, see Appendix D for Cox proportional models with covariates and with time-varrying treatment effects). The second row presents the estimates from the first row divided by the mean outcome in the control group. Rows three and four present the number of observations per experimental group. The first, second, and third columns present the results of Mentoring intervention. Columns four, five, and six show the results from the Challenges program. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimates of the average treatment effect of Mentoring are estimated using Equation 2. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. Standard errors of the estimates from Cox proportional hazard model are estimated using the partial likelihood function (coxph implementation in R).
We find that participation in the Mentoring program significantly increases the probability of getting a job in the tech industry, by 13 percentage points within both 4 and 12 months after the program’s conclusion. The short-term effect’s estimate has a standard error of 5 percentage points, whereas the longer-term effect’s standard error is reduced to 2.5 percentage points, thanks to the proportional hazard model’s functional form assumption and our time-series data. The Challenges program also significantly enhances the likelihood of securing a tech job, although its impact is somewhat less pronounced than that of the Mentoring program. We estimate that participation in Challenges increases the chance of securing a tech job by 8.9 percentage points within 4 months (with a standard error of 4), and by 7.1 percentage points within 12 months (with a standard error of 2).
In the second row of Table 2, we express the average treatment effects as a percentage of the control group’s mean outcome. As previously noted, mentor selection leads to higher average outcomes in the Mentoring control group compared to the Challenges control group. As a result, when expressed in percentage terms, the impact of both programs appears quite similar. Participation in Mentoring increases the odds of securing a tech job within 12 months by 37%, whereas in Challenges by 47%. Importantly, we find no evidence of the treatment effect diminishing over time when comparing the programs’ effectiveness 4 and 12 months after their conclusion.
Finally, columns 1 and 4 of Table 2 display the estimated impact of participating in Dare IT programs on the probability of acquiring any new job within 4 months post-graduation. We do not observe any statistically significant difference in this outcome between treatment and control. Therefore, our results suggest that the treatment and control groups of both experiments managed to secure new jobs at similar rates. However, those in the treatment groups were more likely to find employment in the technology sector.
In Appendix F, we show the results based on the outcomes survey. We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
5.2 Heterogeneous treatment effects
We explore heterogeneity in treatment effects by conducting an analysis of treatment effects across different subgroups using the Cox proportional hazard model. We categorize applicants to the Mentoring and Challenges programs into non-overlapping groups based on four different types of characteristics: STEM education, length of professional experience, type of residence, and age. For each of these categories, we estimate a hazard model that includes treatment and subgroup indicators along with their interaction. The results of this analysis are presented in Table 3. It’s important to note that applicants in each of the non-overlapping groups may have other varying characteristics; to account for this, we include a full set of controls and interactions in Appendix I in Table 10. Furthermore, here we present only selected subgroups; Appendix I presents all the interaction terms.
Table 3: Heterogeneity in treatment effects: Mentoring vs. Challenges.
Mentoring Challenges
Baseline CATE Baseline CATE Diff. Mentoring - Challenges
STEM degree
Yes 0.437 (0.02) 0.030 (0.03) 0.191 (0.02) 0.030 (0.03) -0.001 (0.04)
No 0.239 (0.02) 0.232 (0.03) 0.135 (0.01) 0.086 (0.02) 0.146 (0.03)
Difference subgroups 0.199 (0.03) -0.202 (0.05) 0.055 (0.02) -0.055 (0.03) -0.147 (0.05)
Length of professional experience
Less than 5 years 0.429 (0.03) 0.008 (0.04) 0.119 (0.01) 0.136 (0.03) -0.128 (0.05)
More than 5 years 0.344 (0.02) 0.142 (0.03) 0.178 (0.01) 0.021 (0.02) 0.121 (0.03)
Difference subgroups -0.085 (0.03) 0.134 (0.05) 0.059 (0.02) -0.115 (0.03) 0.248 (0.06)
Type of residence
Small town 0.307 (0.03) 0.285 (0.05) 0.161 (0.02) 0.160 (0.03) 0.126 (0.06)
Warsaw 0.353 (0.02) 0.075 (0.03) 0.148 (0.01) 0.051 (0.02) 0.023 (0.03)
Difference subgroups -0.046 (0.03) 0.211 (0.05) 0.013 (0.02) 0.109 (0.04) 0.102 (0.07)
Age
Over 30 years old 0.261 (0.02) 0.125 (0.03) 0.130 (0.01) 0.114 (0.02) 0.011 (0.04)
Under 30 years old 0.428 (0.02) 0.117 (0.04) 0.178 (0.01) 0.025 (0.02) 0.092 (0.04)
Difference subgroups -0.167 (0.03) 0.008 (0.05) -0.049 (0.02) 0.089 (0.03) 0.011 (0.05)
Note: Estimates of average treatment effects conditioned on subgroups in Mentoring and Challenges. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction. The first and third columns show the subgroup’s average outcome within control subjects. The second and fourth columns show the difference between treatment and control in the subgroup. The last column shows the difference between Mentoring and Challenges of the treatment effects. Standard errors are in parentheses.
First, we observe that participants with a STEM degree have higher baseline outcomes than those without a STEM degree. Both programs, Mentoring and Challenges, are more effective for participants without STEM degrees, bridging the outcome gap across these two groups. Specifically, in the case of Mentoring, we find that the treatment entirely eliminates the STEM outcomes gap. When comparing the two programs, we find no statistically significant difference for participants with STEM degrees, while for those without STEM degrees, Mentoring leads to nearly 15 percentage points higher treatment effects than Challenges. Second, when focusing on the subgroups with more or less than 5 years of professional experience, we find that Challenges is more effective for participants starting their professional career, with nearly 13 percentage points higher treatment effects in this subgroup. Conversely, for participants with longer professional experiences, Mentoring appears more effective. Third, we find that participants from smaller towns benefit more from both programs. However, this difference is particularly pronounced in the case of Mentoring. Participants from smaller towns have lower baselines than participants from Warsaw, but after the treatment, the gap is entirely closed. In fact, Mentoring participants from smaller towns achieve, on average, higher outcomes than Mentoring participants from Warsaw. Finally, participants below 30 years old have higher baselines than those over 30 years old. However, both programs help close the gap between the two age groups by having higher treatment effects for older participants.
Additional characteristics from Mentoring.
In the case of Mentoring, we observe several additional characteristics; these come from the Mentoring registration survey, which was more extensive than the Challenges one, and we also observe from characteristics of mentors. In Table 4, we present additional results documenting heterogeneity in treatment effects along those additional dimensions.
Table 4: Heterogeneity in treatment effects: Mentoring.
Variable value is 1 Variable value is 0
Baseline CATE Baseline CATE Diff. Subgroups
Mentee characteristics
Mother 0.348 (0.02) 0.288 (0.04) 0.398 (0.02) 0.068 (0.03) 0.220 (0.05)
Applied before 0.384 (0.03) 0.141 (0.05) 0.330 (0.02) 0.121 (0.03) 0.020 (0.05)
Mentor characteristics
Managerial exp. 0.285 (0.02) 0.223 (0.03) 0.414 (0.02) 0.012 (0.03) 0.211 (0.05)
Long exp. tech 0.288 (0.02) 0.226 (0.04) 0.389 (0.02) 0.052 (0.04) 0.174 (0.05)
1st time mentor 0.354 (0.02) 0.131 (0.03) 0.342 (0.02) 0.112 (0.03) 0.019 (0.05)
Note: Estimates of average treatment effects conditioned on subgroups in Mentoring. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction. Applied before takes the value of 1 if the mentee has applied to the program in the past, but did not get selected into the program, and reapplied, and the value of zero if this is the first time applying to the program. Long exp. in tech takes the value of 1 if the mentor worked for over 5 years in the technology sector (the median value), and zero otherwise; 1st time mentor takes the value of 1 if this is the first time the mentor provides formalized mentoring service, and zero otherwise. Columns 2 and 3 describe the values for subgroups where the subgroup indicator takes the value 1, while columns 4 and 5 when the indicator takes the value of 0. The last column is the difference in the effectiveness of Mentoring across the two subgroups. Columns 1 and 3 are average outcomes in control groups, and columns 2 and 4 are the estimates of the average treatment effect. Standard errors are in parentheses.
First, we observe that the Mentoring program is particularly effective for mothers. Mothers who were selected by mentors but randomized into the control group had a baseline probability of 35% of finding a job in technology, compared to 40% for those without children. However, the average treatment effect for mothers is significantly larger at 29 percentage points, compared to just 7 percentage points for those without children. As a result, Mentoring not only enables mothers to catch up with applicants without children but even surpass them in the probability of getting a job in technology. In contrast, we do not observe a high difference in the effectiveness of Mentoring between first-time applicants and those who applied in previous editions.
Turning our attention to the characteristics of mentors, we find that mentors with managerial experience and those with a long history in the technology sector (over 5 years) tend to select, on average, less promising candidates, as indicated by baseline outcome probabilities of 29% for both groups, compared to the baseline of 41% for mentors without managerial experience and 39% for those with short tenure in tech. However, these mentors with managerial experience and long tech sector tenure are highly effective, with their mentees achieving a probability of over 50% of securing a job in technology. Interestingly, we also find that first-time mentors are as effective as mentors who have had previous mentoring experience. These findings are consistent with the notion that effective mentors leverage their knowledge, skills, and relationships accumulated over years of working in the technology sector, rather than improving their mentoring capabilities through the practice of mentoring.
Patterns of treatment effects heterogeneity are consistent with some of the mechanisms through which the two programs might impact employability that we hypothesized in Section 3. Participants from smaller towns and those without family or friends working in tech (see Table 10) might have less extensive professional networks in the technology sector, and mentors with longer experience in tech are likely to have richer networks. Thus, the finding that participants with these characteristics benefit more from Mentoring is consistent with the mechanisms that Mentoring helps mentees by bolstering their professional network and enhancing their capacity to utilize these networks effectively. As for Challenges, we hypothesized that the program would equip participants with credible signals of their technical skills. This mechanism is consistent with the finding that participants without STEM degrees and those with shorter professional experience, thus fewer other signals on their resumes, exhibit larger benefits.
6 Counterfactual program assignment policies
The number of spots in each program is limited. The capacity of Mentoring is constrained by the availability of mentors, while in Challenges, we capped spots to 300 to accommodate potential operational complexities and logistical issues. Considering the substantial treatment effects of both programs, one might posit that relaxing these capacity constraints will create significant value. Nonetheless, capacity constraints will likely remain a long-term issue for the mentoring program and may persist, albeit to a lesser degree, for Challenges. Consequently, Dare IT will continue being faced with the question of which applicants to accept for each program. With heterogeneous treatment effects, admission policies favoring applicants with high predicted benefits could enhance Dare IT programs’ impact. This section aims to quantify such potential gains.
We carry out three analyses: (i.) we estimate gains from the introduction of a policy that, for given capacity levels in each program, assigns applicants to programs to maximize the sum of predicted treatment effects across the two programs; (ii.) we compare the overall impact of Mentoring when mentors select most promising candidates to a case when they select to maximize expected benefits; ((iii). finally, we evaluate several simple policies based on heuristics.
An assignment policy is a mapping from applicants’ characteristics $X\in\mathcal{X}$ and programs capacity levels $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT. $\mathcal{Q}$ is the grid of possible capacity levels under consideration. Policy estimation and evaluation are carried out separately. The objective of policy estimation is to obtain a mapping from user characteristics and programs’ capacity level to program assignments’. This mapping $\hat{\pi}$ is estimated in a training set. Across all analyses, we obtain training and testing sets by randomly splitting data at the applicant level with equal probability. Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping to the testing set, where for each $i$ in the test set and each $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. Second, for each $Q\in\mathcal{Q}$, we construct the AIPW estimates of the “policy value,” that is, the value of the outcome when treatments are assigned according to $\pi(X_{i},Q)\in\mathcal{P}$. We do that using cross-fitting in the testing sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds (we use 100 repetitions with 10 folds and consider the median outcome across repetitions) [^16] We follow the approach suggested in Chernozhukov et al. (2017).. Throughout this section, we use Cox proportional hazard model with the full set of interactions as the outcome model and the logit model as the propensity model. Appendix J provides details of the two models and of the cross-fitting procedure.
6.1 Optimal assignment policies
Let $z_{ip}\in\left\{0,1\right\}$ be a variable indicating whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments maximizing the sum of treatment effects given by Equation 3:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\thickspace\forall p\text{ \& }\sum_{p\in\mathcal{P}}z_{ip}=1\thickspace\forall i,$ (3)
where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$. The first constraint ensures that the capacity limits are not violated, and the second one is that each applicant is assigned to one program. A policy that induces allocation $z_{ip}^{*}$ is what we call optimal policy. In Algorithm 1 in Appendix K, we provide further details of the estimation of optimal policy using training data, where we let $\hat{\pi}$ denote the estimate.
We estimate the optimal policy at three capacity levels: first, with the existing capacity (allowing 13% of all applicants into Mentoring and 15% into Challenges) – Optimal 1, second, increasing the capacity of Challenges to accommodate 50% of applicants – Optimal 2, and third, a scenario where each program can admit all applicants – Optimal 3. We compare these optimal policies against several alternatives: a policy accepting all applicants to Mentoring, a policy enrolling all applicants in Challenges (both these policies have increased capacities to accept all applicants), and the existing allocation policy (Status quo). We use the average treatment effect as the metric to compare policies. The results are presented in Figure 3.
Figure 3: Evaluation of counterfactual policies.
Note: Estimates of values of policies. The policy value is the mean probability of having a tech job across all participants and is measured by the mean of AIPW scores under allocation induced by each policy across all participants. Standard errors clustered at a participant level. Whiskers show confidence intervals. Optimal 1 has the existing capacity levels. Optimal 2 increases the capacity of Challenges to 50%, and Optimal 3 has no capacity limits.
First, we find that the optimal policy, with existing capacity levels, markedly improves the average treatment effect when compared to the Status quo, with an increase from 12 to 22 percentage points. Furthermore, this policy is also substantially more effective than policies that assign all applicants to Mentoring or to Challenges. This result follows from the high degree of treatment effect heterogeneity across both programs and shows that a targeting policy can substantially augment Dare IT’s effectiveness. Second, we find that the policy expanding the number of Challenges spots to 50% of applicants (Optimal 2) increases the average treatment effect to 17 percentage points and Optimal 3 to 15 percentage points. Such an increment highlights that the there are sizeable subpopulations amongst Dare IT applicants that have higher treatment effects. Moreover, we find only a small and statistically insignificant difference in policy value when contrasting Optimal 2 to the Optimal 3 policy.
In Table 5, we present the gains from optimal policy per assignment group. A program assignment group consists of all applicants that under a policy are assigned to this program, i.e., Mentoring assignment group are all applicants that are assigned to Mentoring. We are interested in comparing the benefits from the participation in each program of each of these groups under. We find that the group allocated to Mentoring substantially gains from their allocation, yielding an ATE of 0.199 (S.E. 0.017). This group would also substantially benefit from Challenges with the average treatment effect of 0.276 (S.E. 0.019). The difference between assigning them to Mentoring over Challenges is significant -0.077 (S.E. 0.021). However, the group assigned to Challenges benefits from this assignment even more. The average treatment effect of Challenges amounts to 0.244 (S.E. 0.017) compared to 0.109 (S.E. 0.012) of Mentoring. Lastly, the non-participating group has substantially lower benefits from assignment to Mentoring or to Challenges, average treatment effects of 0.101 (S.E. 0.006) and 0.073 (S.E. 0.003) respectively.
Table 5: Gains from the optimal assignment
Mentoring group Challenges group Nothing group
Share 0.15 0.50 0.35
ATE under assignment to:
Optimal 0.199 (0.017) 0.244 (0.017)
Mentoring 0.199 (0.017) 0.109 (0.012) 0.101 (0.006)
Challenges 0.276 (0.019) 0.244 (0.017) 0.073 (0.003)
Difference in ATE between assignments:
Mentoring - Challenges -0.077 (0.021) -0.135 (0.025) 0.028 (0.006)
Note: Gains from the optimal assignment by assignment group from Optimal 1 - 13% of applicants get be admitted to Mentoring and 15% to Challenges. Groups are defined as all applicants assigned to the specific program. The columns describe the outcomes of groups under various allocations - the group membership is determined using Algorithm 1 and is constant across the comparisons here, the allocations change. The first row is the share of applicants assigned to a group. Rows two to four describe the average treatment effect of the group under various assignments. The 5th row is the average treatment effect under Mentoring less the average treatment effect under Challenges. Standard errors are next to each value in parentheses and are clustered at the applicant level.
When comparing the three policies based on optimal allocation, it’s evident that the current capacity levels limit Dare IT’s impact but that additional spots do not create more value beyond a certain capacity. To further unpack this, in Figure 4, we analyze the value of optimal policy under a sequence of capacity values and compare with random assignment to either Mentoring or Challenges. Purple dots in Figure 4, show the average probability of getting a job in technology across all applicants (including those not participating in Dare IT programs), for the total capacity of the two programs ranging from 0 to 100% (in all cases, the two programs have the same capacity, so for example the total capacity of 50% means that, up to 25% of applicants get be assigned to either program). We compare the outcome under these policies with policies that assign applicants randomly to Mentoring (in yellow) and to Challenges (in red).
Figure 4: The value of targeting at various capacity levels
Note: Estimates of mean outcomes across all applicants at various capacity levels of Dare IT programs and allocation mechanisms. The policy value is the predicted share of applicants getting a tech job. The share of applicants getting a spot refers to the total capacity in both programs. Challenges Random (in orange) and Mentoring Random (in red) assign applicants at random between the program and Out of Dare IT. Targeting policy (in purple) follow the optimal policy approach.
Our analysis indicates that policies randomly admitting applicants to either the Mentoring or Challenges programs lead to consistently lower policy values when compared to the targeted policy across all capacity levels. When capacities are low, targeted policies prioritize applicants with the highest benefits. A meaningful difference in the average probability of getting a job across all applicants quickly emerges between the targeted and random policies; consequently, already at the 25% capacity levels, there is an approximately 4 percentage points gap between targeted and random policies. Interestingly, even when capacity constraints are lifted to accommodate all applicants, the targeted strategy continues to lead to higher outcomes than the random policies based on one program. This discrepancy remains because there of the difference in benefits that applicants receive across the two programs, i.e., there are some applicants that benefit more from Challenges than they do from Mentoring.
6.2 Counterfactual selection policies for mentors
Applicants selected by mentors and randomized to be the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2). This suggests that mentors tend to prefer highly promising applicants who are likely to secure a tech job irrespective of their participation in Mentoring. In this section, we examine alternative selection strategies that mentors could adopt and their potential implications for the effectiveness of the mentoring program. Specifically, we consider four scenarios: (i) Highest CATE – targeting applicants predicted to have the highest treatment effects, (ii) Most Promising – targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising – selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated – where mentors select participants that have the highest predicted outcomes if they participate in Mentoring; additionally, we include a case in which mentors select at random (Random Choice) and the observed allocation (Observed) as benchmarks. In each policy, we assume that 36% of applicants participate in Mentoring, which corresponds to the share of applicants that were selected by mentors. [^17] We consider a higher number of participants than the current number of spots in Mentoring to amplify the impact of policies on outcomes across all applicants; the ordering of policies does not depend on the share of applicants admitted to Mentoring as long as this number is constant across policies.
We follow the same off-policy approach as previously. We consider the dataset of all applicants to Mentoring, which consists of 151 individuals who participated in the program, 147 mentors selected but randomized to the control group, and 468 non-selected applicants. We split data into train and test with equal probability. We use the training dataset to estimate the mapping from applicants’ characteristics to assignments. To do that, we train the Cox proportional hazard model in the training data and predict outcomes with and without treatment in the testing data. Then, we use these predictions to determine which applicants get selected under each policy. To evaluate policies, we compute AIPW scores analogously to the optimal policy evaluation (Appendix J.1): using only testing data, we train the propensity model (logit model) and a new outcome model (Cox hazard model) using cross-fitting. We compute 100 cross-fitting rounds and take the median outcome. Table 6 shows the results.
Table 6: Mentors’ selection policies.
Policy Mean ATE Baseline outcome treated group Mean outcome all applicants
Highest CATE 0.178 (0.012) 0.187 (0.008) 0.232 (0.005)
Most Promising 0.062 (0.012) 0.285 (0.009) 0.192 (0.005)
Least Promising 0.127 (0.007) 0.055 (0.005) 0.211 (0.005)
Highest Chances if Treated 0.130 0.012 0.237 (0.008) 0.218 (0.005)
Observed 0.096 (0.012) 0.275 (0.008) 0.205 (0.005)
Random 0.094 (0.011) 0.191 (0.008) 0.203 (0.005)
Note: Outcomes and treatment effects across different mentors’ selection rules. Mean ATE is estimated using AIPW estimator with cross-fitting. Baseline outcome treated group is the predicted outcome without treatment for the group that receives treatment in the considered policy. Mean outcome all applicants corresponds to the average probability of getting a job in tech across all applicants, where the group assigned to treatment under considered policy receives treatment, and the remaining applicants do not participate. Standard errors clustered at the applicant level.
We find that prioritizing applicants with the highest anticipated treatment benefits results in higher average outcomes for all applicants and a more significant treatment effect than other strategies. Using the Highest CATE policy, 23% of applicants secure a tech job, a noticeable improvement over the 20% success rate under the Random or Observed policies. The average treatment effect under Highest CATE is an 18 percentage points increase in the probability of having a tech job for the treated group. Interestingly, targeting applicants with the highest chances of having a job if not treated (Most Promising) leads to treatment effects that are lower than selecting applicants randomly. Moreover, a policy aimed at the seemingly less promising applicants (Least Promising) — which goes against current mentor preferences — produces an average treatment effect of 13 percentage points. This results in an average outcome for all applicants surpassing those achieved by the Most Promising and Random policies. The policy of targetting applicants that have the highest chances of getting a tech job if treated (Highest Chances if Treated) results in a 13 percentage points average treatment effect, which is similar to Least Promising.
74% of applicants assigned to Mentoring under Highest CATE are also assigned to the program under Highest Chances if Treated, which is expected considering the fact that these two policies are most promising. However, there is a low overlap between Highest CATE and Least Promising only 7%. Furthermore, there is a substantial overlap between the observed selection by mentors and the Highest CATE policy – 54%, but a very low one between Observed and Least Promising, only 4%. These observations suggest that a group of applicants has low chances of securing a tech job when not treated and are systematically overlooked by mentors, even though their benefits from the program would be high. We analyze the overlaps and the characteristics of selected candidates in Appendix M.
6.3 Simple priority rules
Optimal assignment policy increases the average probability of getting a tech job across all applicants compared to the status quo. However, a non-parametric policy requiring constraint integer optimization is difficult to implement in Dare IT’s context. In this section, we examine whether a simple priority rule based on heuristics can enhance the average effectiveness of Dare IT’s programs.
Consider the admission policy wherein Dare IT first screens participants based on their priority level and then randomly assigns them to either the Mentoring or Challenges program (with the probabilities proportional to the number of spots in either of the programs). We propose the following priority groups: Priority 1 – Applicants who are from small towns, do not have a degree in a STEM field, and do not have a master’s or postgraduate degree; Priority 2 – Applicants who are from small towns, do not have a STEM degree, but have a master’s or postgraduate degree; Priority 3 – Applicants who are not from small towns and do not have a STEM degree; Priority 4 – Applicants who have a STEM degree.
The proposed priority groups are non-overlapping and together encompass all applicants. Thus, under this proposed admission policy, Dare IT would first consider the applicants in the first group. If there is more capacity after admitting the applicants in the first group, they will move to the second group, and so on. We are interested in evaluating whether the groups of applicants with high priority have higher predicted treatment effects than groups with lower priority levels and whether the high-priority groups are large enough to fill in all the existing spots in the two programs.
Given that this policy is defined using heuristics instead of data, we do not split the dataset for policy estimation and separate evaluation. Instead, we use the entire dataset to evaluate this pre-defined policy. To estimate the value of the policy, we consider both experimental populations and the remaining applicants. For each applicant, we estimate outcomes under three conditions: Mentoring, Challenges, and Out of Dare IT. We do that using cross-fitting as described in Appendix J.1. The outcome model is the proportional hazard Cox model, and the propensity model is the multinomial logit model; we combine them to obtain AIPW estimates. Admitted applicants are assigned with probability 45% to Mentoring and 55% to Challenges (reflecting the proportions of treated subjects in the dataset). Once we have the allocation, we compute the average treatment effect per group using the AIPW predictions. Figure 5 shows the results.
Figure 5: Evaluation of simple policies.
Note: Estimates of the average treatment effect per priority group; AIPW estimates obtained using cross-fitting. The treatment effect is the percentage point increase in the probability of a tech job. The red dotted line is the average treatment effect across the entire population. Whiskers around point estimates represent confidence intervals. Standard errors are clustered at the participant level.
We find that groups 1 and 2 have predicted average treatment effects that are higher than the average across the entire population of applicants (red scatter line) and higher than groups 3 and 4. Furthermore, group 1 encompasses 14% of applicants and group 2 10%. The combined capacity of Mentoring and Challenges is 28%. Thus, the two groups can almost fill in all the spots in the two programs. We find that this simple priority rule can increase Dare IT’s effectiveness.
7 Conclusion
Changing labor markets push workers to want to transition into high-growth sectors, and facilitating these transitions is an important goal for labor policy. However, effectively designing such policies requires principled evaluation of various training programs in terms of their overall impact and their heterogeneity. In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector. This way we identify a class of interventions that can effectively narrow the gender gap in technology.
We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by more than 40% four months after the end of the program. Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
Our findings reveal strong patterns of heterogeneity in treatment effects. The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age. For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring. Furthermore, certain groups benefit distinctly from one program over the other. Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring.
For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit. We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects. Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness. Specifically, we estimate a seven percentage points increase in the average probability of getting a job in the technology sector across all applicants. Notably, such targeted policy utilizing the current combined 28% capacity of both programs, yields outcomes comparable to policies that accommodate all applicants into either Mentoring or Challenges. This showcases that in this context targeting admission has a very high potential impact.
A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees. To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs. This resulted in three categories: applicants not chosen by mentors, those shortlisted but placed in the control group through randomization, and those shortlisted and included in the treatment group. Importantly, all applicants underwent an assessment of their technical skills before being presented to mentors. By the end of our observation period, there was a 20 percentage point difference in securing tech jobs between those selected but placed in control and the non-selected applicants, with the latter securing job in technology at the probability of 22%. This underscores that mentors lean towards candidates they deem more promising. While this approach boosts the success rate of Mentoring, reflected in the proportion of participants landing tech jobs, it does not necessarily maximize impact since many of these candidates might have secured tech roles even without Mentoring. Through an off-policy evaluation, we explored potential alternative mentor selection strategies. Our findings suggest that mentors focusing on participants with the greatest treatment effects, rather than those already poised for success, can significantly amplify the efficacy of Mentoring. We estimate that when mentors select 40% of participants based on the highest treatment effect, the share of applicants eventually finding a job in technology is 31% compared to 21% when they choose participants that have the highest outcomes if treated, and 24% when they select participants at random.
This study comes with certain limitations worth noting. First, the limited participant size in both programs led to wide confidence intervals in several estimates, leaving room for uncertainty regarding the exact impacts of Mentoring and Challenges. Secondly, we could not pinpoint the exact mechanisms driving the outcomes for both programs. Although our findings suggest that Mentoring likely enhances participants’ professional networks and teaches how to utilize them, and Challenges supports participants in the development of a credible signal of their skills, a more in-depth exploration is required to understand the specific mechanisms through which the training provided by both programs impacts employability. Such an analysis could also help scale up Mentoring by identifying the crucial aspects of the program and directing Dare IT resources to them. Lastly, the participants from the first edition of Challenges may not be representative of those in later editions. Early adopters of such programs might experience greater benefits. To validate the program’s continued efficacy, subsequent evaluations are essential.
Appendix
Appendix A Challenges program
The Challenges program is designed for the same population as the Mentoring program: individuals who possess technical competencies but have not yet secured employment in the technology sector. The primary aim of the Challenges program is to assist these individuals in developing credible credentials that can effectively signal their skill set. In its first edition, the Challenges program provided two distinct tracks: User Experience Design (UX) and Front-end Development in React (front-end). The selection of these specific tracks was influenced by several determinants. Primarily, there was a significant volume of applications for the Mentoring program related to these two domains. Furthermore, Dare IT possessed both internal expertise and established partnerships in these areas. Lastly, a pronounced market demand existed for professionals in these specializations in Poland.
To achieve the program’s scalability goal and streamline the application process, unlike the Mentoring program, the review of the minimum technical skills during application was based solely on applicants’ declaration. The registration website explicitly mentioned the necessity for an advanced understanding of Figma for those interested in UX design, and React for those leaning towards front-end engineering as prerequisites for the Challenges program. The applicants had to confirm that they indeed have the minimum skill level. Additionally, English proficiency was mandated due to the inclusion of course content delivered in English.
The Challenges program spanned 16 weeks. Initially, the first two weeks focused on an introductory event and a refresher course in Figma and React. The concluding two weeks were allocated for the refinement of the final projects, HR workshops, and a closing event. Central to the program, the intermediate 12 weeks were primarily aimed at the progression of an individual project. Within both specializations, experts from partner organizations, specifically Flying Bisons and Grover – Polish tech companies, were asked to prepare six tasks. These tasks, inspired by their recent undertakings, encapsulated the comprehensive procedure of designing a mobile application for the UX domain and detailed the facets of front-end engineering for the respective specialization.
The program is based on two-week cycles culminating on Thursdays when submissions are expected and news assignments are posted. Every subsequent Friday, post-submission, an evaluative session with a subject-matter specialist takes place. During this session, a selection of submissions undergoes review, identifying prevalent errors and underscoring industry best practices. This grants participants the opportunity to amend and refine their submissions from the previous week. The domain experts, either mentors from Dare IT or other professionals with an appropriate level of seniority, carry out the reviews. Each assignment has been designed to take approximately eight hours of time to finalize. As the program approaches its conclusion, participants engage with a Human Resources specialist. This guidance is primarily aimed at instructing participants on leveraging their newly-acquired qualifications, both as skill indicators in job applications and as a resource during subsequent interviews.
In the first week of the UX path, participants conceptualize their project. They are given the flexibility to either devise their own application idea or choose from two provided project themes, followed by a detailed analysis of market competitors and benchmarks. The subsequent phases guide participants through a structured journey from ideation to delivery. They familiarize themselves with tools like Personas, User Journey Maps, and User Stories, ensuring a systematic organization of the information gathered. As the program advances, participants delve deeper into content hierarchy, Information Architecture, and user pathways, culminating in the creation of UX wireframes. The latter weeks emphasize usability testing, where participants craft interactive prototypes, refining them based on feedback. Crucially, they also submit project presentations and receive feedback on them.
In the case of the front-end path, in the first weeks, participants created reusable UI components, often termed Atoms. This initial phase emphasized the foundational importance of efficient design. In the subsequent weeks, attendees delved into topics of server communication: creating APIs, handling different requests, and DELETE queries. Following these modules, the middle segment of the program trained participants in creating models and form-based data collection. Later, the curriculum shifted towards data visualization using charts and the integration of real-time user feedback systems, e.g., notifications. These modules were designed to underscore the advantages of adept data presentation and heightened user experience. Finally, the program introduced participants to TypeScript, showcasing its emerging prevalence in modern web applications.
In addition to the assignments and review sessions, participants were granted access to the exclusive Dare IT community on the Slack platform. During the kick-off meeting, the participants were recommended to collaborate in teams to tackle the assignments. A dedicated process was created to support collaboration; participants were asked whether they wanted to be matched with a group. If they were interested, they received access to a group’s channel. [^18] The emphasis on collaborative efforts and peer support was further amplified in the ensuing editions of the Challenges program, where matching events were organized. Within each path’s Slack channel, distinct discussion forums were established, allowing participants to pose queries, circulate resources, and engage in discourse about assignment components. Activity within this Slack community was consistently high, with frequent discussions across multiple channels.
The first edition of Challenges has attracted a lot of popularity amongst the participants, applicants that we could not accommodate due to capacity limits, and the Polish tech community more broadly, thereby galvanizing its further development. Currently, the program is instituted on a trimestral cycle, extending its offerings to encompass Manual Testing, Automated Testing in Python, Cloud Computing, and User Interface Design. Furthermore, in the summer of 2022, the program was adapted to cater to Ukrainian refugees in Poland. This edition was delivered in Ukrainian, incorporating bespoke modules tailored to guide foreigners in navigating the Polish job market.
Appendix B Covariate balance
In Figure 6 we show the balance of covariates in the mentoring experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. The only difference that is statistically significant is the graduate degree, where 66% of treated subjects have a graduate degree and 55% of control (p-value 0.05).
Figure 6: Covariate balance mentoring.
Note: Balance across treatment and control of the main variables. Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
In Figure 7 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances. None of the differences is statistically significant.
Figure 7: Covariate balance Challenges.
Note: Balance across treatment and control of the main variables. Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
Appendix C Short-term average treatment effects estimates from the AIPW estimator
In Table 7, we present the estimates of the average treatment effects using the Augmented Inverse Propensity Weighing estimator (Robins et al., 1994). AIPW is a doubly-robust method which means that it adjusts for covariates in the outcome model and the propensity score. We use the grf implementation of the AIPW estimator, as in Athey et al. (2019). We use the full set of covariates as controls. We find that the estimates from the AIPW estimator are very similar to the estimates obtained using the difference-in-means estimator.
Table 7: Average treatment effect estimates.
Mentoring Challenges
Tech job 4 months New job Tech job 4 months New job
ATE AIPW 0.129 0.037 0.088 0.0437
(0.05) (0.06) (0.04) (0.04)
ATE 0.129 0.045 0.089 0.040
(0.05) (0.05) (0.04) (0.05)
Mean treatment 0.421 0.467 0.284 0.333
(0.04) (0.04) (0.03) (0.03)
Mean control 0.293 0.422 0.196 0.293
(0.04) (0.04) (0.03) (0.03)
No. of obs. treatment 152 152 183 183
No. of obs. control 147 147 225 225
Note: The estimates of the average treatment effects of both the mentoring program and Challenges on the probability of finding a new job in technology and finding any new job. The first row shows the estimates using the augmented inverse propensity weighting (AIPW) estimator. The second row shows estimates of the average treatment effect using the difference-in-means estimator. Rows three and four present mean outcomes in treatment and control groups and rows five and six the number of observations per experimental group. The first and second columns present the results of the mentoring intervention on the probability of finding a job in technology and the probability of finding any new job. Columns three and four show the results of the impact of participation in Challenges on the probability of finding a job in technology and any new job. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimates the average treatment effect of mentoring estimated using equation 2. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. AIPW variance estimated using grf implementation of AIPW estimator (see Athey et al. (2019) section 4 for details).
Appendix D Long term treatment effects
In this, section we present the estimates from two Cox proportional hazard models for both Mentoring and Challenges. First, we estimate a model with a full set of controls and second, we consider a model where the treatment variable has time-varying coefficients and the full set of controls. Table 8 shows the results. We estimate the coefficient of 0.26 in the case of Mentoring, which corresponds to a 29% increase in the probability of having a tech job. We find higher treatment effects in the case of Challenges, a coefficient of 0.37, corresponding to a 45% increase. We also consider the model in which we analyze the impact of participation in Dare IT programs across time - columns 2 and 4 of Table 8. We confirm the observation from Figure 2, treatment effects (especially in the case of Challenges) are realized predominantly after the end of the two programs. Furthermore, we do not observe attenuation of the treatment effect over time.
Table 8: Long-term treatment effects
Dependent variable: tech job
Mentoring Mentoring time-varying Challenges Challenges time-varying
Treated ${}^{***}$ (0.0501) ${}^{***}$ (0.065)
Treated:during program 0.196 (0.130) --0.074 (0.239)
Treated: 4 months after the program ${}^{***}$ (0.071) ${}^{***}$ (0.094)
Treated: 5 months plus after the program ${}^{***}$ (0.082) ${}^{***}$ (0.096)
Full set of controls x x x x
Observations 4,275 4,275 5,490 5,490
Note: Selected coefficient estimate from the proportional hazard models. The first column is the mentoring experiment. The second column is the mentoring experiment with time-varying coefficients. The third column is the challenges experiment. In the last column, the challenges experiment with time-varying coefficients. In Appendix G, we present the Cox proportional hazard model diagnostics.
Appendix E Comparison of salary levels from Glassdoor
We consider two types of jobs as a job in technology:
  1. 1.
    All jobs in tech companies other than finance, regulatory, legal, accounting, and HR, where tech companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online sales).
  2. 2.
    Jobs in non-tech companies that involve software development and testing, IT support, and data analytics. In our context, this category includes jobs in banks and management consulting agencies.
We analyze whether the two categories of jobs pay similar salaries. To do that we searched for salary estimates of all unique job titles that fall into the two categories using Glassdoor.com. We considered only exact matches and salary estimates for jobs in Poland. In total, we have 265 unique salary estimates. We find that the mean monthly salary for tech jobs in tech firms (category 1) is PLN 8634 and for tech jobs in non-tech firms it is PLN 8866. The difference is statistically insignificant (p-value 0.9). Figure 8 shows histograms of of salary estimates.
Figure 8: Histograms of salary estimates from Glassdoor.
Note: Histograms of salary estimates from Glassdoor. In green, we show tech jobs in tech firms (category 1), in blue tech jobs in non-tech firms (category 2).
Appendix F Outcomes measured in surveys
We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes. There were three surveys: two main ones carried out approximately four months after the end of each program, and a shorter survey right after the end of the mentoring program from which we obtained information on salaries in jobs found during or right after the mentoring program. As discussed in Section 4, results based on the outcomes survey suffer from a selection problem. Thus, we interpret the findings based on this data as suggestive evidence.
The number of job offers.
Figure 9 shows the number of job offers received by survey respondents broken by experimental groups. We find that subjects in the control group (in green) were more likely to report receiving zero job offers than the subjects in the treated group (in orange); for any non-zero range, there is a higher share of the treatment group rather than the control group. The largest difference between the experimental groups is amongst subjects reporting more than ten job offers.
Figure 9: Reported numbers of job offers: Mentoring.
Note: The number of job offers received by subjects in the mentoring experiment that responded to the survey. Each bar corresponds to the share of respondents from an experimental group. Orange represents the treated group, and green represents the control group.
In Figure 10, we present evidence of the differences in the number of job offers across treatment and control groups of the Challenges experiment. We find results analogous to those from the mentoring program. The lowest category, here less than two job offers, has a higher share of control group respondents than treated group respondents. In contrast, treated subjects are more likely to report a higher number of job offers. The difference between treatment and control is more pronounced in the mentoring program.
Figure 10: Reported numbers of job offers: Challenges.
Note: The number of job offers received by subjects in the Challenges experiment that responded to the survey. Each bar corresponds to the share of respondents from an experimental group. Orange represents the treated group, and green represents the control group.
Salary ranges.
Figure 11 presents the share of respondents to mentoring outcomes surveys declaring a salary within a salary range (ranges were pre-defined in the survey). Orange bars indicate the treatment group, and the green ones are the control group. The right panel is based on the survey carried out right after the end of the program, while the left panel is from the survey conducted four months after the end of the program.
Figure 11: Salary ranges: Mentoring.
Note: Reported salary ranges by subjects who found new jobs after the start of the mentoring program. Orange represents the treated group, and green represents the control group. The length of a bar represent the share of respondents declaring a salary within the range. The right panel survey was conducted right after the end of the program; the left panel survey was conducted four months after the program.
Evidence presented in Figure 11 suggests that subjects in the treated group of the mentoring experiment that found a new job are more likely to have a salary in a high salary range.
In Figure 12, we report analogous results from the Challenges outcome survey. We find that a higher share of the survey respondents in the control reported salary in the lowest range as compared to the treatment group. We also find that the opposite is true for the highest salary range.
Figure 12: Reported salary range Challenges.
Note:Reported salary ranges by subjects who found new jobs after the start of the Challenges experiment. The length of a bar represent the share of respondents declaring a salary within the range. Orange represents the treated group, and green represents the control group.
Negotiation.
Respondents that reported finding a job were asked whether they negotiated the conditions of their employment. Amongst the mentoring group, 33% of treated subjects that found a new job negotiated the terms of the new job, while only 24% of the control group did so. The difference is statistically insignificant.
Appendix G Proportional hazards assumption
The Cox proportional hazards model assumes the hazard rate of an event is a product of a baseline hazard and an exponential function of the covariates. Thus, in the analysis, we assume the effects of different predictors are constant over time. To test that, we correlate Schoenfeld residuals from the baseline Mentoring and Challenges models with months. Table 9 presents the results. We do not find any statistically significant correlations.
Table 9: Proportional hazards assumption test
Mentoring model Challenges model
chisq1 df1 p1 chisq2 df2 p2
Treatment 0.232 1 0.630 1.301 1 0.254
City 0.149 2 0.928 1.863 2 0.394
Degree 1.909 4 0.752 5.832 4 0.212
STEM 0.001 1 0.971 0.002 1 0.968
Prof experience 0.187 1 0.665 1.723 1 0.189
Social science 0.074 1 0.785 - - -
Family friends IT 0.074 1 0.785 - - -
Mother 0.164 1 0.685 - - -
Applied before 0.087 1 0.768 - - -
First time mentor 0.956 1 0.328 - - -
Managerial experience 0.762 1 0.383 - - -
Career changer 0.043 1 0.836 - - -
Years of experience Mentor 0.428 1 0.513 - - -
Age 2.716 3 0.438 - - -
GLOBAL 9.595 20 0.975 13.130 9 0.157
Note: Diagnostics of the proportional hazard model. The first three columns show results for the Mentoring model, and the last three for Challenges. City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is also a discrete variable with levels at: no education mentioned, non-traditional credential, college degree, masters, and postgraduate. Prof experience is the number of years between the start date of the first employment of program participants and 2023.
Appendix H Bounds on the average treatment effect
In this section, we perform three robustness checks on the average treatment effects. First, we consider the AT bounds method introduced by Lee and Weidner (2021). Second, due to changes in LinkedIn profile URLs we could not collect outcome data from 17 Challenges participants. To obtain bounds on the treatment effects, we consider two extreme cases - none of the 17 subjects found a job in technology, and all of them did. We estimate treatment effects in both cases. Lastly, we had four participants drop out of the Mentoring program because they secured jobs in the tech industry before the start of the program. We establish bounds based on this information, assuming two distinct scenarios: (i) all corresponding control group participants also found jobs in technology, and (ii) none of them did.
Lee and Weidner (2021) proposes bounds on the average treatment effects that are designed to be robust in situations where there are many conditioning variables, or they have many values, or when the overlap condition is violated. We use R implementation of their algorithm (Lee and Weidner, ).
In the case of Mentoring we consider the following variables: Social science, STEM, Family or Friends in IT, Mother, Applied Before, First-time mentor, Managerial Experience, Career changer, Mentor’s Years of Experience, Professional Experience. We estimate the bounds on the average treatment effect to be 0.11 to 0.18 with a confidence interval of 0.03 to 0.26. In the case of Challenges, we consider the following characteristics: social science, STEM, Specialization, Residence, Age, Professional experience, Degree. We obtain bounds of 0.10 to 0.12 with a confidence interval of 0.02 to 0.20.
Next, we address the unavailability of LinkedIn data for 17 Challenges participants. We explore two extreme scenarios: one where all 17 did not secure tech-sector jobs and another where all 17 did. We estimate the average treatment effect in these scenarios using a difference-in-means estimator, without controlling for observable features due to the lack of LinkedIn data. When we assign zero outcomes for the missing subjects, the estimated average treatment effect is 6.6 percentage points, with a confidence interval ranging from -1 to 15 percentage points. Assuming positive outcomes for all missing variables, we estimate the average treatment effect at 15 percentage points, with a confidence interval spanning 7 to 20 percentage points.
Finally, we conduct a robustness check of our evaluation of the Mentoring experiment by examining two scenarios: (i) all four control group subjects also found tech jobs, and (ii) none of them did. We estimate treatment effects using the difference-in-means estimator. In the pessimistic scenario, the estimated treatment effect is 12 percentage points, with a confidence interval of 2 to 23 percentage points. In the optimistic scenario, we estimate the treatment effect at 15 percentage points, with a confidence interval spanning 4 to 26 percentage points.
Appendix I Heterogeneous treatment effects
In Table 10 we show estimates from the Cox model including the full list of interactions.
Table 10: Heterogeneous treatment effects of the interventions.
Dependent variable: tech job
Mentoring Challenges
Treatment ${}^{***}$ (0.265) ${}^{***}$ (0.239)
Social science --0.124 (0.125)
STEM ${}^{***}$ (0.099) ${}^{***}$ (0.111)
Family friends IT ${}^{***}$ (0.092)
Small Town $*{}^{*}$ (0.102) $*{}^{*}$ (0.126)
Warsaw 0.088 (0.094) ${}^{***}$ (0.137)
Mother ${}^{***}$ (0.137)
Applied before ${}^{***}$ (0.096)
First time mentor 0.123 (0.084)
Managerial experience ${}^{***}$ (0.087)
Career changer ${}^{***}$ (0.099)
Years of experience mentor --0.003 (0.013)
Non-traditional credential ${}^{***}$ (0.188) ${}^{***}$ (0.179)
Masters ${}^{***}$ (0.171) 0.218 (0.135)
No education mentioned ${}^{**}$ (0.193) $*{}^{*}$ (0.287)
Postgraduate $*{}^{*}$ (0.188) ${}^{***}$ (0.194)
Age: 25-30 $*{}^{*}$ (0.145) ${}^{**}$ (0.155)
Age 31-35 ${}^{***}$ (0.168) 0.170 (0.171)
Age: 36+ ${}^{**}$ (0.198) --0.144 (0.184)
Prof. experience ${}^{***}$ (0.013) 0.008 (0.011)
Treatment×$\times$×Social science 0.273 (0.179)
Treatment×$\times$×STEM ${}^{***}$ (0.136) ${}^{***}$ (0.160)
Treatment×$\times$×Family friends IT $*{}^{*}$ (0.119)
Treatment×$\times$×Small ${}^{***}$ (0.131) ${}^{***}$ (0.173)
Treatment×$\times$×Warsaw ${}^{***}$ (0.129) ${}^{***}$ (0.182)
Treatment×$\times$×Mother ${}^{***}$ (0.164)
Treatment×$\times$×Applied before --0.155 (0.123)
Treatment×$\times$×First time mentor --0.079 (0.108)
Treatment×$\times$×Managerial experience ${}^{***}$ (0.112)
Treatment×$\times$×Career changer ${}^{***}$ (0.125)
Treatment:$\times$×Years of experience mentor 0.001 (0.016)
Treatment×$\times$ Non-traditional credential ${}^{***}$ (0.235) ${}^{***}$ (0.251)
Treatment×$\times$ masters ${}^{***}$ (0.209) --0.267 (0.192)
Treatment×$\times$ No education mentioned ${}^{***}$ (0.347) --0.139 (0.349)
Treatment×$\times$×postgraduate ${}^{**}$ (0.236) ${}^{***}$ (0.243)
Treatment×$\times$×Age: 25-30 ${}^{***}$ (0.203) ${}^{***}$ (0.209)
Treatment×$\times$×Age:31-35 ${}^{***}$ (0.241) ${}^{***}$ (0.234)
Treatment×$\times$×Age: 36+ 0.273 (0.277) --0.013 (0.235)
Treatment×$\times$×Prof experience ${}^{***}$ (0.016) $*{}^{*}$ (0.015)
Observations 4,440 4,965
[…]
Log Likelihood --11,845.370 --6,329.277
Wald Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
LR Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
Score (Logrank) Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
Note: Estimates from the Cox proportional hazard model. Outcome variable tech job.
Appendix J Models used in off-policy sections
This section presents further details about the propensity and outcome models used in Section 6 for policy estimation. To estimate both models, we first join the Mentoring, Challenges, and applicants’ datasets. We obtain a dataset of 1314 individuals who applied to Dare IT, out of whom 151 participated in Mentoring and 166 in Challenges; the remaining 997 are the non-selected Mentoring applicants, late applicants to Challenges, and the control groups. All individuals are observed over the span of 15 months. The dataset is randomly split between training and testing at the individual applicant level, 50-50.
Propensity model.
To estimate the propensity model for policy estimation, we use multinomial logit regression with the full set of covariates that overlap between all datasets. We train the model in the training sample and predict into the testing set. We find that the propensity of being assigned to Mentoring ranges from 4% to 37%, to Challenges from 5% to 32%, and to Out of Dare IT from 53% to 89%. Figure 13 presents the histogram of the three propensities.
Figure 13: Predicted program propensities.
Note: Test set propensity of assignment across groups, in blue to Challenges, in red to Mentoring, and in green to Out of Dare IT.
We also evaluate the adjustment with the propensity weights by compare the standardized absolute mean differences in covaraites values with and without the adjustment. In Figure 14, we show the difference in average values of covariates between Mentoring and Nothing groups, with and without adjustment in each cross-fitting round. The same analysis for the case of Challenges and Nothing is in Figure 15. We find that the propensity model successfully adjusts for the differences in covariates.
Figure 14: Covariate balance: Mentoring.
Note: Differences in mean covariates across Mentoring and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
Figure 15: Covariate balance: Challenges.
Note: Differences in mean covariates across Challenges and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
Outcome model.
We use Cox proportional hazard model to model outcomes. We estimate a model with a full set of covariates and interactions of covariates and treatments. We model outcomes at the person and month level. In Figure 16, we show an evaluation of the model’s predictive performance. First, we train a model on the training sample and predict probabilities of having a job in a specific month using the testing sample; then, we take the average per user and group users based on these averages into quintiles. Finally, we compare the predicted probability within the quintile to the observed probability. We find that the differences are statistically insignificant, and the magnitudes are small. We conclude that the model is well-calibrated.
Figure 16: Comparison of predicted and observed outcomes.
Note: Comparison of predicted and observed outcomes in the test set. Subjects are grouped into quintiles based on the predicted probability of having a tech job. In red, the average per quintile predicted probability. In blue, is the average per quintile observed outcome. Whiskers show confidence intervals.
J.1 Cross-fitting.
We estimate predicted outcomes under a policy using cross-fitting with 10 folds and 100 repetitions. Specifically, for each repetition, $m\in\left\{1,...,100\right\}$, we randomly assign observations into 10 folds $k\in\left\{1,..,10\right\}$. Randomization is at the individual subject level with equal probability. Next, for each fold $k$ we train an outcome and propensity models using data from all other folds and predict outcomes and treatments in the $k$-th fold: $\hat{\mu}_{k}(X_{i})_{P}^{m}$ and $\hat{e}_{k}(X_{i})_{P}^{m}$, where $P\in\left\{0,M,C\right\}$ refers to a program. We use these predictions to construct AIPW estimates; for example, the treatment effect from Mentoring instead of Out of Dare IT writes:
$\tau^{m}_{OM}=\frac{1}{n}\sum_{i=1}^{n}\left[\hat{\mu}_{k}(X_{i})_{M}^{m}+\frac{\mathbb{1}_{W_{i}=M}}{\hat{e}_{k}(X_{i})^{m}_{M}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{M})-\hat{\mu}(X_{i})_{0}^{m}+\frac{\mathbb{1}_{W_{i}=0}}{\hat{e}_{k}(X_{i})^{m}_{O}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{O})\right],$
analogously for Challenges. Finally, following Chernozhukov et al. (2017), we take the median over 100 repetitions so that $\hat{\tau}_{OM}=Median(\tau^{m}_{0M})$.
Appendix K Policy estimation algorithms
Algorithm 1 describes the estimation and evaluation of optimal policy.
Algorithm 1 Optimal policy
1. Considering the dataset of all applicants to both programs, randomly split the dataset into train and test with equal probability at the applicant level, denote the resulting sets as $I^{train}=\left\{1,..,\mathcal{I}^{train}\right\}$ and $I^{test}=\left\{1,..,\mathcal{I}^{test}\right\}$.
2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
3. In the test set, construct predicted treatment effects using predictions from model $\mu_{train}$. Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p)_{train}-\hat{Y}(O)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p)_{train}$ is the predicted outcome under program $p$ constructed using $\mu_{train}$ and $\hat{Y}(O)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months. Compute mean treatment effects per user as: $\tau^{Op,train}_{i}=1/t\times\sum_{t=1}^{15}\tau^{Op,train}_{i,t}$,
4. Assign treatment to maximize treatment effects subject to capacity constraint. Let $Q^{p}$ be the capacity limit of program $p$ and $z_{ip}$ an indicator variable taking the value of one when applicant $i$ is assigned to program $p$ and zero otherwise. We solve the following constrained optimization problem:
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{i=p}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
The first constraint ensures that the capacity constraints are not violated. The second one is that every applicant is assigned to one program. There is no capacity limit on being Out of Dare IT. We use LP Solve algorithm to solve the problem. We obtain optimal allocation $\mathcal{A}_{X,Q}^{*}=\left\{a_{i}^{*},...,a_{\mathcal{I}^{test}}^{*}\right\}$,
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. See Appendix J.1 for details of the cross-fitting procedure. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
6. Obtain $\hat{V}_{X,Q}^{*}=\frac{1}{|I^{test}|}\times\sum_{i=1}^{\mathcal{I}^{test}}\hat{Y}_{i,k}(a^{*})$ as the mean of predicted outcomes under the allocation $\mathcal{A}_{X,Q}^{*}$. Estimate standard errors clustered at the applicant level: $\sigma_{X,Q}^{*}$. $\triangleright$
return $(\mathcal{A}_{X,Q}^{*},\hat{V}_{X,Q}^{*},\sigma_{X,Q}^{*},\hat{Y}_{1,k}(a^{*}),...,\hat{Y}_{\mathcal{I}^{test},k}(a^{*}))$
Appendix L Validate AIPW approach
In Section 6, we used estimated AIPW scores to construct and evaluate counterfactual policies. Now, we compare the estimates of treatment effects using the AIPW approach in the testing set with cross-fitting and the experimental results. Thus, we want to test the hypotheses that $\tau^{AIPW}=\tau^{RCT}$. To estimate $\tau^{AIPW}$, we proceed in the same steps as in Section 6: first, we pool all datasets: Mentoring treatment and control, mentoring non-selected applicants, Challenges treatment and control, challenges late applicants. We split the dataset into train and test, and in the testing dataset, we randomly assign 10 folds. Then for each fold, we train a proportional hazard model for the outcomes and a multinomial logistic regression for propensities using other folds, and then we predict outcomes and propensities for the left-out fold and construct AIPW scores. In the outcome model, we include an indicator variable of whether the applicant was selected by a mentor or not. We repeat the process of cross-fitting 100 times and take the average AIPW score per observation.
To assess Mentoring, we take the difference between predicted AIPW scores under mentoring treatment and under mentoring control, i.e., the estimate of the average treatment effect for the entire sample, analogously for Challenges. We compare obtained quantities with the estimates of the treatment effects using Cox models on the experimental samples. We also compute the difference between the estimates. Figure 17 shows the results
Figure 17: Difference between AIPW and RCT estimates.
Note: Comparison of the average treatment effect estimates using the randomized experiment and the AIPW scores. Point estimates and confidence intervals.
We find that the RCT and AIPW estimates are similar, and the differences are statistically not distinguishable from zero. Recall that the randomization into treatment and control was carried out within experimental samples, that is, the sample of Mentoring applicants who mentors selected and the subset of applicants to Challenges that applied to the program before we closed the registration. Thus, it was not a priori clear that the AIPW procedure with cross-fitting would recover similar estimates of the treatment effect after we pooled all datasets together. The finding that these estimates are very close in magnitude and statistically not different from one another bolsters the credibility of results presented in Section 6.
Appendix M Mentors’ selection strategies
In this section, we dive deeper into four mentors’ strategies: Observed, Highest CATE, Highest if treated, and Least promising. We analyze the degree to which they overlap in terms of which candidates get selected and how the characteristics of potential mentees differ across these strategies.
First, we present the Venn diagram in Figure 18. We observe considerable overlap between Observed, Highest CATE, Highest if treated; approximately half of the candidates selected for Mentoring under Highest CATE are also selected in the observed allocation. Furthermore, almost two-thirds of selections under Highest if treated and Highest CATE overlap. Similarly, 100 out of 138 candidates overlap between Highest if trated and Observed selection. In contrast, there is low overlap between Least promising and the other three policies, with 124 candidates that are selected under Least promising not being selected under any other policy.
Figure 18: Venn diagram of selected policies.
Note: Each policy has 138 selected candidates. Percentages refer to the share of the 138 that fall into the region of the graph.
Next, we consider the following sets of candidates: (i) Observed selection – all candidates selected under the current allocation, (ii) Only least promising – candidates that are selected under the Least promising policy and are not selected under the remaining three policies, (iii) Highest CATE & Observed selection – the candidates that are selected both under the Highest CATE and under Observed selection, and (iv) Highest CATE but not selected - where we consider candidates that are selected under Highest CATE, but under mentors’ rule. We analyze the baseline outcomes, average treatment effects, and means of selected covariates. We present results in Table 11
Table 11: ATEs, baselines, and covariate means and standard errors for different policies
policies ATE Baseline STEM Prof. exp 30 plus Small town
Observed selection $0.095(0.012)$ $0.275(0.008)$ $0.457(0.011)$ $7.072(0.110)$ $0.514(0.168)$ $0.232(0.009)$
Only least promising $0.125(0.008)$ $0.048(0.005)$ $0.076(0.006)$ $6.277(0.160)$ $1.000(0.000)$ $0.588(0.012)$
Highest CATE & Observed selection $0.158(0.017)$ $0.244(0.011)$ $0.493(0.015)$ $8.587(0.161)$ $0.507(0.015)$ $0.400(0.015)$
Highest CATE but not selected $0.208(0.016)$ $0.112(0.011)$ $0.365(0.016)$ $10.095(0.168)$ $0.365(0.016)$ $0.873(0.011)$
Note: Average treatment effects and baselines from Section 6.2 standard errors clustered at the candidate level. STEM is the share of candidates selected under the considered policy that have a degree in STEM, Prof. exp is the average length of professional experience, 30 plus is the share of candidates that are above 30 years old, and Small town is the share of candidates that reside in a smaller town.
Comparing candidates selected only under Least promising to other categories, we find that these applicants are less likely to have a STEM degree, have shorter professional experience, and are more likely to be above 30 years old. Even though both mentors overlook this group and we do not select these candidates under the Highest CATE policy, the estimated treatment effects for this group are high. Furthermore, the treatment effect of 0.125 increase in the probability of getting a tech job is particularly high considering the baseline of only 0.048.
Finally, comparing the candidates that have high estimated treatment effects and get selected to those with high treatment effects who are not selected, we notice that first, mentors select high CATE candidates with higher baselines and lower average treatment effects. They tend not to select enough candidates who have longer professional experience and come from smaller towns. These patterns are consistent with our estimates of the heterogenous treatment effects shown in Section 5.2.

### v4

@@
-
+Susan Athey    Emil Palikot
 Abstract
-We describe the design, implementation, and evaluation of a low-cost (approximately \$15 per person) and scalable program, called Challenges, aimed at aiding women in Poland transition to technology-sector jobs.
-This program helps participants develop portfolios demonstrating job-relevant competencies.
-We conduct two independent evaluations, one of the Challenges program and the other of a traditional mentoring program – Mentoring – where experienced tech professionals work individually with mentees to support them in their job search.
-Exploiting the fact that both programs were oversubscribed, we randomized admissions and measured their impact on the probability of finding a job in the technology sector.
-We estimate that Mentoring increases the probability of finding a technology job within four months from 29% to 42% and Challenges from 20% to 29%, and the treatment effects do not attenuate over 12 months.
-Since both programs are capacity constrained in practice (only 28% of applicants can be accommodated), we evaluate the effectiveness of several alternative prioritization rules based on applicant characteristics.
-We find that a policy that selects applicants based on their predicted treatment effects increases the average treatment effect across the two programs to 22 percentage points.
-We further analyze how alternative prioritization rules compare to the selection that mentors used.
-We find that mentors selected applicants who were more likely to get a tech job even without participating in the program, and the treatment effect for applicants with similar characteristics to those selected by mentors is about half of the effect attainable when participants are prioritized optimally.
+Women remain underrepresented in technology despite many support programs, partly due to limited evidence on which interventions work and how to allocate scarce program slots.
+We evaluate two approaches to facilitating technology-sector transitions for women in Poland: Challenges, a scalable and cheap-to-operate (\$15/person) online program developing job-relevant portfolios (designed and implemented during this research), and one-on-one mentoring.
+The programs address distinct labor market barriers: expanding professional networks and teaching their effective use (Mentoring) versus providing credible skill signals (Challenges).
+Randomizing oversubscribed admissions, we find both programs substantially increase technology employment twelve months post-completion, Mentoring by 15 percentage points (p.p.) and Challenges by 11 p.p.
+Exploiting treatment effect heterogeneity, we show that algorithmic targeting based on predicted treatment effects achieves outcomes 86% larger than random assignment at current capacity levels.
+While mentors successfully identify high-benefit candidates (outcomes 48% above random selection), algorithmic allocation using observable characteristics yields an additional 11% improvement.
+The results demonstrate that (i) relatively low-cost interventions can meaningfully address gender gaps in technology, and (ii) data-driven allocation rules can substantially enhance program effectiveness even when expert judgment performs well.
+JEL: J24, J08, J16, C93
 1 Introduction
-An important concern of labor policymakers is identifying and implementing effective programs to support workers transitioning into occupations and industries that offer attractive jobs.
-One sector that has experienced rapid growth in recent decades and offers many attractive career opportunities is the technology sector.
-However, not all societal groups have benefited equally from opportunities in this sector; in particular, women remain underrepresented.
-Despite a number of programs and initiatives aimed at supporting women in the technology sector, [^1] For example, the United Nations organizes an annual event Girls in ICT Day; The European Commission carries out an annual study: Women in Digital Scoreboard.
-There are various NGOs dedicated to increasing women’s participation in the technology sector; For instance, Women in Technology International is a community of 160,000 women in the technology sector; Girls Who Code operates a network of 10,000 clubs providing support for women in technology.
-Finally, many major technology companies have initiatives aimed at supporting women in technology, including Women in Tech from Salesforce, The Women at Microsoft, or Accenture’s Women in TS&A.
-women held only 30% of technology jobs worldwide in 2021 (United Nations, 2021).
-The shortage of empirical evidence evaluating the effectiveness of specific strategies to address this issue is one of the hurdles in designing a successful policy response.
-In this paper, we ask whether interventions can have an impact on this problem, and if so, which workers benefit most.
-We conduct randomized evaluations of two distinct programs, both implemented in Poland.
-The first, Mentoring, involves one-to-one mentoring and is thus not easy to scale.
-We designed and introduced a new program, Challenges, to be scalable.
-In Challenges, women work in teams to create portfolios to show to employers.
-Second, we ask which women benefit most from each program.
-Using data from the two evaluations, we identify the distinct subgroups with the largest benefits.
-Third, we evaluate the effectiveness of alternative prioritization rules for admission to the program in the presence of capacity constraints, illustrating that there are large gains from the optimal prioritization.
-Fourth, we examine the implicit selection rule used by mentors, who each selected two candidates from a pool of applicants.
-We find that mentors do not select the candidates with highest estimated treatment effects, but rather focus more on candidates who would be likely to succeed even without the program.
-The results from our randomized evaluation suggest that the Challenges program increases the likelihood of successfully transitioning to the technology sector by approximately 45% within four months of graduation.
-[^2] Throughout this paper, we consider technology jobs as all jobs in firms in the technology sector other than those in finance, regulatory, legal, accounting, or human resources or jobs outside of the technology sector that involve software development and testing, IT support, and data analytics.
-Specifically, all jobs in UX and front-end programming are considered technology jobs.
-Challenges helps participants develop portfolios demonstrating capability relevant to the jobs they seek.
-The program was designed in partnership with technology companies, delivered online, and relies on interactions between program participants, which makes it cheap to operate (approximately \$15 per person) and suitable for scaling up.
-The program is also transferable across domains of the technology sector and languages; the first edition, evaluated in this paper, had specializations in user experience design and front-end programming.
-The program’s second edition was offered in Ukrainian instead of in Polish and offered additional specializations in software testing, cloud computing, and user interface design.
-We designed the Challenges program as a scalable alternative to traditional in-person mentoring.
-To develop it, we partnered with Dare IT, a Polish organization offering a non-profit mentoring program – Mentoring, since 2018.
-Dare IT’s mentoring program serves women who, despite having invested time into acquiring IT skills, still have not successfully obtained a technology job.
-There is a high demand for this program, as evidenced by the fact that it has been repeatedly oversubscribed.
-However, scaling it up to meet the demand faces several challenges.
-First, it is difficult to find enough mentors.
-Mentors operate only on a volunteer basis and tend to be women in mid-seniority technology jobs.
-Since few women are in these positions, recruiting more mentors is difficult.
-Second, the one-to-one format of Mentoring makes it operationally demanding.
-Ensuring satisfaction, remedying conflicts, and tracking everyone are complicated tasks that require regular supervision by Dare IT staff.
-The development of Challenges proceeded in two steps.
-First, together with Dare IT, we set up a series of interviews with HR experts and hiring managers of Polish technology companies to gain their perspectives on ways to support women candidates.
-Each of our interviewees mentioned the shortage of practical experience or clear signals of practical skills in women applicants’ profiles.
-Second, along with Dare IT and industry partners, we used the insights from the interviews to create Challenges.
-The program’s core consists of assignments sent every second week that, if completed, result in a job-relevant skill signal that can be added to participants’ resumes or mentioned during an interview.
-The program is designed based on three principles: (i.) it is delivered online, which lowers the costs; (ii.) it primarily involves interactions between participants, addressing the shortage of mentors; and (iii.) it is closely based on recent business cases.
-To operationalize the last principle, industry partners were recruited to help design the assignments and provide domain experts to review and critique selected submissions.
-We launched the registration to the program in December 2021 with 300 available spots; the program became oversubscribed within four hours.
-We conducted two concurrent randomized experiments to evaluate the effectiveness of Challenges and Mentoring.
-Both experiments exploit the fact that the two programs were oversubscribed, so we were able to randomize access and evaluate the impact of participation in a program on the probability of having a technology job.
-To evaluate Mentoring, we asked mentors to select two mentee candidates (instead of the usual one), and within these pairs, we randomized into treatment and control.
-In the case of Challenges, we randomly assigned access to 200 spots in the program.
-Four months after the end of each program, we collected data on the outcomes.
-We estimate that participation in Mentoring increased the probability of having a job in technology from 29% to 42%, while the Challenges program increased this probability from 20% to 28%.
-[^3] The difference in the outcomes of the two control groups is likely because, in Mentoring, mentors selected applicants which were later randomized into the control group.
-The control group of Challenges is a random draw from the pool of applicants.
-We document that the outcomes of non-selected applicants are substantially lower than those of the applicants selected by mentors and later randomized into the control group.
-We repeated the evaluation 12 months after the end of the two programs and found that their impact did not attenuate over time.
-We find that the two programs lead to different patterns of treatment effect heterogeneity.
-Mentoring is particularly beneficial for women from smaller towns that do not have a STEM degree, are mothers, and have been active in the labor market for over five years.
-Working with mentors with managerial experience and a longer tenure in technology also leads to high treatment effects.
-In the case of Challenges, participants who have just entered the labor market without a STEM degree exhibit high treatment effects.
-Different patterns of heterogeneity in treatment effects combined with insights from our discussions with Dare IT, mentors, and industry experts suggest that the two programs support participants’ employability through different mechanisms.
-Mentoring can positively impact the chances of obtaining employment in the technology sector by reinforcing self-efficacy, improving interview skills, or expanding professional networks.
-In contrast, creating portfolio items in the Challenges program that constitute signals of specific skills relevant to the job can counter a pessimistic belief (e.g., based on stereotypes) held by hiring managers, recruiters, or applicants themselves.
-The core motivation for creating the Challenges program was to extend Dare IT’s impact beyond what is possible via the relatively small Mentoring program.
-Although Challenges expanded the number of people who receive Dare IT services, the existing capacity constraints in Mentoring and Challenges present a significant hurdle, as the number of interested applicants greatly exceeds the number of available spots in each program.
-Furthermore, at the time this research was conducted, Dare IT was not using paid advertising to promote its services, so the applicants to Dare IT heard about the program through word-of-mouth recommendations or through an internet search, which underscores that the potential demand for Dare IT services is even higher.
-In the case of Mentoring, the constraint is caused by the limited number of mentors and the operational difficulties of scaling up the program.
-For the Challenges program, the number of spots has been intentionally limited to 300 to maintain appropriate participant support.
-[^4] The first 100 applicants were automatically granted a spot in Challenges and were removed from the analysis conducted in this paper.
-Even though there are plans to incrementally increase the number of available spots in future Challenges cohorts, Dare IT will continue to grapple with the decisions of which applicants to admit to the Challenges program and how to select participants for the Mentoring program in the foreseeable future.
-In the presence of capacity constraints and substantial heterogeneity in treatment effects, admission rules prioritizing applicants with the highest benefits can potentially enhance Dare IT’s impact.
-Adapting the methodology of Athey and Wager (2021) and Zhou et al.
-(2023), we estimate the value of counterfactual assignment policies that maximize outcomes across all applicants, given the constraints on the number of spots available in each program.
-These policies are then evaluated in the test set using doubly-robust estimates obtained through cross-fitting.
-We estimate that the existing policy, which assigns 15% of all applicants to Challenges and 13% in Mentoring, leads to the weighted average treatment effect of 12 percentage points increase in the probability of getting a tech job.
-Considering the case in which the available number of spots in each program does not change, but we change the assignment policy to prioritizes applicants with the highest predicted benefits from each program (we refer to this policy as optimal policy), we show that the average treatment effect increases to 22 percentage points.
-Thus, we find a 10 percentage points incremental benefit of targeting.
-Second, we evaluate the optimal policy at various capacity levels.
-While expanding capacity increases the average outcomes as applicants who previously did not participate are admitted, the incremental value of targeting over random admissions might decrease, as applicants with the highest predicted benefits are allocated first, and the marginal participants have lower predicted benefits.
-However, we estimate that targeting has substantial value even at high capacity levels.
-We show that when the Challenges program is expanded to 50% of all applicants, the incremental benefit of targeting is 5 percentage points, while when there are no capacity limits, the incremental benefit of targeting is to increase the probability of finding a job by 3 percentage points.
-The comparison of random and targeted admissions at the full capacity level, underscores that some participants have higher benefits from one program than from the other; thus, the value of targeting is not only in targeting between programs and out of programs, but also across the two offerings.
-Dare IT ensures high-quality matching by asking mentors to choose mentees from a list of pre-screened candidates.
-When comparing the outcomes of applicants to Mentoring, who were selected by mentors but randomized to the control group, with those not selected, we find that mentors tend to pick promising candidates.
-These candidates will likely find jobs in the technology sector, even without participating in the program.
-However, this approach to selecting participants might not be maximizing the potential impact of Mentoring.
-We use the off-policy approach to evaluate an alternative policy where mentors select candidates with the highest treatment effects.
-We find that the policy of choosing applicants based on their treatment effects leads to an average probability of getting a job in the technology sector of 23% across all applicants to Mentoring (we exclude applicants to Challenges from this analysis), compared to 20% when selecting the most promising candidates.
-This finding suggests that focusing on those most likely to benefit from the program rather than those who seem most promising overall could increase the program’s effectiveness.
-Finally, a non-parametric assignment policy which requires solving a discrete optimization program, might not be realistic in the context of Dare IT.
-Therefore, to create viable alternatives which lead to similar benefits as optimal policies but are easier to implement, we propose simple rules that define priority groups based on observed characteristics (e.g., applicants from smaller cities are prioritized for Mentoring).
-We then evaluate these rules using predicted treatment effects.
-We find that prioritization can substantially increase the two programs’ effectiveness.
+An important concern for labor policymakers is identifying and implementing effective programs to support workers transitioning into high-growth sectors.
+The technology sector exemplifies this challenge: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide.111unitednations2021 estimates the share of women in the technology sector at approximately 30%.
+Many programs exist to support women in technology, including NGO initiatives (Women in Technology International, Girls Who Code), corporate programs (Women in Tech from Salesforce, The Women at Microsoft), and government efforts (UN’s Girls in ICT Day, European Commission’s Women in Digital Scoreboard).
+While women face multiple barriers entering technology, this paper examines two specific labor market frictions.
+[^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (reuben2014gender; moss-racusin2012science), work-family conflict and child penalties (goldin2014grand; kleven2019children; bertrand2010dynamics), and underrepresentation in STEM education pipelines (bayer2016diversity; carrell2010sex).
+We focus on network and information barriers because short-term labor market interventions can directly address these constraints, unlike structural factors requiring longer-term policy responses.
+First, women have weaker professional networks than men, limiting access to job opportunities (greguletz2019women; forret2004networking).
+Second, gender stereotypes lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (cheryan2013stereotypical; del2022more; murciano2022missing).
+Understanding which constraints bind, and for whom, is critical for designing effective policy.
+If network barriers dominate, interventions should expand professional connections; if information frictions dominate, scalable credentialing programs may suffice.
+Alternatively, both frictions may bind simultaneously for different populations, necessitating multiple program types and optimal assignment rules.
+We provide experimental evidence distinguishing between these possibilities.
+We evaluate two interventions implemented in Poland by Dare IT, a social impact organization, each targeting one of these frictions.
+Mentoring pairs participants with senior female technology professionals for four months of one-on-one sessions.
+While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching effective relationship leverage.
+Challenges is a scalable online program (\$15 per participant) we designed during this research to address skill signaling barriers.
+The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, mobile app designs or React applications, that participants reference in job applications.
+[^3] Program development began with structured interviews with hiring managers at twelve Polish technology firms, who uniformly cited the absence of credible practical experience signals in female applicants’ profiles.
+We designed Challenges to be delivered entirely online, centered on peer interaction rather than scarce mentors, and based on realistic business cases to ensure employer credibility.
+The program has been replicated across multiple specializations (software testing, cloud computing, UI design) and delivered in Ukrainian for refugees.
+We exploit both programs’ substantial oversubscription to randomize access and measure employment outcomes via LinkedIn profiles over twelve months.
+[^4] Technology jobs include positions in technology firms (excluding finance, legal, HR, accounting roles) or technical positions elsewhere (software development, IT support, data analytics).
+UX and front-end programming roles are classified as technology jobs regardless of firm type.
+Both programs generate large employment effects, but with markedly different patterns of treatment effect heterogeneity that reveal which constraints bind for different populations.
+Mentoring increases technology employment by 15 percentage points (SE 5.7, from a 46% control baseline); Challenges by 11 percentage points (SE 4.8, from 27%).
+Evidence supports distinct mechanisms.
+Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
+The programs benefit different populations, with heterogeneity patterns aligning with each mechanism.
+Professional experience reveals the starkest contrast: Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers, while Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced workers.
+This indicates that early-career women benefit from portfolio signals that compensate for limited work history, while experienced women benefit from network expansion that leverages their accumulated professional standing.
+Geography provides additional evidence: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12).
+This pattern is consistent with network barriers binding tightly outside major urban centers where local technology ecosystems are sparse, while online portfolio signals function equally regardless of location.
+Mentor characteristics also predict treatment effects: participants matched to mentors with managerial experience or over five years in technology show larger treatment effects, consistent with senior mentors possessing richer networks and better access to hiring decision-makers.
+These differential patterns demonstrate that both network barriers and information frictions constrain technology employment, but for distinct populations defined by career stage or geography.
+The differential treatment effects create an allocation problem when program capacity is limited.
+Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratios for Mentoring; 300 spots filled in four hours for Challenges), requiring selection among applicants.
+We evaluate counterfactual assignment policies using the methods of athey2021policy and zhou2023offline, estimating which applicants should receive which program to maximize aggregate outcomes.
+A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves outcomes 86% larger than random assignment.
+This gain reflects two sources of value: selecting individuals with highest treatment effects and matching individuals to their optimal program type.
+The matching component generates persistent value: even when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment.
+This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion, treatment effect heterogeneity enables welfare-improving assignment even when most applicants can be served.
+We further compare algorithmic targeting to expert human judgment.
+Mentoring relies on experienced female technology professionals to select mentees from pre-screened applicants based on pre-recorded short videos and the assessment of the application materials (cover letter, resume, and registration survey).
+This expert selection substantially outperforms random allocation, achieving outcomes 48% larger.
+Yet algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain preference) achieves outcomes 63% larger than random assignment, an additional 11% improvement over mentor judgment.
+The two approaches, by experts and by the algorithm, achieve high-degree of concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates.
+Despite mentors’ rich soft information, systematic patterns in treatment response remain that simple demographic targeting can exploit.
+The result extends recent work showing algorithms can complement expert decisions in domains from bail (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
+The remainder of the paper proceeds as follows.
+Section 2 reviews related literature.
+Section 3 describes the programs, theoretical framework, and testable predictions.
+Section 4 details the experimental design and data.
+Section 5 presents treatment effect estimates, mechanisms, and heterogeneity analysis.
+Section 6 evaluates counterfactual allocation policies.
+Section 7 concludes.
 2 Literature Review
-This paper relates to several strands of literature.
-First, the paper contributes to the literature on the effectiveness of mentoring in improving career outcomes.
-Several influential papers focus on the effectiveness of mentoring in the academic context: Dennehy and Dasgupta (2017) employ a randomized experiment to show that assigning a female mentor to a female engineering student increases the sense of belonging in the discipline, self-efficacy, motivation, and retention for women majoring in engineering in college.
-Resnjanskij et al.
-(2021), also in a randomized experiment, document a mentoring program’s positive impact on disadvantaged high-school youth’s math grades and career plans.
-Several papers focus on the impact of mentoring on the careers of academic faculty members.
-Gardiner et al.
-(2007) provide empirical evidence that mentored individuals are more likely to stay at a university, receive more grant income, and receive promotions more frequently.
-In more recent studies, Ginther et al.
-(2020) and Blau et al.
-(2010) analyze the effectiveness of a mentorship workshop to support women in research careers.
-In a randomized access design similar to ours, they show that individuals in the treatment group are more likely to receive tenure at their institution, have more publications and citations, and receive more federal grants.
-We contribute to this literature by providing evidence from a randomized trial of the efficacy of mentoring outside of the context of education or academic jobs.
-[^5] The impact of mentoring on career outcomes has also been studied using surveys and other techniques based on observational data.
-For example, Ragins and Cotton (1999), in a survey of over 600 individuals, find that mentored respondents have better career outcomes than non-mentored respondents.
-They also find that the gender composition of the mentoring pair influences the effectiveness of mentoring.
-Similarly, Aryee et al.
-(1996) survey a group of workers in Hong Kong and found that mentored respondents reported a higher number of promotions and were more satisfied with their work than those not mentored.
-Kammeyer-Mueller and Judge (2008) carry out a meta-analysis of this literature and concludes that mentoring improves job and career satisfaction.
-Second, there is rich evidence evaluating the efficacy of offline labor market training programs.
-Card et al.
-(2018) carry out a comprehensive meta-analysis of the estimates from over 200 recent studies.
-They find that on average, the analyzed training programs have close to zero short-term effects on employment and some small positive impact in 2-3 years after completion.
-Programs evaluated in Card et al.
-(2018) generally aim at unemployed individuals looking for blue-collar jobs.
-[^6] Other influential papers evaluating labor training programs have a similar focus: Barnow (1987); Bloom et al.
-(1997); Heckman et al.
-(1999).
-In this category, programs focused on the development of skills on-the-job tend to be more effective (Sianesi, 2008; Lechner and Gerfin, 2000).
-On-the-job training programs do not necessarily need to take the form of long apprenticeships to be effective: Biewen et al.
-(2014) compare a long apprenticeship with short-term on-the-job training, combined with a job search workshop, and find that the short-term program is equally effective or more effective than the much longer program.
-Fein and Hamadyk (2018) use a randomized experiment to evaluate YearUp, a sectoral training program, and find a 40% increase in quarterly earnings in the medium term.
-We contribute to this literature by bringing evidence from a randomized trial of a training program that happens online, is focused on individuals joining a highly paid technology sector, and is very cheap to operate.
-Our program has on-the-job training elements as participants develop content designed by domain experts to closely resemble the content they would produce on the job.
-Third, a few studies have analyzed the efficacy of Massive Open Online Courses (MOOCs) in supporting learners’ job search.
-For instance, Zhenghao et al.
-(2015) shows that a significant majority of Coursera participants reported career benefits, particularly those from economically- and academically-disadvantaged backgrounds.
-However, through a difference-in-differences approach, Castaño-Muñoz and Rodrigues (2021) finds no effect on wages or employment among MOOC learners in Spain and a moderate impact on job retention.
-Hadavand et al.
-(2018) compare participants in a Data Science MOOC that were just above the threshold to graduate with those just below and argue that graduation from this MOOC leads to a 30% increase in the probability of finding a job.
-Our study contributes to this literature by using a randomized experiment approach.
-We also compare an online training program with an in-person Mentoring, identifying diverse patterns of heterogeneity that suggest different priority groups for online versus offline learning.
-Fourth, we contribute to the literature on skill signaling (Spence, 1978).
-Previous research showed that credentials demonstrating capability can significantly boost earnings Tyler et al.
-(2000); Hadavand et al.
-(2018).
-Furthermore, in situations where companies might be reticent about hiring less experienced candidates, establishing trustworthy indicators of ability can bring advantages at the market level Pallais (2014).
-Our research, through a randomized trial, sheds light on the efficacy of a program aimed at developing labor market signals.
-Our subjects are women seeking tech roles who can utilize skill signaling to counter possible negative stereotypes.
-However, since the programs we evaluate incorporate both learning and signaling, we cannot separate the signaling effect.
-Fifth, we document a high degree of heterogeneity in treatment effects from both programs, which we use to define treatment assignment policies.
-We show that targeting based on treatment effects leads to substantially higher average outcomes than random targeting, or targeting based on lowest outcomes in the absence of treatment, or highest outcomes with treatment.
-In doing so, we contribute to a small but growing literature on targeting treatment based on estimated treatment effects.
-An early contribution to this area was made by Ascarza (2018), who using randomized experiments, documents that treatment effect-based targeting is more effective in customer retention than churn-based prediction.
-Similar findings have been noted in marketing and development economics and medical contexts (Hitsch and Misra (2018), Devriendt et al.
-(2021), Haushofer et al.
-(2022), Inoue et al.
-(2023)).
-Lastly, a comparison of assignment policies for financial aid application nudges by Athey et al.
-(2023) highlights the superiority of maximizing treatment effects over targeting based on the predicted risk of non-compliance.
-Finally, this paper adds to the literature on the gender gap in the tech sector.
-Various reasons for low women’s participation in the technology sector have been studied.
-For example, Murciano-Goroff (2022) identify women’s tendency to under-report programming skills, impacting their job prospects.
-Strategies proposed to counter this include increasing role model visibility (Correll and Mackenzie (2016)) and shifting the perception of tech roles being predominantly male (Cheryan et al.
-(2013), Del Carpio and Guadalupe (2022)).
-We contribute to this literature by demonstrating how mentoring and job-training programs can enhance women’s success rates in obtaining tech jobs.
-Notably, our work focuses on the primary outcome of finding a job.
-3 Mentoring and Challenges Programs
-Over recent decades, Poland’s technology sector has witnessed significant growth, culminating in a 2021 revenue of roughly USD 51 billion, which accounts for 8% of the national GDP and provides employment for 250,000 individuals (GUS, 2022a).
-The sector’s growth has contributed to an escalating demand for skilled tech professionals, subsequently driving an approximate 9% wage growth in 2021.
-This led to the sector’s median monthly wage reaching USD 2,600, substantially outpacing the country’s overall median monthly wage for full-time workers, which stands at approximately USD 950 (GUS, 2022b).
-A survey of gender demographics within this industry unveils an underrepresentation of women, who make up 30% of the total Polish tech workforce (Gawlowska-Bujok and Bujok, 2021).
-The European Commission (European Commission, 2021) reports that the tech industry accounts for 1.1% of Polish women active in the labor market, compared to 5.2% of men.
-Dare IT is a social impact organization that aims at reducing the gender gap in the tech sector.
-They do so by helping qualified women transition from other careers.
-Dare IT offers a variety of services for women looking for jobs in tech or making their first steps in the profession, including a mentoring program, career fairs, networking events, and an online social network.
+This paper contributes to the literature on treatment effect heterogeneity and optimal policy assignment.
+Building on recent advances in policy learning using machine learning methods (athey2017econometrics; kitagawa2018should; athey2021policy), we examine whether data-driven targeting can improve program allocation decisions in active labor market policy.
+Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (kitagawa2018should; bhattacharya2012inferring).
+Empirical applications of these methods have demonstrated value across diverse domains: customer retention (ascarza2018retention), marketing (hitsch2018heterogeneous), development programs (haushofer2022targeting), agriculture (kasy2021adaptive), and medical treatment (inoue2023machine).
+We extend this literature to active labor market policy.
+Optimal allocation based on predicted treatment effects achieves outcomes 86% larger than random assignment at 15% capacity and 23% larger even at universal coverage.
+Crucially, targeting based on predicted treatment effects (CATE) dominates targeting based on baseline outcomes or predicted final outcomes, demonstrating that treatment response heterogeneity does not align with baseline heterogeneity.
+The persistent gains at near-universal coverage reflect complementarity between individual characteristics and program type, creating allocation value beyond capacity expansion.
+We further contribute to the growing literature examining whether algorithmic decision-making can complement or improve upon expert human judgment.
+Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), often by identifying systematic patterns that experts miss or cannot consistently apply.
+Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (pallais2014inefficient), we show that observable characteristics predict treatment effects systematically.
+We demonstrate that data-driven targeting can improve upon expert human judgment.
+Mentor-based selection achieves treatment effects 48% larger than random allocation.
+Yet algorithmic targeting using only six observable characteristics achieves an additional 11% improvement.
+The 70% concordance between methods indicates complementarity: algorithms largely validate expert choices while identifying additional high-benefit candidates, demonstrating that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity.
+This paper contributes to the literature on mentoring effectiveness in improving career outcomes.
+Several influential papers focus on mentoring in academic contexts: dennehy2017female; resnjanskij2021can; gardiner2007show; ginther2020can; blau2010can show that mentorship increases sense of belonging among female engineering students, youth’ learning outcomes, and tenure rates, publications, and federal grants for women in research careers.
+[^5] The impact of mentoring on career outcomes has also been studied using observational data.
+ragins1999mentor survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair.
+kammeyer2008quantitative conduct a meta-analysis concluding that mentoring improves job and career satisfaction.
+Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (granovetter1973strength; ioannides2004job; pallais2016referential).
+We provide randomized trial evidence on mentoring effectiveness in labor market transitions outside academic settings, demonstrating large effects on technology employment for women career-changers.
+Crucially, we provide experimental evidence that mentoring causally expands professional networks: Mentoring increases participants’ LinkedIn connections by 52.7 (SE 20.8), a 23% increase from baseline, while the non-network-based Challenges program shows no statistically significant effect.
+We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with observable characteristics of both mentors and mentees predicting treatment effects systematically.
+We contribute to the extensive literature evaluating active labor market training programs.
+card2018works conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion.
+[^6] See also heckman1999economics, card2010active, and crepon2016active for reviews of ALMP effectiveness.
+Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (sianesi2008differential; lechner2000microeconometric), though even intensive apprenticeships show mixed results (biewen2014effectiveness).
+fein2018bridging show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains.
+However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (abebe2021anonymity; bassi2022screening; carranza2020job; athey2024moocs).
+Building on the theoretical work of spence1978job, this literature demonstrates that credentials demonstrating capability can significantly boost earnings (tyler2000estimating; hadavand2018can) and generate market-level benefits when employers face uncertainty about worker quality (pallais2014inefficient).
+Recent evidence further suggests that training effectiveness may depend on complementary interventions: batista2022closing find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
+We provide experimental evidence on training program effectiveness in technology sector transitions, a context involving higher-skilled, career-changing women rather than the unemployed workers typically studied.
+We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in card2018works’ meta-analysis.
+We directly compare two program modalities: personalized one-on-one mentoring versus scalable online project-based training.
+This head-to-head comparison reveals complementarity rather than substitutability: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment.
+Our subjects are women seeking technology roles who can utilize skill signaling to counter gender-based statistical discrimination.
+Finally, this paper adds to the literature on the gender gap in the technology sector.
+The persistent gender wage gap reflects both occupational segregation and within-occupation differences (juhn2017specialization).
+Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (murciano2022missing) and the impact of occupational stereotypes (cheryan2013stereotypical; del2022more).
+Strategies proposed to counter these barriers include increasing role model visibility (correll2016succeed) and shifting perceptions of technology roles.
+We contribute to this literature by demonstrating how mentoring and credentialing programs can enhance women’s success rates in obtaining technology jobs.
+3 Experimental Interventions and Theoretical Framework
+Labor market transitions face two fundamental frictions: imperfect information about worker abilities and search costs in matching workers to opportunities.
+Workers address these through skill signaling (spence1978job) and professional networks (granovetter1973strength; ioannides2004job; pallais2016referential).
+For women entering male-dominated technology sectors, both mechanisms face gender-specific barriers.
+First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (cheryan2013stereotypical; del2022more; murciano2022missing).
+Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (greguletz2019women; forret2004networking).
+Third, women in male-dominated fields face psychological barriers that may reduce job search effectiveness.
+Women systematically under-report their technical abilities on resumes compared to similarly skilled men (murciano2022missing), use less self-promoting language in professional contexts (lerchenmueller2019gender), and may experience reduced confidence when confronting occupational stereotypes (spencer1999stereotype).
+We study the relative importance of these barriers by evaluating two interventions with distinct mechanisms: a mentoring program that addresses network gaps and builds confidence through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
+Context.
+Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (gus2022).
+Despite strong labor demand driving 9% annual wage growth, women held only 30% of technology positions and comprised just 1.1% of the female labor force, compared to 5.2% of men (EC2021; fluff2021).
+This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (unitednations2021).
+3.1 Program Designs and Mechanisms
+We partnered with Dare IT, a Polish social impact organization addressing gender barriers in technology employment, to evaluate two interventions targeting different constraints.
+Both programs serve women with baseline technical skills who have not yet secured technology jobs.
+The programs’ contrasting designs allow us to contrast network effects from signaling effects.
+Mentoring: Network Expansion and Tacit Knowledge.
+Dare IT launched its Mentoring program in 2018 to support women transitioning into technology careers.
+The program pairs participants with experienced female technology professionals for weekly one-on-one sessions over four months.
+Mentors, all women actively working in tech, volunteer their time and are selected based on technical expertise and professional standing.
+From its inception, demand has substantially exceeded capacity.
+The spring 2021 cohort received over 1,000 applications for 150 available spots, a ratio typical of recent cohorts.
+This persistent oversubscription reflects a fundamental scalability constraint: the program’s reliance on volunteer mentors limits expansion, as the pool of qualified female tech professionals willing and able to dedicate four months of weekly mentoring remains finite.
+Mentoring addresses gender-specific network barriers documented in prior research.
+Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (greguletz2019women).
+Female professionals may also hesitate to engage in networking behaviors that appear effective for men (forret2004networking).
+By connecting mentees to established female professionals, the program potentially overcomes both barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
+Our qualitative interviews with mentors reveal four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review.
+Mentors consistently emphasize networking as the primary mechanism, actively facilitating introductions and teaching mentees how to leverage professional connections effectively.
+Challenges: Skill Signaling.
+To overcome Mentoring’s capacity constraints, we designed Challenges as a scalable alternative targeting information frictions exacerbated by gender stereotypes.
+Program development began with structured interviews of hiring managers at twelve prominent Polish technology firms.
+These conversations revealed a systematic pattern: hiring managers perceived female applicants as lacking practical technical experience, even when formal credentials suggested otherwise.
+Managers emphasized that portfolio projects demonstrating hands-on capability could substantially alter hiring decisions, but expressed skepticism about generic online certifications.
+These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (murciano2022missing) and stereotypes about women’s technical competence create additional screening burdens for female candidates (cheryan2013stereotypical).
+A credible skill signal may be particularly valuable for women if it counters these negative stereotypes by providing concrete, verifiable evidence of technical proficiency.
+Challenges provides six bi-weekly assignments in User Experience Design (UX) or front-end programing in React.
+All assignments are designed by industry practitioners, that culminate in a complete portfolio project (UX design or React application; see Figure 1 for an example from the UX path).
+Each assignment replicates realistic business tasks and receives expert evaluation and feedback.
+Participants can add completed portfolios to resumes and reference them in interviews, creating a signal that hiring managers indicated they would find credible.
+The program operates online, accommodates cohorts of 300+ participants, and costs approximately \$15 per person.
+Figure 1: An Example of Challenges Final Product.
+Note: Selected wireframes from a design of a mobile app for pets adoption prepared during Challenges. See: https://www.behance.net/gallery/142047393/Pet-Adoption-App-UXUI for the entire design.
+A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
+3.2 Theoretical Predictions and Empirical Strategy.
+The programs’ differential mechanisms generate testable predictions about the treatment effect heterogeneity.
+The patterns of heterogeneity reveal which populations benefit most from network expansion versus skill signaling, providing evidence on the economic importance of each mechanism.
+If network barriers constrain technology employment for some women, Mentoring should be especially effective for participants with weak existing professional connections: women in smaller cities and towns where local technology networks are sparse, those without family or friends working in technology, and those with longer gaps in employment who may have lost previous professional contacts.
+The network mechanism also predicts effects on observable network measures: in particular, we should observe increases in LinkedIn connection counts among Mentoring participants.
+Additionally, if experienced mentors’ networks generate value, treatment effects should vary with mentor characteristics.
+Mentors with longer technology sector tenure and those in managerial positions likely possess richer professional networks and better access to hiring decision-makers, potentially generating larger effects for their mentees.
+If imperfect information about technical abilities constrains hiring, particularly when exacerbated by gender stereotypes, Challenges should disproportionately benefit women whose traditional credentials provide weak or ambiguous signals: for example, those early in their careers with limited work history.
+For these candidates, the portfolio complements weak existing signals.
+In contrast, women with strong observable credentials may gain less marginal value from an additional portfolio signal.
+We test these predictions by estimating treatment effects across participant subgroups defined by network strength, credential quality, and demographic characteristics.
+Differential patterns of heterogeneity provide evidence on whether these frictions meaningfully constrain employment and which populations each intervention serves most effectively.
+We further examine effects on LinkedIn connection counts to directly test the network expansion mechanism.
+4 Experimental Design and Data
 Mentoring.
-Dare IT launched Mentoring in 2018 as a not-for-profit initiative to support women aspiring to tech jobs.
-Since then, the program has been offered twice per year, and the number of participants has increased in each cohort.
-The program is funded by grants and donations, which allows participation to be free of charge.
-In recent years, demand for the program has been far greater than supply.
-For example, for the spring 2021 cohort, Dare IT received over 1,000 applications for 150 spots.
-The program follows several steps.
-Each cycle begins with Dare IT recruiting potential mentors through past participant referrals, social media engagement, and corporate partnerships.
-All mentors are women who are active professionals in the tech sector, and they contribute their time and expertise to the program on a pro-bono basis.
-Mentors are screened and selected to participate based on an application process; Dare IT considers multiple criteria, including candidates’ expertise in domains that attract the most interest from mentees, length of professional experience, and reputation.
-Simultaneously, the recruitment of potential mentees occurs primarily via social media and word-of-mouth recommendations.
-During this study, Dare IT did not utilize paid advertisements for recruitment.
-Dare IT selects participants based on several criteria.
-Only fully-completed applications by women not currently engaged in a tech role are considered.
-The program primarily targets individuals with necessary skills — quantified by self-reported learning hours in a particular domain and manual review of applications by Dare IT employees — who require additional support in their tech job search.
-Applicants who do not meet a baseline skill level are excluded.
-After Dare IT finalizes the selection process, mentors receive access to a database of eligible applicants.
-They can then select their preferred mentee.
-Once mentors and mentees are paired, they typically meet on a weekly basis for four months in order to support the mentee in their job search.
-Dare IT provides mentors with guidelines and best practices to increase the likelihood of a successful mentorship experience, although mentors have the flexibility to modify these guidelines to cater to their mentee’s needs.
-As a result, mentoring takes various forms, from coaching in job searching and interviewing or resume preparation to reviewing code and developing technical skills.
-Our discussions with mentors and Dare IT staff suggest four potential channels through which Mentoring increases the employability of mentees: (i.) reinforcement of self-efficacy through the provision of emotional support and encouragement, (ii.) help in the development of a professional network by relaying on connections of mentors [^7] Pallais and Sands (2016) emphasizes how referrals can augment job opportunities and delves into the underlying mechanisms., (iii.) explanation of unwritten norms of professional behavior and conduct, especially those intrinsic to the tech sector and, more specifically, during tech recruitment phases, and (iv.) direct transfer of skills through code review and discussions.
+We evaluate the Autumn 2021 cohort, for which recruitment opened in October 2021 and the program launched in January 2022.
+As in previous cohorts, applications substantially exceeded the number of available mentors.
+Even after Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking technical skills at a junior professional level, the applicant-to-mentor ratio remained 4:1.
+This oversubscription allowed us to implement a randomized evaluation while preserving Dare IT’s mentor-mentee matching process.
+We modified the standard selection procedure: instead of each mentor selecting one mentee (the historical practice), we asked mentors to select their two preferred candidates from the pool of eligible applicants.
+We then randomized within each pair, assigning one applicant to treatment (participation in Mentoring) and one to control (no participation).
+This paired randomization design controls for mentor selection preferences while generating exogenous variation in program access.
+Of 160 mentors, 151 agreed to this modified selection process; 9 declined, because they did not want to spend effort selecting mentee who will not end up working with them.
+This yielded an initial sample of 302 applicants (151 treatment, 151 control) across 151 mentor pairs.
+However, 4 treatment-assigned participants secured technology jobs between randomization (November 2021) and program launch (January 2022).
+We excluded these individuals from analysis and allowed the initially-assigned control applicant to participate instead, maintaining program access for mentors who invested time in candidate selection.
+Our final experimental sample comprises 151 treatment participants and 147 control participants.
+Appendix G provides robustness checks: (i.) excluding the 4 mentor pairs without matched controls, (ii.) including the 4 pre-program employment cases with outcomes set to 1, and (iii.) including the 9 non-participating mentors under the assumption their hypothetical mentees did not find technology jobs.
+Results are substantively unchanged across specifications.
+In addition to experimental subjects, we collected outcome data on 468 eligible applicants whom mentors did not select.
+These applicants provide a benchmark for assessing mentor selection effects independent of program participation.
+As shown in Figure 2, selected-but-control participants have substantially higher employment success than non-selected applicants, highlighting the importance of randomization for isolating treatment effects from selection effects.
 Challenges.
-Due to the excess demand for Mentoring, Dare IT explored the possibility of expanding its capacity.
-However, scaling proved to be a complex issue.
-First, selecting more mentors without sacrificing the quality of mentoring is difficult due to the scarcity of women occupying suitable roles in tech.
-Second, implementing a large 1:1 mentoring program presents operational difficulties.
-We partnered with Dare IT to develop a more easily scalable approach.
-Our first step was forming a set of hypotheses on barriers that impede women from securing roles in the Polish tech sector.
-We conducted interviews with hiring managers and HR experts from a dozen prominent tech companies in Poland.
-Each of our interviewees mentioned a shortage of practical experience or signals of practical skills on job application materials provided by female candidates.
-They emphasized that tangible proof of hands-on experience, even a short project, could significantly alter the interview dynamics, typically favoring candidates.
-They also expressed low trust in skill credentials that are available online.
-Many of our interviewees described online credentials as needing to be more informative about the capability to carry out real-world business tasks.
-Motivated by these findings, we developed a new program called Challenges, designed to address these specific concerns.
-The Challenges program comprises six bi-weekly mini-projects.
-Successful completion of all mini-projects results in a portfolio item, e.g., a UX design for a mobile application, which functions as proof of practical experience in the chosen domain (see Figure 1 for an example of wireframes from a completed product).
-The program strongly emphasizes on-the-job training: assignments are curated by domain experts and are readily applicable in a business setting.
-Each weekly mini-project undergoes evaluation and feedback from domain practitioners, including existing mentors and industry professionals, facilitating ongoing skill and knowledge enhancement for participants.
-A more comprehensive discussion of the way the Challenges program is organized and it’s content is in Appendix A.
-Figure 1: Challenges final product.
-Note: Selected wireframes from a design of a mobile app for pets adoption prepared during Challenges. See: https://www.behance.net/gallery/142047393/Pet-Adoption-App-UXUI for the entire design.
-We launched the program in December 2021, setting the start date for the initial cohort in mid-January 2022.
-Dare IT promoted the program on its social media channels to encourage applications.
-Enrollment was capped at 300 participants.
-[^8] The decision to limit program capacity to 300 spots was made to account for potential logistical challenges and to ensure that we could provide personalized assistance to all participants if necessary.
-As the program progressed, it became evident that only a few participants required hands-on support.
-Consequently, in subsequent cohorts, Dare IT increased the number of available spots.
-The first 100 slots were allocated to the initial 100 applicants (excluded from this analysis), with the remaining 200 spots randomly assigned among subsequent applicants.
-The program was oversubscribed within four hours, at which point we closed the registration form.
-Applications received post-closure were not included in the randomization.
-However, we collected data on these late applicants’ outcomes for comparison.
-4 Experimental design and data
-4.1 Experimental design
-Mentoring.
-We evaluate the mentoring program using the cohort of Autumn 2021.
-As in previous cohorts, Dare IT received a significantly higher number of applications than the available number of mentors.
-Even after the screening process to verify eligibility, the ratio of applicants to mentors remained at 4:1.
-The eligibility for the Mentoring program underwent manual verification by Dare IT staff.
-Only applicants who did not currently hold positions in the technology field and who, according to Dare IT’s reviewer assessments, have technical skills at a level typical of a junior employee in a tech company were considered eligible for participation.
-The surplus of eligible applicants compared to the number of available mentors provided us with an opportunity to implement a randomized access experiment.
-In order to maintain high-quality mentor-mentee pairings while incorporating randomized access, we asked mentors to select two potential mentees instead of just one, as they did in the previous program’s cohorts.
-Within these pairs, we then randomly determined which of the two applicants would be granted access to the program (the treatment group) and which would not (the control group).
-Thus, we created a paired experiment.
-In the experiment, we initially had 156 mentor-mentee pairs and an equivalent number of mentor-control pairs.
-However, four pairs did not start the program because the mentee secured a tech job between applying for the program and its commencement.
-In these cases, the mentor-selected applicant, who was initially randomized into the control group, participated in the program.
-[^9] We decided against having idle mentors to ensure broader access to the program, as well as, to respect mentors’ time dedicated to selecting candidates and willingness to volunteer their time in mentoring.
-Consequently, we excluded these mentors and participants from our analysis.
-As a result, our analysis included a total of 152 pairs.
-Additionally, we gathered data on 468 eligible applicants whom mentors did not select for participation.
-Challenges.
-Regarding the Challenges program, we concluded the recruitment process once we received 500 applications that met the established eligibility criteria.
-To qualify for participation, applicants were required to declare that they were not currently employed in the technology sector, possess a fluency level in English (B1 level), and be proficient in Figma or React depending on the specialization.
-The first 100 applicants were granted program spots, while the remaining 400 were randomly assigned to either the treatment or control groups.
-Randomization was conducted with stratification based on variables obtained from the registration survey.
-[^10] These stratification variables included specialization (front-end development or UX), age, type of residence, and employment status (student, employed, neither employed nor student).
-Additionally, we gathered data on 160 applicants who submitted their forms after the registration process had concluded.
-These late applicants were not included in the primary program evaluation.
-However, they are considered in the off-policy evaluation, as detailed in Section 6.
-Furthermore, it’s worth noting that some applicants randomized into the treatment group did not actively participate in the program (did not submit any assignments).
-Nevertheless, since these individuals still received all program materials and had the opportunity to engage in webinars discussing projects, we have included all such individuals in our analysis.
+We launched recruitment in December 2021, with the program scheduled to begin in February 2022.
+To generate initial interest, we guaranteed admission to the first 100 applicants.
+Registration was capped at 500 applicants, but high demand led to 525 applications before we closed the portal.
+Applicants self-certified that they (i.) were not currently employed in the technology sector, (ii.) possessed English language proficiency at B1 level or higher, and (iii.) had working knowledge of either Figma (User Experience Design track) or React (Front-end Development track).
+Unlike Mentoring, we did not manually verify technical skills prior to admission due to scalability considerations.
+We exclude the first 100 guaranteed-admission applicants from our analysis, as they were not subject to randomization.
+Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access).
+Randomization employed stratification on four registration characteristics: track specialization (UX Design vs.
+Front-end Development), age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
+Between randomization (December 2021) and program launch (February 2022), 17 treatment-assigned applicants secured technology employment.
+Consistent with our approach in Mentoring, we exclude these individuals from our primary analysis.
+Our final analytic sample includes 183 treatment participants and 225 control participants.
+Appendix G demonstrates that results are robust to including these 17 early-employment cases with outcomes coded as 1.
+Among the 183 treatment participants, 41 (22%) did not submit any program assignments, suggesting passive non-compliance.
+However, these individuals retained access to all program materials, weekly webinars with industry experts, and the Dare IT Slack community.
+Following intention-to-treat principles, we include them in our analysis regardless of active participation level.
+We also collected outcome data on 160 applicants who submitted forms after enrollment closed.
+These late applicants were not randomized and do not contribute to our primary treatment effect estimates.
+However, they provide an additional comparison group for the off-policy evaluation in Section 6.
+Appendix B reports covariate balance tests across treatment and control groups.
+Of 9 baseline characteristics, only social science degree shows a statistically significant imbalance at the 5% level, consistent with chance given multiple comparisons.
+Our main specifications control for all observed variables to improve precision.
 4.1 Data
-We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted following the program’s conclusion, and information collected from participants’ LinkedIn profiles.
+We have four data sources: applicant registration surveys, mentor registration surveys, outcome surveys conducted four months post-program, and LinkedIn profile data.
 Applicant Registration Survey.
-Applicants in both programs were asked to complete a simple registration survey.
-Data from the survey included the city of residence, age, employment situation, and time spent, which measures how much time the applicant had spent developing their IT skills.
-The mentoring application form was richer; first, applicants were asked if they have family or friends in IT and whether they have children; second, the form also had several open-ended questions related to the applicant’s motivation and the possibility of attaching a short introductory video, but we do not use either in this analysis.
-Additionally, we collected information on participants’ preferred domains: the main categories were UX design, manager (project or product), or developer (front or back-end).
+Both programs collected basic information during registration: city of residence, age, employment situation, and time spent developing IT skills.
+The Mentoring application additionally asked whether applicants have family or friends in IT and whether they have children.
+We also collected information on preferred domains: UX design, manager (project or product), or developer (front or back-end).
 Mentors Registration Survey.
-Mentors were asked a number of questions before the start of the program, which allowed us to construct the following variables: first-time mentor, career changer (which indicates whether the mentor herself had worked outside of IT before moving to the sector), managerial experience (which indicates whether the mentor has any managerial experience or not), and years of experience (years of experience in the tech sector).
-Outcomes survey.
-We surveyed participants of each program and those in control groups four months after the end of the program.
-In the survey, we were primarily interested in the salaries in the new jobs: we asked the question whether the respondent found a new job and in which range the salary was.
-We also inquired about the number of job offers and whether the participant negotiated the proposed offer or accepted the initial one.
-Finally, we carried out a shorter survey right after the end of Mentoring, the “short survey,” where we asked for salaries for the new jobs.
-Altogether we conducted three surveys: the main mentoring survey in which we received 162 responses (with 111 responded from treatment group subjects), the initial survey of the mentoring program (with 194 responses), and the Challenges survey, for which we received responses from 68 subjects in the control group and 61 from the treatment group.
-We analyze the outcomes from surveys in Appendix F.
+Mentors provided information enabling us to construct: first-time mentor (binary), career changer (whether the mentor worked outside tech before transitioning to the sector), managerial experience (binary), and years of experience in the tech sector.
 LinkedIn Data.
-Our primary source of outcome data is the LinkedIn profiles of the participants.
-This data was collected at two different points: first, four months, and second, 12 months after the conclusion of the programs.
-During the initial data collection phase, outcomes for all experimental subjects were gathered.
-In the second phase, Dare IT was unable to observe the outcome of one Mentoring participant, 17 Challenges participants, and 7 Challenges control group members.
-The principal cause of the missing data is the change in the LinkedIn profile address.
-[^11] It’s plausible that the propensity to change LinkedIn’s profile address was impacted by participation in the program.
-In Appendix ate_bounds we provide robustness checks to our primary results by assuming the most pessimistic and optimistic outcomes for the missing subjects.
-We also estimate AT bounds following the methodology in Lee and Weidner (2021).
-The primary outcome metric was whether a participant secured a technology job following a program’s start.
-As new technology jobs, we consider jobs with a starting date of January 2022 or later for Mentoring and February 2022 or later for Challenges that fall into one of the following two categories:
+Our primary outcome measure, pre-registered prior to data collection, is whether participants secured technology-sector employment following program completion.
+We collected employment data from participants’ LinkedIn profiles at two time points: four months post-program (May 2022 for Mentoring, June 2022 for Challenges) and twelve months post-program (May 2023 and June 2023, respectively).
+LinkedIn provides standardized job title metadata for each employment entry.
+The research team and research assistants, blinded to treatment assignment, coded each profile according to our pre-registered definition.
+A technology job is:
   1. 1.
-    All jobs in technology companies other than positions in finance, regulatory, legal, accounting, and HR, where technology companies include firms in software development, testing, and sales; data analytics; IT services; digital marketing; and online platforms (including peer-to-peer platforms, and online shops).
+    Any position in a technology company—firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms—excluding finance, regulatory, legal, accounting, and HR roles.
   2. 2.
-    Jobs in non-technology companies that involve software development and testing, IT support, and data analytics.
-In our context, this category includes jobs in banks and management consulting agencies.
-Based on these categories, we define three outcomes: new job, tech job 4 months, and tech job 12 months.
-New job refers to any job, whether in technology or not, that started between January and May 2022 for Mentoring and between February and June for Challenges.
-Tech job 4 months denotes a job that aligns with category (1) or (2) and is also classified as a new job.
-Finally, tech job 12 months is a job, corresponding to either category (1) or (2), that began after January 2022 for Mentoring and February for Challenges and before May 2023.
-We also collect the date on which subjects got their new jobs.
-In some cases, we do not observe the month in which the new job started, but only the year; for such observations, we assign the start date as December 2022 or May 2023, depending on whether the job started in 2022 or 2023.
-We also leveraged LinkedIn to gather data on the highest education level attained by participants, which was organized into the following categories: bachelor’s degree, master’s degree, postgraduate degree, non-traditional credential, and other.
-The other category includes profiles lacking education information or information that could not be classified into any of the pre-existing categories.
-Furthermore, we compiled data on college majors, forming two binary variables: STEM and social science or business.
-The former includes all STEM degrees, excluding social sciences, while the latter comprises degrees in social science, business, finance, or management.
-Lastly, we deduced a person’s age from a combination of age reported in the registration survey and the commencement date of their latest educational degree.
-Finally, we also compute their professional experience length by subtracting the start date of their first job recorded on LinkedIn from 2022.
-Table 1 summarizes the key variables for both programs.
+    Technical positions in non-technology companies involving software development and testing, IT support, or data analytics (e.g., software engineers at banks, data analysts at consulting firms).
+Self-employment, freelancing, and volunteer positions were not coded as new employment.
+Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference (p=0.9), supporting their equivalence.
+Data collection and attrition.
+We obtained complete LinkedIn data for all Mentoring participants at both measurement points.
+For Challenges, we collected data on all participants at four months and 384 of 408 participants (94%) at twelve months.
+Missing data resulted from participants changing LinkedIn profile URLs or privacy settings between recruitment and follow-up waves.
+Appendix G provides bounds and estimates under extreme attrition assumptions; main results remain statistically significant under all scenarios.
+Outcome variables.
+We define three binary measures: new job (any employment starting January 2022 or later for Mentoring, February 2022 or later for Challenges), tech job (technology-sector position per above criteria), and non-tech job (new employment outside technology).
+In a few cases, subjects reported only the start year of new employment on their LinkedIn profiles, we conservatively assign December 2022 or May 2023 as the start date.
+We also collected from LinkedIn: highest education level (bachelor’s, master’s, postgraduate, non-traditional credential, or other), college major (constructing binary variables for STEM and social science/business), age (combining registration survey data with degree completion dates), and professional experience length (from first job start date to 2022).
+Table 1 summarizes key variables.
 Table 1: Summary Statistics of the Main Variables: Treatment, Control, and Applicants
@@
 Mentoring: mentor characteristics
-First time mentor 152 0.467 0.501 0 1 145 0.434 0.497 0 1
-Managerial experience 152 0.520 0.501 0 1 145 0.510 0.502 0 1
-Career changer 152 0.691 0.464 0 1 145 0.710 0.455 0. 1
-Years of tech experience 152 5.856 3.777 2 35 145 6.021 3.794 2 35
+First time mentor 151 0.47 0.50 0 1 145 0.43 0.50 0 1
+Managerial experience 151 0.52 0.50 0 1 145 0.51 0.50 0 1
+Career changer 151 5.86 3.79 2 35 145 6.02 3.79 2 35
+Years of tech experience 151 0.70 0.46 0 1 145 0.71 0.46 0 1
 Mentoring: experimental groups
-New job 152 0.467 0.501 0 1 147 0.422 0.496 0 1
-Tech job 4 months 152 0.421 0.495 0 1 147 0.293 0.456 0 1
-Tech job 12 months 151 0.609 0.499 0 1 147 0.456 0.500 0 1
-Social science 152 0.132 0.339 0 1 147 0.136 0.344 0 1
-STEM 152 0.487 0.501 0 1 147 0.531 0.501 0 1
-Family friends IT 152 0.697 0.461 0 1 147 0.701 0.460 0 1
-Warsaw 152 0.276 0.449 0 1 147 0.259 0.439 0 1
-Mother 152 0.257 0.438 0 1 147 0.259 0.439 0 1
-Applied before 152 0.243 0.431 0 1 147 0.231 0.423 0 1
-Years of professional experience 152 7.053 5.049 0 31 147 7.204 4.816 0 23
+New job 151 0.65 0.48 0 1 147 0.57 0.50 0 1
+Non-tech job 151 0.04 0.20 0 1 147 0.12 0.32 0 1
+Tech job 151 0.61 0.49 0 1 147 0.46 0.50 0 1
+LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
+STEM 151 0.49 0.50 0 1 147 0.53 0.50 0 1
+Student 151 0.14 0.35 0 1 147 0.16 0.37 0 1
+Over 30 151 0.47 0.50 0 1 147 0.50 0.50 0 1
+Small town 151 0.25 0.44 0 1 147 0.24 0.43 0 1
+Applied before 151 0.24 0.43 0 1 147 0.23 0.42 0 1
+Mother 151 0.26 0.44 0 1 147 0.26 0.44 0 1
+High proficiency 151 0.54 0.50 0 1 147 0.57 0.50 0 1
+Family or friends in IT 151 0.70 0.46 0 1 147 0.70 0.46 0 1
+UX 151 0.42 0.50 0 1 147 0.44 0.50 0 1
+Masters 151 0.48 0.50 0 1 147 0.41 0.49 0 1
+Years of professional experience 151 6.99 5.01 0 31 147 7.20 4.82 0 23
 Challenges: experimental groups
-New job 183 0.333 0.473 0 1 225 0.293 0.456 0 1
-Tech job 4 months 183 0.284 0.452 0 1 225 0.196 0.398 0 1
-Tech job 12 months 166 0.380 0.487 0 1 218 0.266 0.443 0 1
-Social science 183 0.208 0.407 0 1 225 0.267 0.443 0 1
-STEM 183 0.421 0.495 0 1 225 0.391 0.489 0 1
-Years of professional experience 183 7.645 4.976 0 31 225 7.671 6.237 0 31
+New job 166 0.57 0.50 0 1 218 0.50 0.50 0 1
+Non tech job 166 0.19 0.40 0 1 218 0.23 0.42 0 1
+Tech job 166 0.38 0.49 0 1 218 0.27 0.44 0 1
+LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
+STEM 183 0.42 0.50 0 1 225 0.39 0.49 0 1
+Social science 183 0.21 0.41 0 1 225 0.27 0.44 0 1
+Warsaw 183 0.20 0.40 0 1 225 0.20 0.40 0 1
+Years of professional experience 183 7.64 4.98 0 31 225 7.67 6.24 0 49
+UX 183 0.49 0.50 0 1 225 0.54 0.50 0 1
+Over 30 183 0.47 0.50 0 1 225 0.44 0.50 0 1
+Small town 183 0.61 0.49 0 1 225 0.63 0.48 0 1
 Applicants both programs
-Tech job 4 months - - - - - 628 0.152 0.359 0 1
-Tech job 12 months - - - - - 628 0.279 0.449 0 1
-STEM - - - - - 628 0.196 0.367 0 1
-Years of professional experience - - - - - 628 7.474 6.318 0 37
+Tech job 12 months - - - - - 628 0.28 0.45 0 1
+STEM - - - - - 628 0.20 0.40 0 1
+Warsaw - - - - - 628 0.31 0.46 0 1
+Small town - - - - - 628 0.39 0.49 0 1
+Years of professional experience - - - - - 606 7.47 6.32 0 37
+Over 30 - - - - - 628 0.14 0.35 0 1
+Masters degree - - - - - 628 0.31 0.46 0 1
+Mentoring applicant - - - - - 628 0.75 0.44 0 1
 Note: Summary statistics of selected variables.
@@
 Appendix B shows the covariate balance between treatment and control.
+Outcomes Survey.
+We surveyed participants and control group members four months post-program, asking about new employment, salary ranges, number of job offers, and negotiation behavior.
+We conducted three surveys: the main Mentoring survey (162 responses, 111 from treatment), an initial Mentoring survey (194 responses), and the Challenges survey (68 control, 61 treatment).
+Appendix E analyzes survey outcomes.
 5 Effectiveness of Dare IT Programs
@@
 5.1 Average Treatment Effects
-In Figure 2, we display the proportions of participants employed in technology jobs over the span of 16 months, post-application to the program.
+In Figure 2, we display the proportions of participants employed in technology jobs over the span of thirteen months post-application to the program.
 This is illustrated across all experimental groups.
@@
 For those observations, we assign December of that year as the starting month.
-Consequently, the average outcomes at the 4th month after the end of either program are lower than the average of tech job 4 months across the experimental groups.
+across the experimental groups.
 In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT).
@@
 In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points.
-The difference in average outcomes across the three groups appears to be substantial, particularly in the case of Mentoring.
-However, the subjects in the Mentoring group were selected by mentors; thus, the program’s impact on the probability of getting a tech job might be compounded by the influence of the non-random selection of the sample.
-In the right figure, we decompose the effect of selection from the impact of Mentoring.
-We show the outcomes across applicants selected by mentors and randomized into Mentoring, selected, but randomized into control, and those that were not selected (the non-selected group consists of the non-selected applicants to Mentoring, and the control group, and late applicants to Challenges).
-We observe that in the 16th month, the probability of having a job in technology differs by over 20 percentage points between selected and not-treated and non-selected applicants, and it is substantially larger than the difference between the two groups that were selected but randomized into treatment or control.
-On average, mentors selected candidates that were more likely to get a job in technology.
-This observation highlights that, in this context, the ability to carry out a randomized trial is crucial for obtaining reliable estimates of the program’s effectiveness.
-Figure 2: Mean outcome per group over time.
+This differential timing, with Mentoring effects emerging during the program while Challenges effects manifest several months post-completion, suggests distinct mechanisms.
+These trajectories are consistent with the distinction between immediate network-based job matching versus the creation of portfolio signals and gradual employer response to them.
+The right panel decomposes selection from treatment effects in Mentoring.
+Applicants selected by mentors but randomized to control achieve substantially higher technology employment (43% at month thirteen) than non-selected applicants (19%), confirming that mentors identify promising candidates.
+Furthermore, the treatment-control gap among selected applicants is smaller than the selection effect itself, underscoring the importance of randomization for isolating program impacts from mentor selection.
+Figure 2: Share of Applicants with a Tech Job per Group Over Time.
 Note: Share of subjects with a tech job per group across time.
@@
 "Selected and treated" group (in green) shows the outcomes of the treatment group in Mentoring, "Selected but not treated" group (in blue) is the control group of Mentoring "Not selected" group (in red) includes applicants to Mentoring that were not selected by mentors, both experimental groups of Challenges, and late applicants to Challenges.
-In Table 2, we present the estimates of the average treatment effects for both the Mentoring and Challenges programs.
-Our analysis covers both short-term and longer-term outcomes: securing a job in tech within 4 months and 12 months after the programs, respectively.
-[^14] Tech job 4 months refers to 4 months after the end of a program which corresponds to approximately 8 months after application; analogously for Tech job 12 months.
-Additionally, we assess the influence of the programs on the acquisition of any new job within 4 months following the end of a program.
-For the short-term outcomes, we employ the difference-in-means estimator.
-[^15] Given the paired design of the mentoring experiment, we estimate variance as $\hat{V}(\hat{\tau}^{M})=\frac{1}{J\times(J-1)}\sum_{g=1}^{J}(\hat{\tau}^{M}_{j}-\hat{\tau}^{M})^{2}.$ (2) Athey and Imbens (2017) show that this is a conservative estimator of variance in a paired experiment.
-In the case of Challenges, we use the Splawa-Neyman estimator of variance.
-When estimating the longer-term effect, we utilize the fact that we observe monthly outcomes to estimate a Cox proportional hazard model (Cox, 1972).
-In Appendix C, we provide an estimate for short-term outcomes using the Augmented Inverse Propensity Weighting estimator (Robins et al., 1994; Athey et al., 2019).
-Furthermore, Appendix D contains alternative specifications of the Cox proportional hazard model, and the results align with those shown in Table 2.
-The test results of proportionality of hazard rates can be found in Appendix G.
-Table 2: Average treatment effect estimates.
+Table 2 presents estimated average treatment effects.
+We estimate the averate treatment effects using the difference-in-means estimator (columns 1-3, 5-7) and using Cox proportional hazard models that exploit monthly outcome variation (columns 4, 8).
+Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns in Figure 2.
+Both programs generate substantial, statistically significant increases in technology employment.
+Mentoring increases the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group baseline of 45.6%.
+Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the 26.6% control baseline.
+Both estimates use difference-in-means; results are nearly identical using Cox proportional hazard models that exploit monthly variation in job start dates (14.5pp for Mentoring, 11.2pp for Challenges).
+Table 2: Average Treatment Effects of Mentoring and Challenges.
+Difference-in-means Prop. hazard Difference-in-means Prop. hazard
 Mentoring Challenges
-Difference-in-means Proportional hazard Difference-in-means Proportional hazard
-New job Tech job 4 months Tech job 12 months New job Tech job 4 months Tech job 12 months
-ATE 0.045 0.129 0.126 0.040 0.089 0.071
-(0.05) (0.05) (0.025) (0.05) (0.04) (0.02)
-ATE % baseline 10.749 43.941 36.887 13.636 45.306 46.972
-(14.23) (17.09) (7.34) (12.00) (14.07) (10.01)
-No. of obs. treatment 152 152 2265 183 183 2490
-No. of obs. control 147 147 2205 225 225 3330
-Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology within 4 months, and within 12 months.
-The first row shows estimates of the average treatment effect using the difference-in-means estimator in the case of new job and tech job 4 months (see Appendix C for AIPW estimates).
-The average treatment effect on Tech job 12 months is estimated using Cox proportional hazard model (we present estimates from a model with no covariates, see Appendix D for Cox proportional models with covariates and with time-varrying treatment effects).
-The second row presents the estimates from the first row divided by the mean outcome in the control group.
-Rows three and four present the number of observations per experimental group.
-The first, second, and third columns present the results of Mentoring intervention.
-Columns four, five, and six show the results from the Challenges program.
+New job Tech job Non-tech job Tech job Cox New job Tech job Non-tech job Tech job Cox
+ATE 0.078 0.153*** -0.076** 0.145*** 0.072 0.113** -0.041 0.112**
+SE (0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
+ATE % baseline 13.576 33.676** -65.641** 31.718*** 14.458 42.646** -17.600 42.232**
+SE (9.893) (14.783) (26.720) (11.992) (10.268) (18.131) (17.978) (19.301)
+No. obs. treated 151 151 151 2265 166 166 166 2656
+No. obs. control 147 147 147 2205 218 218 218 3488
+Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology, and a new job not in technology.
+Columns 1-3 and 5-7 present estimates using the difference-in-means estimator and columns 4 and 8 using a proportional hazard model with a full set of covariates.
+The first row shows estimates of the average treatment effect in the difference in the probability of reporting employment at month 15.
+The third row presents the estimates from the first row divided by the mean outcome in the control group.
 Standard errors are presented below the average treatment effect estimates.
-The variance of the difference-in-means estimates of the average treatment effect of Mentoring are estimated using Equation 2.
 The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator.
-Standard errors of the estimates from Cox proportional hazard model are estimated using the partial likelihood function (coxph implementation in R).
-We find that participation in the Mentoring program significantly increases the probability of getting a job in the tech industry, by 13 percentage points within both 4 and 12 months after the program’s conclusion.
-The short-term effect’s estimate has a standard error of 5 percentage points, whereas the longer-term effect’s standard error is reduced to 2.5 percentage points, thanks to the proportional hazard model’s functional form assumption and our time-series data.
-The Challenges program also significantly enhances the likelihood of securing a tech job, although its impact is somewhat less pronounced than that of the Mentoring program.
-We estimate that participation in Challenges increases the chance of securing a tech job by 8.9 percentage points within 4 months (with a standard error of 4), and by 7.1 percentage points within 12 months (with a standard error of 2).
-In the second row of Table 2, we express the average treatment effects as a percentage of the control group’s mean outcome.
-As previously noted, mentor selection leads to higher average outcomes in the Mentoring control group compared to the Challenges control group.
-As a result, when expressed in percentage terms, the impact of both programs appears quite similar.
-Participation in Mentoring increases the odds of securing a tech job within 12 months by 37%, whereas in Challenges by 47%.
-Importantly, we find no evidence of the treatment effect diminishing over time when comparing the programs’ effectiveness 4 and 12 months after their conclusion.
-Finally, columns 1 and 4 of Table 2 display the estimated impact of participating in Dare IT programs on the probability of acquiring any new job within 4 months post-graduation.
-We do not observe any statistically significant difference in this outcome between treatment and control.
-Therefore, our results suggest that the treatment and control groups of both experiments managed to secure new jobs at similar rates.
-However, those in the treatment groups were more likely to find employment in the technology sector.
-In Appendix F, we show the results based on the outcomes survey.
+*** p<0.01, ** p<0.05, * p<0.1.
+Job Search Selectivity.
+Estimates in Column 3 of Table 2 suggest that Mentoring operates partly through increased job search selectivity.
+While Mentoring increases overall employment by only 7.8 percentage points (SE 5.7), statistically insignificant, it increases technology employment by 15.3 percentage points and decreases non-technology employment by 7.6 percentage points (SE 3.1).
+This implies that approximately half of Mentoring’s impact reflects participants becoming more selective, holding out for technology positions rather than accepting readily available employment in other sectors.
+In contrast, Challenges shows no significant effect on non-technology employment (-4.1pp, SE 4.2), consistent with this program operating through skill signaling rather than changing search behavior or persistence.
+The differential patterns support distinct mechanisms: Mentoring increases both participants’ capability to access technology jobs (through networks and tacit knowledge) and their willingness to persist in pursuing them (for example, through confidence-building), while Challenges primarily improves application success conditional on applying.
+This finding provides evidence consistent with the self-efficacy channel hypothesized in Section 3: mentors help women maintain focus on technology careers.
+In Appendix E, we show the results based on the outcomes survey.
 We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group.
 The differences between treatment and control are greater in the case of Mentoring than Challenges.
+5.2 Impact of Program Participation on the Number of LinkedIn Connections
+We directly test the network mechanism by examining whether program participation increases professional network size, measured by LinkedIn connection counts.
+This analysis distinguishes network expansion from other channels: if Mentoring works solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges.
+Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions or employer contact, should not.
+The data support the network mechanism.
+Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group baseline of 225, a 23% increase.
+In contrast, Challenges produces a statistically insignificant reduction of 11.4 connections (SE 17.8) from a baseline of 232.
+The differential treatment effect of 64 connections between programs (SE 28) provides direct evidence that Mentoring expands professional networks while Challenges does not.
+To understand where these effects appear in the connection distribution, Table 3 presents quantile regression estimates.
+Mentoring generates large, statistically significant increases at the 25th percentile (62 connections, SE 27) and median (92 connections, SE 35), but no detectable effect at the 75th percentile.
+This pattern indicates the program particularly benefits participants with initially sparse networks.
+In contrast, Challenges shows no significant effects at any quantile, confirming this program does not operate through networking channels.
+Table 3: Treatment Effects on LinkedIn Professional Networks.
+Quantile Treatment Effect Mentoring Treatment Effect Challenges Difference: Mentoring and Challenges
+0.25 62** 5 57*
+(27.159) (13.391) (30.281)
+0.5 92*** 5 87*
+(35.426) (29.470) (46.082)
+0.75 5 -8 13
+(47.004) (50.706) (69.141)
+Note: The table provides the quantile regression results for the Challenges and Mentoring programs.
+The treatment effects of the experimental groups on LinkedIn connections are shown at the 25th, 50th, and 75th quantiles.
+The difference column captures the difference in the treatment effects between the Mentoring and Challenges programs.
+Estimates are obtained by regressing the treatment indicator on the number of connections.
+Standard errors are in parentheses.
+*** p<0.01, ** p<0.05, * p<0.1.
+The magnitude of network expansion is economically meaningful.
+A 92-connection increase at the median represents substantial growth from the control baseline of 225, comparable to the network size differences between mid-career and senior professionals in our sample.
+Moreover, the increase occurs despite LinkedIn’s 500-connection display cap, which likely attenuates our estimates: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, suggesting the true effect exceeds our measured 52.7-connection average.
+This network expansion cannot be solely attributed to new employment.
+Both Mentoring and Challenges participants secure technology jobs at high rates, yet only Mentoring participants expand their networks.
+Had job acquisition been the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable increases in both programs.
+The absence of connection growth among Challenges participants, despite their 38% technology employment rate, indicates that Mentoring’s network effects operate independently of (and plausibly prior to) job acquisition.
+This timing is consistent with mentors actively facilitating introductions during the program, as reported in our qualitative interviews, rather than networks expanding passively after job placement.
 5.3 Heterogeneous Treatment Effects
-We explore heterogeneity in treatment effects by conducting an analysis of treatment effects across different subgroups using the Cox proportional hazard model.
-We categorize applicants to the Mentoring and Challenges programs into non-overlapping groups based on four different types of characteristics: STEM education, length of professional experience, type of residence, and age.
-For each of these categories, we estimate a hazard model that includes treatment and subgroup indicators along with their interaction.
-The results of this analysis are presented in Table 3.
-It’s important to note that applicants in each of the non-overlapping groups may have other varying characteristics; to account for this, we include a full set of controls and interactions in Appendix I in Table 10.
-Furthermore, here we present only selected subgroups; Appendix I presents all the interaction terms.
+The hypothesized contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program.
+We exploit treatment effect heterogeneity to provide evidence consistent with these mechanisms: if Mentoring operates primarily through network expansion (as suggested by LinkedIn connections analysis), effects should concentrate among participants with weak existing networks; if Challenges works through skill signaling, effects should be largest for those with weaker pre-treatment credentials.
+We estimate heterogeneous treatment effects using Cox proportional hazard models that estimate treatment effects in subgroups defined by the pre-treatment characteristics of interest.
+We control for the full covariate set.
+Table 4 presents heterogeneous treatment effects across three participant characteristics that map to credential strength and network access.
+These patterns provide evidence on which populations each program serves most effectively and tests the distinct mechanisms we hypothesize.
+The clearest differential pattern emerges from professional experience.
+Challenges generates 28 percentage points larger treatment effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced participants (SE 11).
+The cross-program differential is 49 percentage points (SE 11).
+This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal their technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records.
+Educational background reveals similar patterns.
+Recall that all participants were pre-screened for technical competency, so heterogeneity by educational background likely reflects differences in credential signals and professional networks, and to a lesser degree underlying ability.
+Among Mentoring participants, those without STEM degrees show 19 percentage points larger effects than those with STEM backgrounds (SE 11).
+This pattern is consistent with the network access: non-STEM participants followed educational paths outside technology-adjacent fields, limiting opportunities to build connections with peers entering the sector.
+In contrast, Challenges shows no STEM heterogeneity (difference: 0.3pp, SE 11).
+This null result demonstrates that portfolio signals are equally valuable regardless of educational credentials—employers evaluate demonstrated capability rather than inferring skill from degrees.
+Geography provides additional evidence on network mechanisms.
+Mentoring generates 26 percentage point treatment effects for participants from small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, p=0.11).
+This pattern aligns with network barriers binding more tightly outside major urban centers, where local technology ecosystems are sparse and professional connections to the sector are limited.
+Mentor networks based in Warsaw and other major cities provide greater marginal value when participants lack local access to technology professionals.
+Challenges shows no significant geographic heterogeneity (difference: -2pp, SE 1.2), consistent with online portfolio signals functioning equally regardless of location, employers evaluate portfolios based on technical content rather than applicant geography.
 Table 4: Heterogeneity in Treatment Effects: Mentoring vs. Challenges.
 Mentoring Challenges
-Baseline CATE Baseline CATE Diff. Mentoring - Challenges
+Baseline CATE Baseline CATE Diff. Mentoring – Challenges
 STEM degree
-Yes 0.437 (0.02) 0.030 (0.03) 0.191 (0.02) 0.030 (0.03) -0.001 (0.04)
-No 0.239 (0.02) 0.232 (0.03) 0.135 (0.01) 0.086 (0.02) 0.146 (0.03)
-Difference subgroups 0.199 (0.03) -0.202 (0.05) 0.055 (0.02) -0.055 (0.03) -0.147 (0.05)
+Yes 0.538 (0.057) 0.038 (0.079) 0.283 (0.059) 0.12 (0.099) -0.081 (0.082)
+No 0.362 (0.058) 0.228 (0.07) 0.259 (0.035) 0.116 (0.057) 0.111 (0.068)
+Difference subgroups 0.176 (0.081) -0.189 (0.106) 0.024 (0.068) 0.003 (0.114) -0.193 (0.106)
 Length of professional experience
-Less than 5 years 0.429 (0.03) 0.008 (0.04) 0.119 (0.01) 0.136 (0.03) -0.128 (0.05)
-More than 5 years 0.344 (0.02) 0.142 (0.03) 0.178 (0.01) 0.021 (0.02) 0.121 (0.03)
-Difference subgroups -0.085 (0.03) 0.134 (0.05) 0.059 (0.02) -0.115 (0.03) 0.248 (0.06)
-Type of residence
-Small town 0.307 (0.03) 0.285 (0.05) 0.161 (0.02) 0.160 (0.03) 0.126 (0.06)
-Warsaw 0.353 (0.02) 0.075 (0.03) 0.148 (0.01) 0.051 (0.02) 0.023 (0.03)
-Difference subgroups -0.046 (0.03) 0.211 (0.05) 0.013 (0.02) 0.109 (0.04) 0.102 (0.07)
-Age
-Over 30 years old 0.261 (0.02) 0.125 (0.03) 0.130 (0.01) 0.114 (0.02) 0.011 (0.04)
-Under 30 years old 0.428 (0.02) 0.117 (0.04) 0.178 (0.01) 0.025 (0.02) 0.092 (0.04)
-Difference subgroups -0.167 (0.03) 0.008 (0.05) -0.049 (0.02) 0.089 (0.03) 0.011 (0.05)
-Note: Estimates of average treatment effects conditioned on subgroups in Mentoring and Challenges.
-The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction.
-The first and third columns show the subgroup’s average outcome within control subjects.
-The second and fourth columns show the difference between treatment and control in the subgroup.
-The last column shows the difference between Mentoring and Challenges of the treatment effects.
+More than 5 years 0.412 (0.049) 0.202 (0.066) 0.31 (0.039) -0.003 (0.065) 0.205 (0.063)
+Less than 5 years 0.556 (0.075) 0 (0.087) 0.19 (0.042) 0.281 (0.075) -0.281 (0.086)
+Difference subgroups -0.144 (0.089) 0.203 (0.109) 0.120 (0.057) -0.284 (0.099) 0.486 (0.106)
+Small Town
+Small town 0.429 (0.085) 0.26 (0.101) 0.256 (0.067) 0.149 (0.13) 0.111 (0.108)
+Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166 (0.089) -0.133 (0.093)
+Difference subgroups -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
+Note: Estimates of the average treatment effects conditioned on subgroups in Mentoring and Challenges.
+The outcome variable is Tech job, and estimates are obtained using Cox proportional hazards models per subgroup with a full set of covariates.
+The first and third columns show the average outcome per subgroup within the control groups.
+The second and fourth columns show the difference between the treatment and control groups in each subgroup.
+The last column shows the difference in the treatment effects between Mentoring and Challenges.
 Standard errors are in parentheses.
-First, we observe that participants with a STEM degree have higher baseline outcomes than those without a STEM degree.
-Both programs, Mentoring and Challenges, are more effective for participants without STEM degrees, bridging the outcome gap across these two groups.
-Specifically, in the case of Mentoring, we find that the treatment entirely eliminates the STEM outcomes gap.
-When comparing the two programs, we find no statistically significant difference for participants with STEM degrees, while for those without STEM degrees, Mentoring leads to nearly 15 percentage points higher treatment effects than Challenges.
-Second, when focusing on the subgroups with more or less than 5 years of professional experience, we find that Challenges is more effective for participants starting their professional career, with nearly 13 percentage points higher treatment effects in this subgroup.
-Conversely, for participants with longer professional experiences, Mentoring appears more effective.
-Third, we find that participants from smaller towns benefit more from both programs.
-However, this difference is particularly pronounced in the case of Mentoring.
-Participants from smaller towns have lower baselines than participants from Warsaw, but after the treatment, the gap is entirely closed.
-In fact, Mentoring participants from smaller towns achieve, on average, higher outcomes than Mentoring participants from Warsaw.
-Finally, participants below 30 years old have higher baselines than those over 30 years old.
-However, both programs help close the gap between the two age groups by having higher treatment effects for older participants.
 Additional Characteristics from Mentoring.
-In the case of Mentoring, we observe several additional characteristics; these come from the Mentoring registration survey, which was more extensive than the Challenges one, and we also observe from characteristics of mentors.
-In Table 4, we present additional results documenting heterogeneity in treatment effects along those additional dimensions.
+The Mentoring registration survey was more extensive than the Challenges one, allowing us to observe more characteristics.
+We also obtain the characteristics of mentors from their registration survey.
+In Table 5, we present additional results documenting heterogeneity in treatment effects along these dimensions.
 Table 5: Heterogeneity in Treatment Effects: Mentoring.
-Variable value is 1 Variable value is 0
+Characteristic variable value is 0 Characteristic variable value is 1
 Baseline CATE Baseline CATE Diff. Subgroups
 Mentee characteristics
-Mother 0.348 (0.02) 0.288 (0.04) 0.398 (0.02) 0.068 (0.03) 0.220 (0.05)
-Applied before 0.384 (0.03) 0.141 (0.05) 0.330 (0.02) 0.121 (0.03) 0.020 (0.05)
+Mothers 0.237 (0.07) 0.348 (0.097) 0.532 (0.048) 0.071 (0.062) 0.277 (0.115)
+Applied before 0.5 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143 (0.06) -0.004 (0.124)
 Mentor characteristics
-Managerial exp. 0.285 (0.02) 0.223 (0.03) 0.414 (0.02) 0.012 (0.03) 0.211 (0.05)
-Long exp. tech 0.288 (0.02) 0.226 (0.04) 0.389 (0.02) 0.052 (0.04) 0.174 (0.05)
-1st time mentor 0.354 (0.02) 0.131 (0.03) 0.342 (0.02) 0.112 (0.03) 0.019 (0.05)
-Note: Estimates of average treatment effects conditioned on subgroups in Mentoring.
-The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazard models, which include treatment indicator, subgroup indicator, and interaction.
-Applied before takes the value of 1 if the mentee has applied to the program in the past, but did not get selected into the program, and reapplied, and the value of zero if this is the first time applying to the program.
-Long exp.
-in tech takes the value of 1 if the mentor worked for over 5 years in the technology sector (the median value), and zero otherwise; 1st time mentor takes the value of 1 if this is the first time the mentor provides formalized mentoring service, and zero otherwise.
-Columns 2 and 3 describe the values for subgroups where the subgroup indicator takes the value 1, while columns 4 and 5 when the indicator takes the value of 0.
-The last column is the difference in the effectiveness of Mentoring across the two subgroups.
-Columns 1 and 3 are average outcomes in control groups, and columns 2 and 4 are the estimates of the average treatment effect.
+Managerial exp. 0.378 (0.056) 0.259 (0.073) 0.549 (0.059) 0.002 (0.074) 0.257 (0.104)
+Long exp. in tech 0.39 (0.063) 0.277 (0.085) 0.512 (0.054) 0.021 (0.066) 0.257 (0.107)
+1st time mentor 0.444 (0.062) 0.155 (0.08) 0.476 (0.055) 0.1 (0.073) 0.056 (0.108)
+Note: Estimates of the average treatment effects conditioned on subgroups in the Mentoring program.
+The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazards models, which include the treatment indicator and the set of controls.
+We estimate the model per subgroup.
+"Long exp.
+in tech" takes the value of 1 if the mentor has worked for over 5 years in the technology sector (the median value).
+Columns 2 and 3 describe the values for subgroups when the subgroup indicator takes the value 0, while columns 4 and 5 describe the values when the indicator takes the value 1.
+The last column shows the difference in the effectiveness of Mentoring between the two subgroups.
+Columns 1 and 3 are the average outcomes in the control groups, and columns 2 and 4 are the estimates of the average treatment effect.
 Standard errors are in parentheses.
-First, we observe that the Mentoring program is particularly effective for mothers.
-Mothers who were selected by mentors but randomized into the control group had a baseline probability of 35% of finding a job in technology, compared to 40% for those without children.
-However, the average treatment effect for mothers is significantly larger at 29 percentage points, compared to just 7 percentage points for those without children.
-As a result, Mentoring not only enables mothers to catch up with applicants without children but even surpass them in the probability of getting a job in technology.
-In contrast, we do not observe a high difference in the effectiveness of Mentoring between first-time applicants and those who applied in previous editions.
-Turning our attention to the characteristics of mentors, we find that mentors with managerial experience and those with a long history in the technology sector (over 5 years) tend to select, on average, less promising candidates, as indicated by baseline outcome probabilities of 29% for both groups, compared to the baseline of 41% for mentors without managerial experience and 39% for those with short tenure in tech.
-However, these mentors with managerial experience and long tech sector tenure are highly effective, with their mentees achieving a probability of over 50% of securing a job in technology.
-Interestingly, we also find that first-time mentors are as effective as mentors who have had previous mentoring experience.
-These findings are consistent with the notion that effective mentors leverage their knowledge, skills, and relationships accumulated over years of working in the technology sector, rather than improving their mentoring capabilities through the practice of mentoring.
-Patterns of treatment effects heterogeneity are consistent with some of the mechanisms through which the two programs might impact employability that we hypothesized in Section 3.
-Participants from smaller towns and those without family or friends working in tech (see Table 10) might have less extensive professional networks in the technology sector, and mentors with longer experience in tech are likely to have richer networks.
-Thus, the finding that participants with these characteristics benefit more from Mentoring is consistent with the mechanisms that Mentoring helps mentees by bolstering their professional network and enhancing their capacity to utilize these networks effectively.
-As for Challenges, we hypothesized that the program would equip participants with credible signals of their technical skills.
-This mechanism is consistent with the finding that participants without STEM degrees and those with shorter professional experience, thus fewer other signals on their resumes, exhibit larger benefits.
+Mentors with managerial experience generate 26 percentage point larger treatment effects than those without (SE 10).
+Similarly, mentors with over five years of technology sector experience produce 26 percentage point larger effects (SE 11).
+These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements.
+Interestingly, mentor selection patterns reveal that experienced, managerial mentors choose less promising candidates.
+Control group participants matched to managerial mentors have 38% baseline technology employment compared to 55% for those matched to non-managerial mentors.
+Yet treated participants of managerial mentors achieve higher outcomes (64% vs.
+64%), because the treatment effect is substantially larger.
+This suggests that experienced mentors either (i.) deliberately select candidates who need more help, recognizing their own capacity to assist difficult cases, or (ii.) value different candidate attributes than baseline employment probability.
+Notably, first-time mentors are as effective as repeat mentors (difference: 6pp, SE 11).
+This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks, technical knowledge, and industry standing—not from mentoring skill developed through practice.
+A senior engineer with extensive connections generates value on her first mentoring experience; repeated mentoring does not improve outcomes.
+Mothers and Work-Family Barriers.
+The largest heterogeneity we observe is for mothers.
+Among non-mothers, Mentoring generates 7 percentage point treatment effects (SE 6); among mothers, the effect is 35 percentage points (SE 10), a difference of 28 percentage points (SE 12).
+Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 50% for non-mothers.
+Mentoring completely eliminates this gap: treated mothers achieve 59% employment, surpassing non-mothers.
+This large effect for mothers is consistent with Mentoring addressing multiple barriers simultaneously.
+First, mothers may have weaker networks if career interruptions weakened professional connections.
+Second, work-family conflict concerns may reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful mentors in tech may counter this.
+This finding contributes to understanding gender gaps in technology beyond representation.
+Even among women who have acquired technical skills and are actively seeking technology employment, motherhood creates additional barriers.
+Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that confidence and network deficits also play important roles, and relatively light-touch mentoring interventions can substantially mitigate these barriers.
+The contrasting heterogeneity across programs aligns with our theoretical framework.
+Network barriers might bind most tightly for geographically isolated women and mothers, the groups showing the largest Mentoring effects.
+Information frictions bind most for early-career women with limited work history, those showing the largest Challenges effects.
+The programs’ differential impacts across populations indicate that successful policies must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly.
+Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in which constraints bind most tightly, with implications for optimal program assignment that we explore in Section 6.
 6 Counterfactual Program Assignment Policies
 The number of spots in each program is limited.
-The capacity of Mentoring is constrained by the availability of mentors, while in Challenges, we capped spots to 300 to accommodate potential operational complexities and logistical issues.
-Considering the substantial treatment effects of both programs, one might posit that relaxing these capacity constraints will create significant value.
-Nonetheless, capacity constraints will likely remain a long-term issue for the mentoring program and may persist, albeit to a lesser degree, for Challenges.
-Consequently, Dare IT will continue being faced with the question of which applicants to accept for each program.
-With heterogeneous treatment effects, admission policies favoring applicants with high predicted benefits could enhance Dare IT programs’ impact.
-This section aims to quantify such potential gains.
-We carry out three analyses: (i.) we estimate gains from the introduction of a policy that, for given capacity levels in each program, assigns applicants to programs to maximize the sum of predicted treatment effects across the two programs; (ii.) we compare the overall impact of Mentoring when mentors select most promising candidates to a case when they select to maximize expected benefits; ((iii).
-finally, we evaluate several simple policies based on heuristics.
-An assignment policy is a mapping from applicants’ characteristics $X\in\mathcal{X}$ and programs capacity levels $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT.
+The availability of mentors restricts the capacity of Mentoring.
+In Challenges, we limited the available spots to 300 to accommodate potential operational complexities and logistical problems.
+Considering the substantial treatment effects of both programs, relaxing these capacity constraints will increase Dare IT’s impact.
+However, capacity constraints are likely to remain a long-term issue for the Mentoring program and may persist, albeit to a lesser extent, for Challenges.
+Consequently, Dare IT will continue to face the question of which applicants to accept for each program.
+Given the heterogeneous treatment effects, admission policies that favor applicants with high predicted benefits could enhance the impact of Dare IT programs.
+This section aims to quantify these potential gains.
+We perform three analyses: (i.) we estimate gains from the introduction of a program admission policy that, given capacity levels in each program, assigns applicants to programs, maximizing the sum of predicted treatment effects of programs’ participants; (ii.) we evaluate the gains from relaxing capacity constraints with the targeting policy in place; (iii.) we compare the overall impact of Mentoring when mentors select the most promising candidates to a case where they select to following alternative strategies.
+An assignment policy is a mapping from the characteristics of the applicants $X\in\mathcal{X}$ and the capacity levels of the programs $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT.
 $\mathcal{Q}$ is the grid of possible capacity levels under consideration.
-Policy estimation and evaluation are carried out separately.
-The objective of policy estimation is to obtain a mapping from user characteristics and programs’ capacity level to program assignments’.
-This mapping $\hat{\pi}$ is estimated in a training set.
-Across all analyses, we obtain training and testing sets by randomly splitting data at the applicant level with equal probability.
-Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping to the testing set, where for each $i$ in the test set and each $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$.
-Second, for each $Q\in\mathcal{Q}$, we construct the AIPW estimates of the “policy value,” that is, the value of the outcome when treatments are assigned according to $\pi(X_{i},Q)\in\mathcal{P}$.
-We do that using cross-fitting in the testing sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds (we use 100 repetitions with 10 folds and consider the median outcome across repetitions) [^16] We follow the approach suggested in Chernozhukov et al.
-(2017)..
-Throughout this section, we use Cox proportional hazard model with the full set of interactions as the outcome model and the logit model as the propensity model.
-Appendix J provides details of the two models and of the cross-fitting procedure.
-6.1 Optimal assignment policies
-Let $z_{ip}\in\left\{0,1\right\}$ be a variable indicating whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments maximizing the sum of treatment effects given by Equation 3:
+Policy estimation and evaluation are performed separately in train and test sets.
+Across all analyses, we obtain train and test sets by randomly splitting data at the applicant level with equal probability ($\mathcal{I}_{train}$ and $\mathcal{I}_{test}$ are resulting sets of applicants in the train and test data sets, respectively).
+The objective of policy estimation is to obtain a mapping from applicant characteristics and programs’ capacity levels to program assignments.
+We use the train set to obtain the estimate of the policy $\pi$, which we denote $\hat{\pi}$.
+Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping $\hat{\pi}$ to the test set, where for each $i$ in the test set and $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of the applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$.
+Second, we construct the AIPW estimate of the “policy value,” that is, the outcome value when treatments are assigned according to $\hat{\pi}(X_{i},Q)\in\mathcal{P}$.
+We do that using cross-fitting at the applicant level in the test sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds; specifically, we use 100 repetitions with 5 folds and consider the median outcome across repetitions in (chernozhukov2017double).
+Throughout this section, we combine the two experimental data sets with data on non-selected applicants to Mentoring and late applicants to Challenges.
+Thus, we have 997 applicants who did not participate in any program within our sample, resulting in a total sample of 1314 applicants and 19,000 applicant-month observations.
+For both policy estimation and evaluation, we use survival forests to estimate outcome models and regression forests for the propensity models.
+The propensity model adjusts for differences between participants and non-participants determined by selection to Mentoring and differences in the populations of applicants to Mentoring and to Challenges.
+Appendix H provides details of the models used in policy estimation.
+Appendix I, presents details of the models used in the policy evaluation, the cross-fitting procedure, and validates the approach by comparing the off-policy estimates with the randomized controlled trial estimates.
+Our procedure recovers the experimental estimates.
+Outcome measure.
+In all the preceding analyses, we measured treatment effects as changes in the probability of securing technology employment.
+For policy evaluation, we adopt an alternative metric that captures additional dimensions of program impact: the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period.
+The RMST metric offers two advantages for comparing policies.
+First, it incorporates both whether participants find jobs and how quickly they do so, a policy that helps participants find jobs faster, even if it doesn’t change the ultimate employment probability, generates economic value.
+Second, RMST estimates aggregate more naturally across heterogeneous populations and policy scenarios, facilitating comparisons of the assignment rules.
+We report RMST gains in months: negative values indicate participants find technology employment sooner relative to non-participants.
+An RMST gain of -3 months means treated participants find technology jobs an average of three months earlier.
+We use survival forests to estimate average treatment effects (athey2019generalized; cui2023estimating).
+6.1 Targeted Assignment Policies
+Let $z_{ip}\in\left\{0,1\right\}$ be a variable that indicates whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments that maximize the sum of treatment effects given by Equation 2:
 $\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\thickspace\forall p\text{ \& }\sum_{p\in\mathcal{P}}z_{ip}=1\thickspace\forall i,$ (2)
 where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$.
-The first constraint ensures that the capacity limits are not violated, and the second one is that each applicant is assigned to one program.
-A policy that induces allocation $z_{ip}^{*}$ is what we call optimal policy.
-In Algorithm 1 in Appendix K, we provide further details of the estimation of optimal policy using training data, where we let $\hat{\pi}$ denote the estimate.
-We estimate the optimal policy at three capacity levels: first, with the existing capacity (allowing 13% of all applicants into Mentoring and 15% into Challenges) – Optimal 1, second, increasing the capacity of Challenges to accommodate 50% of applicants – Optimal 2, and third, a scenario where each program can admit all applicants – Optimal 3.
-We compare these optimal policies against several alternatives: a policy accepting all applicants to Mentoring, a policy enrolling all applicants in Challenges (both these policies have increased capacities to accept all applicants), and the existing allocation policy (Status quo).
-We use the average treatment effect as the metric to compare policies.
-The results are presented in Figure 3.
-Figure 3: Evaluation of counterfactual policies.
-Note: Estimates of values of policies.
-The policy value is the mean probability of having a tech job across all participants and is measured by the mean of AIPW scores under allocation induced by each policy across all participants.
-Standard errors clustered at a participant level.
-Whiskers show confidence intervals.
-Optimal 1 has the existing capacity levels.
-Optimal 2 increases the capacity of Challenges to 50%, and Optimal 3 has no capacity limits.
-First, we find that the optimal policy, with existing capacity levels, markedly improves the average treatment effect when compared to the Status quo, with an increase from 12 to 22 percentage points.
-Furthermore, this policy is also substantially more effective than policies that assign all applicants to Mentoring or to Challenges.
-This result follows from the high degree of treatment effect heterogeneity across both programs and shows that a targeting policy can substantially augment Dare IT’s effectiveness.
-Second, we find that the policy expanding the number of Challenges spots to 50% of applicants (Optimal 2) increases the average treatment effect to 17 percentage points and Optimal 3 to 15 percentage points.
-Such an increment highlights that the there are sizeable subpopulations amongst Dare IT applicants that have higher treatment effects.
-Moreover, we find only a small and statistically insignificant difference in policy value when contrasting Optimal 2 to the Optimal 3 policy.
-In Table 5, we present the gains from optimal policy per assignment group.
-A program assignment group consists of all applicants that under a policy are assigned to this program, i.e., Mentoring assignment group are all applicants that are assigned to Mentoring.
-We are interested in comparing the benefits from the participation in each program of each of these groups under.
-We find that the group allocated to Mentoring substantially gains from their allocation, yielding an ATE of 0.199 (S.E.
-0.017).
-This group would also substantially benefit from Challenges with the average treatment effect of 0.276 (S.E.
-0.019).
-The difference between assigning them to Mentoring over Challenges is significant -0.077 (S.E.
-0.021).
-However, the group assigned to Challenges benefits from this assignment even more.
-The average treatment effect of Challenges amounts to 0.244 (S.E.
-0.017) compared to 0.109 (S.E.
-0.012) of Mentoring.
-Lastly, the non-participating group has substantially lower benefits from assignment to Mentoring or to Challenges, average treatment effects of 0.101 (S.E.
-0.006) and 0.073 (S.E.
-0.003) respectively.
-Table 5: Gains from the optimal assignment
-Mentoring group Challenges group Nothing group
-Share 0.15 0.50 0.35
+The first constraint ensures that capacity limits are not violated, and the second is that each applicant is assigned to one allocation.
+A policy that induces allocation $z_{ip}^{*}$ is what we call the targeted policy.
+$\tau_{i}^{p}$’s are estimated using the AIPW estimator trained on the train set.
+In Algorithm 1 in Appendix H, we provide further details of the estimation of the targeted policy.
+The outcome of interest, policy value, is the participants’ average treatment effect given program allocation induced by the policy $\pi$.
+Specifically, we use the AIPW estimator to estimate the average difference in outcomes between being selected to participate based on the assignment policy and not participating in any Dare IT program.
+This emphasis on impact, rather than average outcomes of participants, aligns with Dare IT’s mission to create the greatest benefit for their participants.
+6.2 Value of Targeted Policies
+Given limited program capacity, which applicants should be assigned to which program to maximize overall impact?
+We estimate the targeted policy at three capacity levels: the 15/15/70 policy approximates existing capacity (assigning 15% of all applicants to Mentoring and 15% to Challenges); the 33/33/34 policy increases capacity to 33% for each program; and the 50/50/0 policy allows each program to admit 50% of applicants, ensuring a program slot for every applicant.
+We compare these policies against random allocation as a benchmark (random policy allocates all applicants at random to either Mentoring or Challenges).
+In Section 6.3, we examine how mentor-based selection compares to the algorithm-driven policy in greater detail.
+The policies vary in both capacity levels and whether participants are systematically targeted.
+To understand which applicant types are assigned to each program under data-driven targeting, Figure 3 presents standardized differences in baseline characteristics between optimally assigned individuals and those not assigned to treatment.
+We implement the 15/15/70 targeted allocation policy and compare covariate distributions of each assigned group to the not-assigned group.
+Circles represent comparisons for Mentoring; triangles represent Challenges.
+Dashed vertical lines mark the $\pm$0.1 threshold for imbalance.
+Figure 3: Characteristics of Optimally Assigned Groups
+Notes: Standardized differences in baseline characteristics between optimally assigned individuals and those not assigned under the 15/15/70 allocation policy.
+Circles show Mentoring vs Not Assigned; triangles show Challenges vs Not Assigned.
+The solid vertical line at zero indicates perfect balance; dashed lines mark the $\pm$0.1 threshold.
+Standardized differences are calculated as the difference in means divided by the pooled standard deviation.
+Positive values indicate the optimally assigned group has higher values on that characteristic.
+The targeted allocation exhibits clear differentiation in targeting patterns across programs.
+Applicants selected for Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old.
+Notably, although the algorithm does not use information about prior mentor selection, previously mentor-selected applicants are disproportionately targeted by the algorithm—suggesting the model captures similar signals to those used by mentors.
+For Challenges, the algorithm shows more balanced selection across characteristics.
+Assigned applicants are moderately more likely to have been previously selected by mentors but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience compared to those not selected.
+Main result.
+Table 6 shows gains from data-driven targeting, with effect sizes varying by program and capacity constraint.
+Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, which corresponds to an 86% improvement over random allocation’s 2.33-2.33 months.
+These gains exhibit a strong capacity-selectivity tradeoff: as capacity expands from 15% to 50%, per-person treatment effects decline (from 4.34-4.34 to 2.86-2.86 months overall), though even universal coverage (50/50/0) achieves effects 23% larger than random allocation.
+Thus, we find that there are gains from targeting across programs.
+The targeting gains differ across programs.
+Mentoring exhibits the largest effects under selective policies (5.55-5.55 months at 15% capacity versus 2.31-2.31 under random), while Challenges shows more modest differentiation (3.13-3.13 versus 2.35-2.35).
+Notably, even at 50% capacity where the algorithm must accept marginal applicants with lower predicted treatment effects, Mentoring maintains treatment effects 54% larger than random allocation, demonstrating that observable characteristics contain substantial predictive power for treatment response beyond what random assignment captures.
+Table 6: Targeted Allocation Policies: Treatment Effects by Program
+Capacity Allocation Mentoring Challenges Overall
+(Ment. / Chal. / None) Effect Effect Effect
+Panel A: Benchmark
+Random Allocation -2.31 (0.022) -2.35 (0.021) -2.33 (0.022)
+Panel B: Targeted Allocation Policies
+15% / 15% / 70% -5.55 (0.086) -3.13 (0.069) -4.34 (0.055)
+33% / 33% / 34% -4.33 (0.052) -2.68 (0.041) -3.50 (0.033)
+50% / 50% / 0% -3.55 (0.039) -2.17 (0.027) -2.86 (0.024)
+Notes: Treatment effects measured as change in months to employment (negative values = faster job finding).
+Standard errors in parentheses.
+Each row shows a capacity-constrained allocation policy specifying the share of applicants assigned to Mentoring, Challenges, or neither program.
+Random Allocation assigns each program to a random 50% of applicants.
+Table 7 examines treatment effects within each assignment group under the 15/15/70 targeted policy.
+An assignment group consists of all applicants allocated to a specific program (Mentoring, Challenges, or neither) under this policy.
+For each group, we estimate counterfactual treatment effects under alternative program assignments to assess whether the targeted allocation correctly matches individuals to their highest-benefit program.
+The results reveal effective targeting.
+The Mentoring assignment group receives large benefits from their targeted allocation (5.55-5.55 months, SE = 0.086), substantially exceeding what they would gain from Challenges (3.12-3.12 months, SE = 0.062).
+The difference of 2.43 months (SE = 0.106) is statistically significant.
+The Challenges assignment group also benefits most from their assigned program (3.13-3.13 months, SE = 0.069) compared to Mentoring (1.87-1.87 months, SE = 0.050); a difference of 1.26 months (SE = 0.085).
+This pattern demonstrates benefits from targeting across both programs rather than only selecting to a program.
+The Out of Dare IT group (70% of applicants receiving no program assignment) exhibits substantially smaller treatment effects under either counterfactual assignment: 1.72-1.72 months (SE = 0.024) for Mentoring and 2.02-2.02 months (SE = 0.022) for Challenges.
+While this group would experience gains from program participation, their predicted effects are smaller than those optimally assigned to treatment.
+Table 7: Gains From Assignment Policy per Assignment Group
+Mentoring group Challenges group Out of Dare IT group
+Share 0.15 0.15 0.70
 ATE under assignment to:
-Optimal 0.199 (0.017) 0.244 (0.017)
-Mentoring 0.199 (0.017) 0.109 (0.012) 0.101 (0.006)
-Challenges 0.276 (0.019) 0.244 (0.017) 0.073 (0.003)
+Targeted -5.552 (0.086) -3.132 (0.069)
+Mentoring -5.552 (0.086) -1.872 (0.050) -1.720 (0.024)
+Challenges -3.122 (0.062) -3.132 (0.069) -2.017 (0.022)
 Difference in ATE between assignments:
-Mentoring - Challenges -0.077 (0.021) -0.135 (0.025) 0.028 (0.006)
-Note: Gains from the optimal assignment by assignment group from Optimal 1 - 13% of applicants get be admitted to Mentoring and 15% to Challenges.
-Groups are defined as all applicants assigned to the specific program.
-The columns describe the outcomes of groups under various allocations - the group membership is determined using Algorithm 1 and is constant across the comparisons here, the allocations change.
-The first row is the share of applicants assigned to a group.
-Rows two to four describe the average treatment effect of the group under various assignments.
-The 5th row is the average treatment effect under Mentoring less the average treatment effect under Challenges.
-Standard errors are next to each value in parentheses and are clustered at the applicant level.
-When comparing the three policies based on optimal allocation, it’s evident that the current capacity levels limit Dare IT’s impact but that additional spots do not create more value beyond a certain capacity.
-To further unpack this, in Figure 4, we analyze the value of optimal policy under a sequence of capacity values and compare with random assignment to either Mentoring or Challenges.
-Purple dots in Figure 4, show the average probability of getting a job in technology across all applicants (including those not participating in Dare IT programs), for the total capacity of the two programs ranging from 0 to 100% (in all cases, the two programs have the same capacity, so for example the total capacity of 50% means that, up to 25% of applicants get be assigned to either program).
-We compare the outcome under these policies with policies that assign applicants randomly to Mentoring (in yellow) and to Challenges (in red).
-Figure 4: The value of targeting at various capacity levels
-Note: Estimates of mean outcomes across all applicants at various capacity levels of Dare IT programs and allocation mechanisms.
-The policy value is the predicted share of applicants getting a tech job.
-The share of applicants getting a spot refers to the total capacity in both programs.
-Challenges Random (in orange) and Mentoring Random (in red) assign applicants at random between the program and Out of Dare IT.
-Targeting policy (in purple) follow the optimal policy approach.
-Our analysis indicates that policies randomly admitting applicants to either the Mentoring or Challenges programs lead to consistently lower policy values when compared to the targeted policy across all capacity levels.
-When capacities are low, targeted policies prioritize applicants with the highest benefits.
-A meaningful difference in the average probability of getting a job across all applicants quickly emerges between the targeted and random policies; consequently, already at the 25% capacity levels, there is an approximately 4 percentage points gap between targeted and random policies.
-Interestingly, even when capacity constraints are lifted to accommodate all applicants, the targeted strategy continues to lead to higher outcomes than the random policies based on one program.
-This discrepancy remains because there of the difference in benefits that applicants receive across the two programs, i.e., there are some applicants that benefit more from Challenges than they do from Mentoring.
-6.2 Counterfactual selection policies for mentors
-Applicants selected by mentors and randomized to be the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2).
-This suggests that mentors tend to prefer highly promising applicants who are likely to secure a tech job irrespective of their participation in Mentoring.
-In this section, we examine alternative selection strategies that mentors could adopt and their potential implications for the effectiveness of the mentoring program.
-Specifically, we consider four scenarios: (i) Highest CATE – targeting applicants predicted to have the highest treatment effects, (ii) Most Promising – targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising – selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated – where mentors select participants that have the highest predicted outcomes if they participate in Mentoring; additionally, we include a case in which mentors select at random (Random Choice) and the observed allocation (Observed) as benchmarks.
-In each policy, we assume that 36% of applicants participate in Mentoring, which corresponds to the share of applicants that were selected by mentors.
-[^17] We consider a higher number of participants than the current number of spots in Mentoring to amplify the impact of policies on outcomes across all applicants; the ordering of policies does not depend on the share of applicants admitted to Mentoring as long as this number is constant across policies.
-We follow the same off-policy approach as previously.
-We consider the dataset of all applicants to Mentoring, which consists of 151 individuals who participated in the program, 147 mentors selected but randomized to the control group, and 468 non-selected applicants.
-We split data into train and test with equal probability.
-We use the training dataset to estimate the mapping from applicants’ characteristics to assignments.
-To do that, we train the Cox proportional hazard model in the training data and predict outcomes with and without treatment in the testing data.
-Then, we use these predictions to determine which applicants get selected under each policy.
-To evaluate policies, we compute AIPW scores analogously to the optimal policy evaluation (Appendix J.1): using only testing data, we train the propensity model (logit model) and a new outcome model (Cox hazard model) using cross-fitting.
-We compute 100 cross-fitting rounds and take the median outcome.
-Table 6 shows the results.
-Table 6: Mentors’ selection policies.
-Policy Mean ATE Baseline outcome treated group Mean outcome all applicants
-Highest CATE 0.178 (0.012) 0.187 (0.008) 0.232 (0.005)
-Most Promising 0.062 (0.012) 0.285 (0.009) 0.192 (0.005)
-Least Promising 0.127 (0.007) 0.055 (0.005) 0.211 (0.005)
-Highest Chances if Treated 0.130 0.012 0.237 (0.008) 0.218 (0.005)
-Observed 0.096 (0.012) 0.275 (0.008) 0.205 (0.005)
-Random 0.094 (0.011) 0.191 (0.008) 0.203 (0.005)
-Note: Outcomes and treatment effects across different mentors’ selection rules.
-Mean ATE is estimated using AIPW estimator with cross-fitting.
-Baseline outcome treated group is the predicted outcome without treatment for the group that receives treatment in the considered policy.
-Mean outcome all applicants corresponds to the average probability of getting a job in tech across all applicants, where the group assigned to treatment under considered policy receives treatment, and the remaining applicants do not participate.
-Standard errors clustered at the applicant level.
-We find that prioritizing applicants with the highest anticipated treatment benefits results in higher average outcomes for all applicants and a more significant treatment effect than other strategies.
-Using the Highest CATE policy, 23% of applicants secure a tech job, a noticeable improvement over the 20% success rate under the Random or Observed policies.
-The average treatment effect under Highest CATE is an 18 percentage points increase in the probability of having a tech job for the treated group.
-Interestingly, targeting applicants with the highest chances of having a job if not treated (Most Promising) leads to treatment effects that are lower than selecting applicants randomly.
-Moreover, a policy aimed at the seemingly less promising applicants (Least Promising) — which goes against current mentor preferences — produces an average treatment effect of 13 percentage points.
-This results in an average outcome for all applicants surpassing those achieved by the Most Promising and Random policies.
-The policy of targetting applicants that have the highest chances of getting a tech job if treated (Highest Chances if Treated) results in a 13 percentage points average treatment effect, which is similar to Least Promising.
-74% of applicants assigned to Mentoring under Highest CATE are also assigned to the program under Highest Chances if Treated, which is expected considering the fact that these two policies are most promising.
-However, there is a low overlap between Highest CATE and Least Promising only 7%.
-Furthermore, there is a substantial overlap between the observed selection by mentors and the Highest CATE policy – 54%, but a very low one between Observed and Least Promising, only 4%.
-These observations suggest that a group of applicants has low chances of securing a tech job when not treated and are systematically overlooked by mentors, even though their benefits from the program would be high.
-We analyze the overlaps and the characteristics of selected candidates in Appendix M.
-6.3 Simple priority rules
-Optimal assignment policy increases the average probability of getting a tech job across all applicants compared to the status quo.
-However, a non-parametric policy requiring constraint integer optimization is difficult to implement in Dare IT’s context.
-In this section, we examine whether a simple priority rule based on heuristics can enhance the average effectiveness of Dare IT’s programs.
-Consider the admission policy wherein Dare IT first screens participants based on their priority level and then randomly assigns them to either the Mentoring or Challenges program (with the probabilities proportional to the number of spots in either of the programs).
-We propose the following priority groups: Priority 1 – Applicants who are from small towns, do not have a degree in a STEM field, and do not have a master’s or postgraduate degree; Priority 2 – Applicants who are from small towns, do not have a STEM degree, but have a master’s or postgraduate degree; Priority 3 – Applicants who are not from small towns and do not have a STEM degree; Priority 4 – Applicants who have a STEM degree.
-The proposed priority groups are non-overlapping and together encompass all applicants.
-Thus, under this proposed admission policy, Dare IT would first consider the applicants in the first group.
-If there is more capacity after admitting the applicants in the first group, they will move to the second group, and so on.
-We are interested in evaluating whether the groups of applicants with high priority have higher predicted treatment effects than groups with lower priority levels and whether the high-priority groups are large enough to fill in all the existing spots in the two programs.
-Given that this policy is defined using heuristics instead of data, we do not split the dataset for policy estimation and separate evaluation.
-Instead, we use the entire dataset to evaluate this pre-defined policy.
-To estimate the value of the policy, we consider both experimental populations and the remaining applicants.
-For each applicant, we estimate outcomes under three conditions: Mentoring, Challenges, and Out of Dare IT.
-We do that using cross-fitting as described in Appendix J.1.
-The outcome model is the proportional hazard Cox model, and the propensity model is the multinomial logit model; we combine them to obtain AIPW estimates.
-Admitted applicants are assigned with probability 45% to Mentoring and 55% to Challenges (reflecting the proportions of treated subjects in the dataset).
-Once we have the allocation, we compute the average treatment effect per group using the AIPW predictions.
-Figure 5 shows the results.
-Figure 5: Evaluation of simple policies.
-Note: Estimates of the average treatment effect per priority group; AIPW estimates obtained using cross-fitting.
-The treatment effect is the percentage point increase in the probability of a tech job.
-The red dotted line is the average treatment effect across the entire population.
-Whiskers around point estimates represent confidence intervals.
-Standard errors are clustered at the participant level.
-We find that groups 1 and 2 have predicted average treatment effects that are higher than the average across the entire population of applicants (red scatter line) and higher than groups 3 and 4.
-Furthermore, group 1 encompasses 14% of applicants and group 2 10%.
-The combined capacity of Mentoring and Challenges is 28%.
-Thus, the two groups can almost fill in all the spots in the two programs.
-We find that this simple priority rule can increase Dare IT’s effectiveness.
+Mentoring - Challenges -2.431 (0.106) 1.259 (0.085) 0.297 (0.032)
+Note: ATEs are cross-fitted AIPW estimates from causal survival forests (RMST changes up to the evaluation horizon).
+Negative values indicate earlier job finding (months sooner) relative to out of Dare IT programs.
+Standard errors in parentheses are GRF mean-of-variance plug-in SEs aggregated within arm.
+Policy value across capacity levels.
+When comparing the average treatment effects under targeted assignments and the random assignment, we find that there are subsets of applicants who benefit substantially from targeting.
+However, when the capacity of the two programs increases, the additional participants who get in have lower treatment effects, as suggested by the comparison of targeted policies in Table 6.
+Thus, there are diminishing returns to targeting additional slots.
+To identify the level of capacity that realizes most of the available gains from targeting, we estimate the outcomes from targeted assignment under a sequence of capacity values and compare the resulting average probability of getting a job in technology with this outcome under random assignments to Mentoring and to Challenges.
+Figure 4: Policy Value Across Capacity Levels: Targeted vs Random Allocation
+Notes: Average treatment effects under targeted and random allocation policies as capacity varies from 5% to 50% per program.
+The solid curve shows targeted allocation that assigns individuals with highest predicted treatment effects to each program; the dashed line shows random assignment of 50% to each program.
+Treatment effects measured as change in months to employment (negative = faster job finding).
+Shaded region indicates 95% confidence interval for targeted policy.
+Policy value evaluated using cross-fitted AIPW estimates on test set.
+At each capacity level, targeted allocation maximizes expected policy value by solving the assignment problem in Equation 2.
+Figure 4 presents the capacity-selectivity tradeoff by plotting average treatment effects as program capacity varies from 5% to 50% per program.
+The solid curve shows targeted allocation; the dashed horizontal line represents random assignment.
+Several patterns emerge.
+First, targeted allocation consistently outperforms random allocation across all capacity levels, with gains largest when capacity is most constrained.
+At 15% capacity per program, the targeted policy achieves treatment effects of 4.3-4.3 months compared to random allocation’s 2.3-2.3 months, a 2-month improvement.
+Second, the targeted policy exhibits sharply diminishing returns as capacity expands: per-person effects decline from approximately 4.5-4.5 months at 5% capacity to 2.9-2.9 months at 50% capacity.
+This reflects the algorithm’s sequential assignment of highest-benefit individuals first, with marginal participants exhibiting progressively smaller treatment effects.
+Third, even at universal coverage (50% per program, accommodating all applicants), targeted allocation maintains a 0.6-month advantage over random assignment—demonstrating that complementarity between individual characteristics and program type generates value beyond pure capacity expansion.
+6.3 Counterfactual Mentors’ Selection Policies
+Applicants selected by mentors and randomly assigned to the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2).
+This suggests that mentors identify promising applicants who are likely to secure a tech job regardless of their participation in Mentoring.
+Table 8 compares characteristics of selected and non-selected applicants (panel A), and selected by different groups of mentors (panels B and C).
+Baseline is the share of control group applicants that found a job.
+Table 8: Differences in Applicant Characteristics: Selection and Mentor Experience
+Baseline Small town Warsaw Over 30 Prof. exp. Social science STEM Applied before Mother High skill Family/friends in IT
+Panel A: Selected vs. Not Selected
+Selected 0.456 0.245 0.268 0.487 7.097 0.134 0.510
+(0.041) (0.025) (0.026) (0.029) (0.284) (0.020) (0.029) (–) (–) (–) (–)
+Not 0.286 0.476 0.282 0.147 7.489 0.207 0.182
+(0.021) (0.023) (0.021) (0.016) (0.288) (0.019) (0.018) (–) (–) (–) (–)
+Panel B: Mentor Managerial Experience
+High 0.378 0.229 0.281 0.588 7.595 0.137 0.497 0.261 0.288 0.614 0.699
+(0.057) (0.034) (0.036) (0.040) (0.417) (0.028) (0.041) (0.036) (0.037) (0.039) (0.037)
+Low 0.549 0.266 0.245 0.378 6.503 0.133 0.531 0.210 0.231 0.497 0.699
+(0.059) (0.037) (0.036) (0.041) (0.384) (0.028) (0.042) (0.034) (0.035) (0.042) (0.038)
+Panel C: Mentor Total Experience (above vs. below median)
+Above 0.390 0.239 0.299 0.590 8.197 0.171 0.564 0.282 0.308 0.573 0.718
+(0.064) (0.040) (0.043) (0.046) (0.442) (0.035) (0.046) (0.042) (0.043) (0.046) (0.042)
+Below 0.512 0.251 0.240 0.419 6.330 0.112 0.480 0.207 0.229 0.547 0.687
+(0.054) (0.033) (0.032) (0.037) (0.364) (0.024) (0.037) (0.030) (0.031) (0.037) (0.035)
+Notes: Means shown with standard errors in parentheses.
+Baseline refers to the average outcome when not participating in Mentoring.
+Panel A compares applicants who were selected vs.
+not selected.
+Panel B compares applicants selected by mentors with vs.
+without managerial experience.
+Panel C compares applicants selected by mentors above vs.
+below the median years of experience.
+Panel A of Table 8 reveals systematic differences between applicants selected by mentors and those not selected.
+Selected applicants exhibit substantially stronger baseline employment prospects (0.456 vs 0.286), despite having similar years of professional experience (7.1 vs 7.5 years).
+Selected applicants are more likely to be over 30 years old (0.487 vs 0.147) and to hold STEM degrees (0.510 vs 0.182), while being considerably less likely to come from small towns (0.245 vs 0.476).
+They are also less likely to have social science backgrounds (0.134 vs 0.207).
+Panel B compares mentors with versus without managerial experience.
+Mentors with managerial experience select applicants with weaker baseline employment prospects (0.378 vs 0.549) but who are substantially older (0.588 vs 0.378) and have more professional experience (7.6 vs 6.5 years).
+Panel C shows that mentors with above-median total career experience exhibit similar but more pronounced patterns.
+They select applicants with weaker baselines (0.390 vs 0.512) who are older (0.590 vs 0.419), have substantially more professional experience (8.2 vs 6.3 years), and are more likely to hold a STEM degree (0.564 vs 0.480).
+The consistency across Panels B and C suggests a common selection philosophy among more experienced mentors: they look beyond current employability to identify "high-potential, high-barrier".
+Less experienced mentors appear to select more conventionally, favoring applicants already demonstrating stronger baseline employability.
+We use our off-policy tools to examine alternative selection strategies that mentors could adopt and their potential implications for the mentoring program’s effectiveness.
+Specifically, we consider four scenarios: (i) Highest CATE—targeting applicants predicted to have the highest treatment effects, (ii) Most Promising—targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising—selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated—selecting participants with the highest predicted outcomes if they participate in Mentoring.
+Additionally, we include two benchmarks: mentors selecting at random (Random Choice) and the observed allocation (Observed).
+In each of the four policies, we assume that 15% of applicants participate in Mentoring, which is an approximate share of participants to applicants observed by DareIT in recent years.
+Table 9 examines the extent to which different allocation policies select the same individuals Each cell shows the percentage of individuals in the row policy who also appear in the column policy.
+First, we consider the relationship between mentor-based selection (Current) and the Highest CATE algorithmic policy.
+Among current mentor-selected participants, 70.4% would also be assigned to Mentoring under the Highest CATE algorithm, demonstrating substantial concordance between expert judgment and data-driven targeting (note that we do not use the information on whether a candidate was selected or not in the policy estimation).
+Current mentor selections align most closely with Highest CATE (70.4%), followed by Most Promising (29.6%), with little overlap with Least Promising (18.3%) or Highest Chances (19.7%).
+This pattern suggests mentors implicitly optimize for treatment effects rather than baseline characteristics or final outcomes, though with some cream-skimming tendency.
+The 29.6% overlap with Most Promising indicates approximately one-third of mentor selections favor strong baseline candidates, while the predominant 70% overlap with Highest CATE confirms that mentors primarily target based on perceived program responsiveness.
+The low overlap with Least Promising (18.3%) suggests mentors, whether by design or implicit bias, rarely select individuals with the weakest baseline labor market attachment, even when such individuals might benefit from treatment.
+Table 9: Overlap Between Allocation Policies
+Policy H. CATE Most Prom. Least Prom. H. Chances Current
+Highest CATE 100.0
+Most Promising 28.5 100.0
+Least Promising 18.2 0.0 100.0
+Highest Chances 27.9 26.7 31.5 100.0
+Current 70.4 29.6 18.3 19.7 100.0
+Notes: Each cell shows the percentage of individuals in the row policy that also appear in the column policy.
+Only lower triangle shown to avoid redundancy; upper triangle omitted.
+Sample sizes in parentheses after policy names.
+Diagonal entries are 100% by construction (bold).
+H.
+CATE = Highest CATE; H.
+Chances = Highest Chances; Most/Least Prom.
+= Most/Least Promising.
+Table 10 compares treatment effects across six allocation policies for the Mentoring program.
+The results demonstrate that targeting based on predicted treatment effects substantially outperforms alternative strategies.
+The Highest CATE policy achieves treatment effects of 3.59-3.59 months (SE = 0.033), the largest magnitude across all policies.
+Random allocation serves as the baseline benchmark, generating treatment effects of 2.20-2.20 months (SE = 0.040).
+Current mentor-based selection performs remarkably well, achieving 3.25-3.25 months (SE = 0.100), a 48% improvement over random allocation.
+This validates that mentors successfully identify high-benefit participants through their evaluative selection process.
+Table 10: Allocation Policy Comparison: Mentoring Program
+Policy Baseline Time Treatment Effect
+(months) (months)
+Current 11.09 (0.131) -3.25 (0.100)
+Random 11.36 (0.059) -2.20 (0.040)
+Data-driven policies:
+Highest CATE 11.08 (0.086) -3.59 (0.033)
+Most Promising 9.71 (0.044) -2.27 (0.075)
+Least Promising 13.44 (0.085) -1.94 (0.069)
+Highest Chances 11.77 (0.152) -2.14 (0.079)
+Notes: Comparison of allocation policies for the Mentoring program.
+Baseline Time is the predicted months to employment without treatment (lower is better); Treatment Effect is the predicted change in months from treatment (more negative = faster employment).
+Standard errors in parentheses.
+Current reflects mentor-selected participants.
+Random is random assignment benchmark.
+Highest CATE maximizes treatment effects; Most/Least Promising target best/worst baseline outcomes; Highest Chances minimizes post-treatment time to employment.
+Notably, the Highest CATE algorithmic policy achieves an additional 10.4% improvement over mentor selection, increasing treatment effects from 3.25-3.25 to 3.59-3.59 months.
+We obtain the 0.34-month gain using only six readily-observable characteristics (education, location, age, experience) while mentors access rich soft information from interviews and holistic assessment.
+The overlap analysis (Table 9) reveals that 70.4% of mentor-selected participants would also be assigned under Highest CATE, demonstrating substantial concordance between expert judgment and data-driven targeting.
+The finding that simple observable characteristics can meaningfully improve upon expert selection—even when experts demonstrate strong performance—underscores the value of systematic data-driven approaches as complements to human judgment.
+Policies optimizing objectives other than treatment effects yield substantially weaker results.
+Most Promising, which targets applicants with the strongest baseline prospects (9.71 months to employment), achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation despite successfully cream-skimming the highest-potential candidates.
+This demonstrates that strong baseline prospects do not predict strong treatment responsiveness.
+Conversely, Least Promising targets individuals with weakest baselines (13.44 months) but generates the smallest treatment effects (1.94-1.94 months), revealing an unfortunate equity-efficiency tradeoff: those who need help most benefit least from the program.
+Highest Chances, which optimizes final employment outcomes, achieves 2.14-2.14 months, marginally worse than random assignment.
+These comparisons underscore that maximizing program impact requires explicitly targeting predicted treatment effects rather than baseline characteristics or final outcomes, as treatment response heterogeneity does not align with baseline heterogeneity.
 7 Conclusion
@@
 This way we identify a class of interventions that can effectively narrow the gender gap in technology.
-We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by more than 40% four months after the end of the program.
+We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by over 30% twelve months after the end of the program.
+We estimate that Mentoring increases technology employment from 46% to 61%, while Challenges increases it from 27% to 38%, with treatment effects persisting over the full observation period.
 Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
@@
 Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring.
+These differential patterns provide evidence on the underlying mechanisms: Mentoring expands professional networks (increasing LinkedIn connections by 52.7, a 23% increase), while Challenges provides credible skill signals.
+Notably, network expansion occurs only in Mentoring, and heterogeneity patterns align with each program’s mechanism: those with weaker networks benefit most from Mentoring, while those with weaker credentials benefit most from Challenges.
 For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit.
@@
 Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness.
-Specifically, we estimate a seven percentage points increase in the average probability of getting a job in the technology sector across all applicants.
-Notably, such targeted policy utilizing the current combined 28% capacity of both programs, yields outcomes comparable to policies that accommodate all applicants into either Mentoring or Challenges.
-This showcases that in this context targeting admission has a very high potential impact.
+Specifically, we estimate that optimal targeting achieves outcomes 86% larger than random assignment at current capacity levels (30% of applicants).
+Notably, the value of targeting persists even as capacity expands: at 50% capacity per program, targeted allocation maintains outcomes 23% larger than random assignment.
+This demonstrates that targeting generates value both through selecting who participates and through matching individuals to their highest-benefit program.
+The complementarity between individual characteristics and program type creates allocation gains beyond simple capacity expansion.
 A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees.
 To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs.
-This resulted in three categories: applicants not chosen by mentors, those shortlisted but placed in the control group through randomization, and those shortlisted and included in the treatment group.
-Importantly, all applicants underwent an assessment of their technical skills before being presented to mentors.
-By the end of our observation period, there was a 20 percentage point difference in securing tech jobs between those selected but placed in control and the non-selected applicants, with the latter securing job in technology at the probability of 22%.
-This underscores that mentors lean towards candidates they deem more promising.
-While this approach boosts the success rate of Mentoring, reflected in the proportion of participants landing tech jobs, it does not necessarily maximize impact since many of these candidates might have secured tech roles even without Mentoring.
-Through an off-policy evaluation, we explored potential alternative mentor selection strategies.
-Our findings suggest that mentors focusing on participants with the greatest treatment effects, rather than those already poised for success, can significantly amplify the efficacy of Mentoring.
-We estimate that when mentors select 40% of participants based on the highest treatment effect, the share of applicants eventually finding a job in technology is 31% compared to 21% when they choose participants that have the highest outcomes if treated, and 24% when they select participants at random.
+Comparing selected-but-control applicants to non-selected applicants reveals that mentors successfully identify promising candidates: mentor-selected participants achieve outcomes 48% larger than random allocation.
+However, mentors appear to prioritize candidates with strong baseline prospects, those likely to secure technology employment even without the program, rather than those who would benefit most from participation.
+We evaluate an alternative algorithmic allocation policy that selects participants based on predicted treatment effects using only six readily-observable characteristics (education, location, age, experience).
+Despite mentors’ rich soft information from interviews and holistic assessment, algorithmic targeting achieves outcomes 63% larger than random allocation, representing an additional 11% improvement over mentor judgment.
+This finding is particularly striking given that algorithmic and mentor-based selections exhibit 70% concordance: the algorithm largely validates expert choices while identifying additional high-benefit candidates.
+The results demonstrate that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity, and that simple demographic targeting can complement human judgment to enhance program effectiveness.
 This study comes with certain limitations worth noting.
@@
 In Figure 5 we show the balance of covariates in the mentoring experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances.
-The only difference that is statistically significant is the graduate degree, where 66% of treated subjects have a graduate degree and 55% of control (p-value 0.05).
 Figure 5: Covariate Balance Mentoring.
@@
 In Figure 6 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances.
-None of the differences is statistically significant.
+The only variable that has a statistically significant differnece is social science.
 Figure 6: Covariate Balance Challenges.
@@
 Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
-Appendix C Short-term average treatment effects estimates from the AIPW estimator
-In Table 7, we present the estimates of the average treatment effects using the Augmented Inverse Propensity Weighing estimator (Robins et al., 1994).
-AIPW is a doubly-robust method which means that it adjusts for covariates in the outcome model and the propensity score.
-We use the grf implementation of the AIPW estimator, as in Athey et al.
-(2019).
-We use the full set of covariates as controls.
-We find that the estimates from the AIPW estimator are very similar to the estimates obtained using the difference-in-means estimator.
-Table 7: Average treatment effect estimates.
-Mentoring Challenges
-Tech job 4 months New job Tech job 4 months New job
-ATE AIPW 0.129 0.037 0.088 0.0437
-(0.05) (0.06) (0.04) (0.04)
-ATE 0.129 0.045 0.089 0.040
-(0.05) (0.05) (0.04) (0.05)
-Mean treatment 0.421 0.467 0.284 0.333
-(0.04) (0.04) (0.03) (0.03)
-Mean control 0.293 0.422 0.196 0.293
-(0.04) (0.04) (0.03) (0.03)
-No. of obs. treatment 152 152 183 183
-No. of obs. control 147 147 225 225
-Note: The estimates of the average treatment effects of both the mentoring program and Challenges on the probability of finding a new job in technology and finding any new job.
-The first row shows the estimates using the augmented inverse propensity weighting (AIPW) estimator.
-The second row shows estimates of the average treatment effect using the difference-in-means estimator.
-Rows three and four present mean outcomes in treatment and control groups and rows five and six the number of observations per experimental group.
-The first and second columns present the results of the mentoring intervention on the probability of finding a job in technology and the probability of finding any new job.
-Columns three and four show the results of the impact of participation in Challenges on the probability of finding a job in technology and any new job.
-Standard errors are presented below the average treatment effect estimates.
-The variance of the difference-in-means estimates the average treatment effect of mentoring estimated using equation 2.
-The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator.
-AIPW variance estimated using grf implementation of AIPW estimator (see Athey et al.
-(2019) section 4 for details).
-Appendix D Long term treatment effects
-In this, section we present the estimates from two Cox proportional hazard models for both Mentoring and Challenges.
-First, we estimate a model with a full set of controls and second, we consider a model where the treatment variable has time-varying coefficients and the full set of controls.
-Table 8 shows the results.
-We estimate the coefficient of 0.26 in the case of Mentoring, which corresponds to a 29% increase in the probability of having a tech job.
-We find higher treatment effects in the case of Challenges, a coefficient of 0.37, corresponding to a 45% increase.
-We also consider the model in which we analyze the impact of participation in Dare IT programs across time - columns 2 and 4 of Table 8.
-We confirm the observation from Figure 2, treatment effects (especially in the case of Challenges) are realized predominantly after the end of the two programs.
-Furthermore, we do not observe attenuation of the treatment effect over time.
-Table 8: Long-term treatment effects
-Dependent variable: tech job
-Mentoring Mentoring time-varying Challenges Challenges time-varying
-Treated ${}^{***}$ (0.0501) ${}^{***}$ (0.065)
-Treated:during program 0.196 (0.130) --0.074 (0.239)
-Treated: 4 months after the program ${}^{***}$ (0.071) ${}^{***}$ (0.094)
-Treated: 5 months plus after the program ${}^{***}$ (0.082) ${}^{***}$ (0.096)
-Full set of controls x x x x
-Observations 4,275 4,275 5,490 5,490
-Note: Selected coefficient estimate from the proportional hazard models.
-The first column is the mentoring experiment.
-The second column is the mentoring experiment with time-varying coefficients.
-The third column is the challenges experiment.
-In the last column, the challenges experiment with time-varying coefficients.
-In Appendix G, we present the Cox proportional hazard model diagnostics.
-Appendix E Comparison of salary levels from Glassdoor
+Appendix C Kaplan Meier Survival Curves
+We present Kaplan Meier estimates of the survival function $S(t)$, defined as the probability of not yet obtaining a technology job by month $t$.
+Figure 7 shows the survival functions for the Mentoring program.
+Relative to both the control group and applicants, participants in the Mentoring treatment group find technology jobs more rapidly: their survival curve falls more steeply, indicating earlier entry into technology jobs.
+The difference between treated and control groups emerges within the first few months and grows steadily over the 15-month horizon.
+We can also notice that there is a substantial gap between the applicants and both experimental groups which emerges already during the first four months,
+Figure 7: Survival Curves Mentoring Program.
+Note: Kaplan Meier survival curves for the Mentoring program, including treated participants, control group, and applicants. A steeper decline indicates earlier job placement.
+Figure 8 presents analogous estimates for the Challenges program.
+Treated participants also transition into jobs more quickly than the control group or applicants, though the gap is smaller in magnitude than in the Mentoring program.
+Figure 8: Survival Curves Challenges Program.
+Note: Kaplan Meier survival curves for the Challenges program, including treated participants, control group, and applicants.
+Treated participants transition to jobs earlier, though the effect is smaller than in Mentoring.
+Appendix D Comparison of salary levels from Glassdoor
 We consider two types of jobs as a job in technology:
@@
 Note: Histograms of salary estimates from Glassdoor. In green, we show tech jobs in tech firms (category 1), in blue tech jobs in non-tech firms (category 2).
-Appendix F Outcomes measured in surveys
+Appendix E Outcomes measured in surveys
 We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes.
@@
 The difference is statistically insignificant.
-Appendix G Proportional hazards assumption
+Appendix F Proportional hazards assumption
 The Cox proportional hazards model assumes the hazard rate of an event is a product of a baseline hazard and an exponential function of the covariates.
@@
 Prof experience is the number of years between the start date of the first employment of program participants and 2023.
-Appendix H Bounds on the average treatment effect
-In this section, we perform three robustness checks on the average treatment effects.
-First, we consider the AT bounds method introduced by Lee and Weidner (2021).
-Second, due to changes in LinkedIn profile URLs we could not collect outcome data from 17 Challenges participants.
-To obtain bounds on the treatment effects, we consider two extreme cases - none of the 17 subjects found a job in technology, and all of them did.
-We estimate treatment effects in both cases.
-Lastly, we had four participants drop out of the Mentoring program because they secured jobs in the tech industry before the start of the program.
-We establish bounds based on this information, assuming two distinct scenarios: (i) all corresponding control group participants also found jobs in technology, and (ii) none of them did.
-Lee and Weidner (2021) proposes bounds on the average treatment effects that are designed to be robust in situations where there are many conditioning variables, or they have many values, or when the overlap condition is violated.
-We use R implementation of their algorithm (Lee and Weidner, ).
-In the case of Mentoring we consider the following variables: Social science, STEM, Family or Friends in IT, Mother, Applied Before, First-time mentor, Managerial Experience, Career changer, Mentor’s Years of Experience, Professional Experience.
-We estimate the bounds on the average treatment effect to be 0.11 to 0.18 with a confidence interval of 0.03 to 0.26.
-In the case of Challenges, we consider the following characteristics: social science, STEM, Specialization, Residence, Age, Professional experience, Degree.
-We obtain bounds of 0.10 to 0.12 with a confidence interval of 0.02 to 0.20.
-Next, we address the unavailability of LinkedIn data for 17 Challenges participants.
-We explore two extreme scenarios: one where all 17 did not secure tech-sector jobs and another where all 17 did.
-We estimate the average treatment effect in these scenarios using a difference-in-means estimator, without controlling for observable features due to the lack of LinkedIn data.
-When we assign zero outcomes for the missing subjects, the estimated average treatment effect is 6.6 percentage points, with a confidence interval ranging from -1 to 15 percentage points.
-Assuming positive outcomes for all missing variables, we estimate the average treatment effect at 15 percentage points, with a confidence interval spanning 7 to 20 percentage points.
-Finally, we conduct a robustness check of our evaluation of the Mentoring experiment by examining two scenarios: (i) all four control group subjects also found tech jobs, and (ii) none of them did.
-We estimate treatment effects using the difference-in-means estimator.
-In the pessimistic scenario, the estimated treatment effect is 12 percentage points, with a confidence interval of 2 to 23 percentage points.
-In the optimistic scenario, we estimate the treatment effect at 15 percentage points, with a confidence interval spanning 4 to 26 percentage points.
-Appendix I Heterogeneous treatment effects
-In Table 10 we show estimates from the Cox model including the full list of interactions.
-Table 10: Heterogeneous treatment effects of the interventions.
-Dependent variable: tech job
+Appendix G Bounds on the Average Treatment Effect Estimates
+In this section, we perform robustness checks on the average treatment effect estimates.
+First, we consider several cases in which we change which subjects are included in the analysis.
+Second, we estimate AT bounds following lee2021bounding.
+Third, we estimate the average treatment effects at 4 months after the programs.
+Robustness to Sample Construction and Attrition.
+Our main analysis excludes participants who secured technology employment between randomization and program launch, as well as mentors who declined to participate in the experimental design.
+While these exclusions yield conservative estimates, we verify that our conclusions are robust to alternative sample definitions.
+Table 12 presents treatment effect estimates under seven alternative specifications.
+For each, we report sample sizes, mean outcomes by treatment status, and the estimated average treatment effect using difference-in-means without covariate adjustment.
+All estimates use the twelve-month tech employment outcome.
+We consider four specifications for Mentoring.
+First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where both candidates were assigned to treatment due to control participants securing early employment.
+Second, restricting to the 126 perfectly matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3).
+Third, including the 4 participants who found jobs before program launch—coding their outcomes as 1—increases the estimated effect to 16.4 percentage points.
+Fourth, additionally including the 9 non-participating mentors under the assumption their hypothetical mentees would not have found jobs reduces the estimate to 13.0 percentage points.
+All four specifications yield treatment effects between 11.9 and 16.4 percentage points, each statistically significant at the 5% level.
+For Challenges, we examine three specifications.
+Our baseline estimate (11.3 percentage points, SE 4.8) excludes 17 treatment participants who secured technology jobs before program launch.
+Including these individuals and coding their outcomes as 1 increases the effect to 17.1 percentage points—as expected, since we add successful job-finders to the treatment group.
+Finally, we address the 17 treatment and 7 control participants with missing twelve-month LinkedIn data.
+Under the conservative assumption that all missing treatment observations had outcomes of 0 and all missing control observations had outcomes of 1—the scenario most unfavorable to finding a treatment effect—we still estimate a positive effect of 10.5 percentage points (SE 4.6, p=0.02).
+Table 12: Robustness to Sample Construction and Attrition
+Specification Sample Size Mean Outcome ATE
+Treated Control Treated Control (SE)
+Panel A: Mentoring
+   Main specification 151 147 0.609 0.456 0.153***
+   Balanced pairs only 126 126 0.603 0.484 0.119*
+   Include pre-program employment 155 147 0.619 0.456 0.164***
+   Include non-participants 164 147 0.585 0.456 0.130**
+Panel B: Challenges
+   Main specification 166 218 0.380 0.266 0.113**
+   Include pre-program employment 183 218 0.437 0.266 0.171***
+   Pessimistic attrition bounds 200 227 0.400 0.295 0.105**
+Note:Estimates of average treatment effects on twelve-month technology employment under alternative sample definitions.
+Main specifications correspond to Table 2.
+Balanced pairs only (Mentoring) restricts to the 126 mentor pairs with both treatment and control observations.
+Include pre-program employment adds participants who secured technology jobs between randomization and program launch.
+Include non-participants (Mentoring) adds 9 mentors who declined experimental participation, assigning their hypothetical mentees outcomes of 0.
+Pessimistic attrition bounds assign outcomes of 0 to 17 missing treatment observations and 1 to 7 missing control observations.
+All estimates use difference-in-means without covariate adjustment.
+Standard errors in parentheses.
+*** p<0.01, ** p<0.05, * p<0.1.
+These robustness checks establish three key points.
+First, our main estimates are insensitive to reasonable alternative sample construction decisions, with all specifications yielding economically large and statistically significant treatment effects.
+Second, the range of point estimates across specifications (11.9 to 16.4 for Mentoring, 10.5 to 17.1 for Challenges) provides additional bounds on the true treatment effects that are consistent with our main findings.
+Third, even under extremely conservative assumptions about missing data—assumptions, the Challenges program shows significant positive impacts.
+This addresses potential concerns that selective attrition might bias our twelve-month estimates.
+Lee Bounds for Selective Attrition.
+Our twelve-month outcome measurement experienced modest attrition in the Challenges program: 6% had missing LinkedIn data due to profile URL changes or privacy setting modifications.
+We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of lee2009training.
+We consider the most conservative case, in which subjects that found jobs before the start of the program are not considered; thus, we focus only on the attrition during the program, due to missing LinkedIn data.
+The Lee bounding procedure addresses selective attrition by considering worst-case scenarios about which individuals are missing.
+Since treatment experienced higher attrition, the conservative assumption for estimating positive effects is that the best-performing treatment participants were most likely to attrite (e.g., successful job-finders who deactivated LinkedIn).
+To construct the lower bound, we trim 6.8% of observations from the upper tail of the observed treatment outcome distribution, the proportion representing excess attrition relative to control, and compare the mean of this trimmed distribution to the full control group mean.
+The upper bound uses all observed treatment outcomes compared to all observed control outcomes.
+We estimate Lee bounds of [0.065, 0.114].
+The upper bound of 0.114 is nearly identical to our main point estimate of 0.113, while the positive lower bound of 0.065 indicates that even under worst-case assumptions about selective attrition, the Challenges program effect remains economically meaningful.
+The 95% confidence interval on the lower bound is [-0.029, 0.159]; while this marginally includes zero due to sampling variability in the observed outcomes, the point estimate is bounded away from zero.
+Average Treatment Effect Estimates at 4 Months.
+Our main analysis uses employment outcomes measured twelve months post-program.
+However, we were unable to collect LinkedIn data for 17 Challenges treatment participants and 7 control participants at the twelve-month measurement due to profile URL changes or privacy setting modifications.
+A natural concern is whether selective attrition biases our estimated treatment effects.
+To address this concern, we re-estimate average treatment effects using employment outcomes measured four months after program completion.
+At this earlier measurement point, we successfully collected complete LinkedIn data for all randomized participants in both programs.
+Any attrition-induced bias is therefore eliminated at this time horizon.
+Table 13 presents the results.
+We find substantively similar treatment effects at the four-month horizon compared to our main twelve-month estimates (reported in Table 2).
+For Mentoring, the four-month treatment effect is 15.3 percentage points (SE 5.7), nearly identical to the twelve-month effect of 14.5 percentage points.
+For Challenges, the four-month effect is 8.9 percentage points (SE 4.3), compared to 11.2 percentage points at twelve months.
+The larger point estimate at twelve months suggests that the Challenges program’s impact may strengthen over time as participants leverage their portfolios in job applications.
+Table 13: Average Treatment Effects of Mentoring and Challenges at 4 Months.
 Mentoring Challenges
-Treatment ${}^{***}$ (0.265) ${}^{***}$ (0.239)
-Social science --0.124 (0.125)
-STEM ${}^{***}$ (0.099) ${}^{***}$ (0.111)
-Family friends IT ${}^{***}$ (0.092)
-Small Town $*{}^{*}$ (0.102) $*{}^{*}$ (0.126)
-Warsaw 0.088 (0.094) ${}^{***}$ (0.137)
-Mother ${}^{***}$ (0.137)
-Applied before ${}^{***}$ (0.096)
-First time mentor 0.123 (0.084)
-Managerial experience ${}^{***}$ (0.087)
-Career changer ${}^{***}$ (0.099)
-Years of experience mentor --0.003 (0.013)
-Non-traditional credential ${}^{***}$ (0.188) ${}^{***}$ (0.179)
-Masters ${}^{***}$ (0.171) 0.218 (0.135)
-No education mentioned ${}^{**}$ (0.193) $*{}^{*}$ (0.287)
-Postgraduate $*{}^{*}$ (0.188) ${}^{***}$ (0.194)
-Age: 25-30 $*{}^{*}$ (0.145) ${}^{**}$ (0.155)
-Age 31-35 ${}^{***}$ (0.168) 0.170 (0.171)
-Age: 36+ ${}^{**}$ (0.198) --0.144 (0.184)
-Prof. experience ${}^{***}$ (0.013) 0.008 (0.011)
-Treatment×$\times$×Social science 0.273 (0.179)
-Treatment×$\times$×STEM ${}^{***}$ (0.136) ${}^{***}$ (0.160)
-Treatment×$\times$×Family friends IT $*{}^{*}$ (0.119)
-Treatment×$\times$×Small ${}^{***}$ (0.131) ${}^{***}$ (0.173)
-Treatment×$\times$×Warsaw ${}^{***}$ (0.129) ${}^{***}$ (0.182)
-Treatment×$\times$×Mother ${}^{***}$ (0.164)
-Treatment×$\times$×Applied before --0.155 (0.123)
-Treatment×$\times$×First time mentor --0.079 (0.108)
-Treatment×$\times$×Managerial experience ${}^{***}$ (0.112)
-Treatment×$\times$×Career changer ${}^{***}$ (0.125)
-Treatment:$\times$×Years of experience mentor 0.001 (0.016)
-Treatment×$\times$ Non-traditional credential ${}^{***}$ (0.235) ${}^{***}$ (0.251)
-Treatment×$\times$ masters ${}^{***}$ (0.209) --0.267 (0.192)
-Treatment×$\times$ No education mentioned ${}^{***}$ (0.347) --0.139 (0.349)
-Treatment×$\times$×postgraduate ${}^{**}$ (0.236) ${}^{***}$ (0.243)
-Treatment×$\times$×Age: 25-30 ${}^{***}$ (0.203) ${}^{***}$ (0.209)
-Treatment×$\times$×Age:31-35 ${}^{***}$ (0.241) ${}^{***}$ (0.234)
-Treatment×$\times$×Age: 36+ 0.273 (0.277) --0.013 (0.235)
-Treatment×$\times$×Prof experience ${}^{***}$ (0.016) $*{}^{*}$ (0.015)
-Observations 4,440 4,965
-[…]
-[…]
-Log Likelihood --11,845.370 --6,329.277
-Wald Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
-LR Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
-Score (Logrank) Test ${}^{***}$ (df = 39) ${}^{***}$ (df = 23)
-Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
-Note: Estimates from the Cox proportional hazard model. Outcome variable tech job.
-Appendix J Models used in off-policy sections
-This section presents further details about the propensity and outcome models used in Section 6 for policy estimation.
-To estimate both models, we first join the Mentoring, Challenges, and applicants’ datasets.
-We obtain a dataset of 1314 individuals who applied to Dare IT, out of whom 151 participated in Mentoring and 166 in Challenges; the remaining 997 are the non-selected Mentoring applicants, late applicants to Challenges, and the control groups.
-All individuals are observed over the span of 15 months.
-The dataset is randomly split between training and testing at the individual applicant level, 50-50.
-Propensity model.
-To estimate the propensity model for policy estimation, we use multinomial logit regression with the full set of covariates that overlap between all datasets.
-We train the model in the training sample and predict into the testing set.
-We find that the propensity of being assigned to Mentoring ranges from 4% to 37%, to Challenges from 5% to 32%, and to Out of Dare IT from 53% to 89%.
-Figure 13 presents the histogram of the three propensities.
-Figure 13: Predicted program propensities.
-Note: Test set propensity of assignment across groups, in blue to Challenges, in red to Mentoring, and in green to Out of Dare IT.
-We also evaluate the adjustment with the propensity weights by compare the standardized absolute mean differences in covaraites values with and without the adjustment.
-In Figure 14, we show the difference in average values of covariates between Mentoring and Nothing groups, with and without adjustment in each cross-fitting round.
-The same analysis for the case of Challenges and Nothing is in Figure 15.
-We find that the propensity model successfully adjusts for the differences in covariates.
-Figure 14: Covariate balance: Mentoring.
-Note: Differences in mean covariates across Mentoring and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
-Figure 15: Covariate balance: Challenges.
-Note: Differences in mean covariates across Challenges and Nothing group with adjustment (in yellow) and without (in blue). Each dot corresponds to one round of cross-fitting.
-Outcome model.
-We use Cox proportional hazard model to model outcomes.
-We estimate a model with a full set of covariates and interactions of covariates and treatments.
-We model outcomes at the person and month level.
-In Figure 16, we show an evaluation of the model’s predictive performance.
-First, we train a model on the training sample and predict probabilities of having a job in a specific month using the testing sample; then, we take the average per user and group users based on these averages into quintiles.
-Finally, we compare the predicted probability within the quintile to the observed probability.
-We find that the differences are statistically insignificant, and the magnitudes are small.
-We conclude that the model is well-calibrated.
-Figure 16: Comparison of predicted and observed outcomes.
-Note: Comparison of predicted and observed outcomes in the test set.
-Subjects are grouped into quintiles based on the predicted probability of having a tech job.
-In red, the average per quintile predicted probability.
-In blue, is the average per quintile observed outcome.
-Whiskers show confidence intervals.
-J.1 Cross-fitting.
-We estimate predicted outcomes under a policy using cross-fitting with 10 folds and 100 repetitions.
-Specifically, for each repetition, $m\in\left\{1,...,100\right\}$, we randomly assign observations into 10 folds $k\in\left\{1,..,10\right\}$.
-Randomization is at the individual subject level with equal probability.
-Next, for each fold $k$ we train an outcome and propensity models using data from all other folds and predict outcomes and treatments in the $k$-th fold: $\hat{\mu}_{k}(X_{i})_{P}^{m}$ and $\hat{e}_{k}(X_{i})_{P}^{m}$, where $P\in\left\{0,M,C\right\}$ refers to a program.
-We use these predictions to construct AIPW estimates; for example, the treatment effect from Mentoring instead of Out of Dare IT writes:
-$\tau^{m}_{OM}=\frac{1}{n}\sum_{i=1}^{n}\left[\hat{\mu}_{k}(X_{i})_{M}^{m}+\frac{\mathbb{1}_{W_{i}=M}}{\hat{e}_{k}(X_{i})^{m}_{M}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{M})-\hat{\mu}(X_{i})_{0}^{m}+\frac{\mathbb{1}_{W_{i}=0}}{\hat{e}_{k}(X_{i})^{m}_{O}}\times(Y_{i}-\hat{\mu}_{k}(X_{i})^{m}_{O})\right],$
-analogously for Challenges. Finally, following Chernozhukov et al. (2017), we take the median over 100 repetitions so that $\hat{\tau}_{OM}=Median(\tau^{m}_{0M})$.
-Appendix K Policy estimation algorithms
+   ATE 0.153 0.089
+   SE (0.057) (0.043)
+   ATE % baseline 33.7 45.3
+   SE % baseline (12.6) (21.8)
+   N Treatment 151 183
+   N Control 147 225
+Note: Estimates of the average treatment effect of Mentoring and Challenges at 4 months after the end of each program. Estimates obtained using the difference-in-means estimator.
+Importantly, these four-month estimates use the full experimental sample without any missing observations, confirming that our main findings are not an artifact of selective attrition.
+The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts card2018works.
+If anything, the point estimates suggest persistent or growing effects, particularly for Challenges, though the conclusions are limited to 12 months from the start of each program.
+Appendix H Policy estimation algorithms
 Algorithm 1 describes the estimation and evaluation of optimal policy.
@@
 Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$.
-See Appendix J.1 for details of the cross-fitting procedure.
+See Appendix LABEL:cross_fit_appendix for details of the cross-fitting procedure.
 Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
@@
 return $(\mathcal{A}_{X,Q}^{*},\hat{V}_{X,Q}^{*},\sigma_{X,Q}^{*},\hat{Y}_{1,k}(a^{*}),...,\hat{Y}_{\mathcal{I}^{test},k}(a^{*}))$
-Appendix L Validate AIPW approach
-In Section 6, we used estimated AIPW scores to construct and evaluate counterfactual policies.
-Now, we compare the estimates of treatment effects using the AIPW approach in the testing set with cross-fitting and the experimental results.
-Thus, we want to test the hypotheses that $\tau^{AIPW}=\tau^{RCT}$.
-To estimate $\tau^{AIPW}$, we proceed in the same steps as in Section 6: first, we pool all datasets: Mentoring treatment and control, mentoring non-selected applicants, Challenges treatment and control, challenges late applicants.
-We split the dataset into train and test, and in the testing dataset, we randomly assign 10 folds.
-Then for each fold, we train a proportional hazard model for the outcomes and a multinomial logistic regression for propensities using other folds, and then we predict outcomes and propensities for the left-out fold and construct AIPW scores.
-In the outcome model, we include an indicator variable of whether the applicant was selected by a mentor or not.
-We repeat the process of cross-fitting 100 times and take the average AIPW score per observation.
-To assess Mentoring, we take the difference between predicted AIPW scores under mentoring treatment and under mentoring control, i.e., the estimate of the average treatment effect for the entire sample, analogously for Challenges.
-We compare obtained quantities with the estimates of the treatment effects using Cox models on the experimental samples.
-We also compute the difference between the estimates.
-Figure 17 shows the results
-Figure 17: Difference between AIPW and RCT estimates.
-Note: Comparison of the average treatment effect estimates using the randomized experiment and the AIPW scores. Point estimates and confidence intervals.
-We find that the RCT and AIPW estimates are similar, and the differences are statistically not distinguishable from zero.
-Recall that the randomization into treatment and control was carried out within experimental samples, that is, the sample of Mentoring applicants who mentors selected and the subset of applicants to Challenges that applied to the program before we closed the registration.
-Thus, it was not a priori clear that the AIPW procedure with cross-fitting would recover similar estimates of the treatment effect after we pooled all datasets together.
-The finding that these estimates are very close in magnitude and statistically not different from one another bolsters the credibility of results presented in Section 6.
-Appendix M Mentors’ selection strategies
-In this section, we dive deeper into four mentors’ strategies: Observed, Highest CATE, Highest if treated, and Least promising.
-We analyze the degree to which they overlap in terms of which candidates get selected and how the characteristics of potential mentees differ across these strategies.
-First, we present the Venn diagram in Figure 18.
-We observe considerable overlap between Observed, Highest CATE, Highest if treated; approximately half of the candidates selected for Mentoring under Highest CATE are also selected in the observed allocation.
-Furthermore, almost two-thirds of selections under Highest if treated and Highest CATE overlap.
-Similarly, 100 out of 138 candidates overlap between Highest if trated and Observed selection.
-In contrast, there is low overlap between Least promising and the other three policies, with 124 candidates that are selected under Least promising not being selected under any other policy.
-Figure 18: Venn diagram of selected policies.
-Note: Each policy has 138 selected candidates. Percentages refer to the share of the 138 that fall into the region of the graph.
-Next, we consider the following sets of candidates: (i) Observed selection – all candidates selected under the current allocation, (ii) Only least promising – candidates that are selected under the Least promising policy and are not selected under the remaining three policies, (iii) Highest CATE & Observed selection – the candidates that are selected both under the Highest CATE and under Observed selection, and (iv) Highest CATE but not selected - where we consider candidates that are selected under Highest CATE, but under mentors’ rule.
-We analyze the baseline outcomes, average treatment effects, and means of selected covariates.
-We present results in Table 11
-Table 11: ATEs, baselines, and covariate means and standard errors for different policies
-policies ATE Baseline STEM Prof. exp 30 plus Small town
-Observed selection $0.095(0.012)$ $0.275(0.008)$ $0.457(0.011)$ $7.072(0.110)$ $0.514(0.168)$ $0.232(0.009)$
-Only least promising $0.125(0.008)$ $0.048(0.005)$ $0.076(0.006)$ $6.277(0.160)$ $1.000(0.000)$ $0.588(0.012)$
-Highest CATE & Observed selection $0.158(0.017)$ $0.244(0.011)$ $0.493(0.015)$ $8.587(0.161)$ $0.507(0.015)$ $0.400(0.015)$
-Highest CATE but not selected $0.208(0.016)$ $0.112(0.011)$ $0.365(0.016)$ $10.095(0.168)$ $0.365(0.016)$ $0.873(0.011)$
-Note: Average treatment effects and baselines from Section 6.2 standard errors clustered at the candidate level.
-STEM is the share of candidates selected under the considered policy that have a degree in STEM, Prof.
-exp is the average length of professional experience, 30 plus is the share of candidates that are above 30 years old, and Small town is the share of candidates that reside in a smaller town.
-Comparing candidates selected only under Least promising to other categories, we find that these applicants are less likely to have a STEM degree, have shorter professional experience, and are more likely to be above 30 years old.
-Even though both mentors overlook this group and we do not select these candidates under the Highest CATE policy, the estimated treatment effects for this group are high.
-Furthermore, the treatment effect of 0.125 increase in the probability of getting a tech job is particularly high considering the baseline of only 0.048.
-Finally, comparing the candidates that have high estimated treatment effects and get selected to those with high treatment effects who are not selected, we notice that first, mentors select high CATE candidates with higher baselines and lower average treatment effects.
-They tend not to select enough candidates who have longer professional experience and come from smaller towns.
-These patterns are consistent with our estimates of the heterogenous treatment effects shown in Section 5.2.
+H.1 Assignment Models
+We pool data from Mentoring experimental participants (N=298), Challenges experimental participants (N=408), non-selected Mentoring applicants (N=468), and late Challenges applicants (N=160), yielding 1,334 total applicants.
+All individuals are observed over fifteen months post-application.
+We randomly split this dataset 50-50 at the applicant level into training (N=667) and testing (N=667) sets.
+Using the training set, we estimate predicted treatment effects for each program relative to non-participation.
+We fit two causal survival forests (athey2019generalized; cui2023estimating): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
+Each causal survival forest is estimated with the following parameters: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months.
+Covariates include: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization.
+These settings follow standard recommendations in the generalized random forests (grf) package (tibshirani2024grf).
+The causal survival forest estimates conditional average treatment effects on the restricted mean survival time (RMST) scale—the average time until technology employment over our observation horizon.
+More negative predicted effects indicate faster job finding under treatment.
+We apply the trained models to the test set to obtain predicted treatment effects $\hat{\tau}_{i}^{M}$ (Mentoring vs.
+nothing) and $\hat{\tau}_{i}^{C}$ (Challenges vs.
+nothing) for each test-set applicant.
+Appendix I Evaluation Models
+While assignment models determine which applicants receive which program under various policies, evaluation models estimate the causal effects of these assignments.
+The credibility of our policy comparisons rests on these evaluation models accurately estimating treatment effects for the allocated applicants.
+If evaluation models correctly estimate treatment effects, our policy value estimates remain valid even if assignment models imperfectly rank applicants.
+We use the held-out test set (N=667) for all policy evaluations.
+This sample includes experimental participants and non-participants from both programs, ensuring we can estimate counterfactual outcomes under different assignment policies.
+All individuals are observed over the same fifteen-month post-application horizon.
+Model specification.
+For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (athey2019generalized; cui2023estimating).
+Within each cross-fitting fold (detailed below), we fit two causal survival forests: one comparing Mentoring participants to applicants who did not participate in any program, and another comparing Challenges participants to non-participants.
+Each causal survival forest uses identical parameters to the assignment models: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months.
+The covariate set is identical to that used in assignment models: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization.
+Model fitting follows the grf package implementation (tibshirani2024grf).
+The causal survival forest produces augmented inverse propensity weighted (AIPW) scores $\psi_{i}^{M}$ and $\psi_{i}^{C}$ for each test-set individual, representing individual-level treatment effect estimates from Mentoring and Challenges relative to non-participation, respectively.
+The forest automatically constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a survival forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates.
+These components are combined via the AIPW formula, yielding estimates that remain consistent if either the propensity or outcome model is correctly specified.
+Cross-fitting procedure.
+To avoid overfitting and ensure valid inference, we employ repeated $K$-fold cross-fitting on the test set.
+Specifically, for each repetition $m\in\{1,\ldots,100\}$, we randomly assign observations into $K=5$ folds.
+Randomization is stratified by experimental group (Mentoring treated, Mentoring control, Challenges treated, Challenges control, non-participants) to ensure each fold contains observations from all groups.
+For each fold $k\in\{1,\ldots,5\}$, we:
+  1. 1.
+    Train causal survival forests using data from all folds except $k$ (the training sample for fold $k$).
+  2. 2.
+    Predict AIPW scores for observations in fold $k$: $\psi_{i,k}^{M,m}$ and $\psi_{i,k}^{C,m}$ for each individual $i$ in fold $k$, along with corresponding variance estimates $\hat{V}_{i,k}^{M,m}$ and $\hat{V}_{i,k}^{C,m}$.
+After completing all folds in repetition $m$, each individual has predicted treatment effects $\psi_{i}^{M,m}$ and $\psi_{i}^{C,m}$ estimated on data that did not include that individual.
+We then aggregate across the 100 repetitions: following chernozhukov2017double, we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$.
+Variance estimates are similarly aggregated by taking the median across repetitions.
+Policy value estimation.
+Given an allocation policy $\pi$ that assigns each test-set individual to Mentoring, Challenges, or neither, we estimate the policy value as:
+$\hat{V}(\pi)=\frac{1}{n_{M}}\sum_{i:\pi(i)=M}\psi_{i}^{M}+\frac{1}{n_{C}}\sum_{i:\pi(i)=C}\psi_{i}^{C}$
+where $n_{M}$ and $n_{C}$ are the number of individuals assigned to Mentoring and Challenges, respectively.
+Standard errors are computed using the aggregated variance estimates from the causal survival forests, accounting for the within-arm averaging.
+For comparisons between policies or programs, we compute appropriate differences in AIPW scores and propagate variances assuming independence across programs.
+This approach separates the assignment decision (which applicants receive which program) from the evaluation of that decision (estimating treatment effects for the assigned groups).
+The validity of our policy comparisons depends on the evaluation models accurately estimating treatment effects, not on the assignment models perfectly ranking applicants.
+I.1 Model Validation: Recovering Experimental Effects
+To validate our evaluation model specification, we assess whether cross-fitted AIPW scores correctly recover experimental average treatment effects.
+This diagnostic tests whether our models produce unbiased treatment effect estimates in the setting where we have the benchmark from the randomized experiment.
+Validation procedure.
+We train causal survival forests following the same process as before, using 5-fold cross-fitting and generate out-of-sample AIPW predictions for each participant; we use the experimental samples.
+We compare these individual-level predictions to the experimental ATE by examining the distribution of deviations: $\delta_{i}=\hat{\tau}_{i}^{AIPW}-\hat{\tau}^{RCT}$.
+Under correct model specification, these deviations should center at zero with dispersion reflecting genuine treatment effect heterogeneity plus estimation noise.
+Figure 14 presents the results.
+The red solid vertical line indicates the RCT benchmark (zero deviation), with shaded regions marking 95% confidence intervals around both the RCT estimate (red) and the mean AIPW deviation (gray).
+The black dashed line shows the mean deviation from the RCT estimate.
+Figure 14: Model Validation: Distribution of Deviations from RCT Estimates
+Refer to caption Refer to caption
+(a) Mentoring Program (b) Challenges Program
+Notes: Distribution of deviations between cross-fitted AIPW predictions and experimental ATE benchmarks.
+Each panel shows $\delta_{i}=\hat{\tau}_{i}^{AIPW}-\hat{\tau}^{RCT}$ for experimental participants.
+The red vertical line marks zero deviation (perfect recovery), with red shading indicating the 95% confidence interval around the RCT estimate.
+The black dashed line shows the mean AIPW deviation, with gray shading for its 95% confidence interval.
+Causal survival forests trained exclusively on experimental samples using 5-fold cross-fitting.
+Mentoring: mean deviation = $0.08$ months (SE = 0.03); Challenges: mean deviation = $0.06$ months (SE = 0.04).
+Deviations statistically indistinguishable from zero in both programs.
+For the Mentoring program (Panel a), AIPW predictions center tightly around the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03).
+The distribution is approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE.
+The modest spread in the distribution reflects both the treatment effect heterogeneity among participants and sampling variation in the predictions.
+For the Challenges program (Panel b), AIPW predictions show similarly strong performance, with a mean deviation of $0.06$ months (SE = 0.04).
+While the distribution exhibits slightly more dispersion than Mentoring—consistent with Challenges’ larger experimental sample and potentially greater heterogeneity—the mean deviation remains statistically indistinguishable from zero.
+The slight positive mean suggests modest underestimation, but the bias is economically negligible (less than 2% of the experimental ATE).
+The cross-fitted causal survival forests accurately recover experimental treatment effects, demonstrating that our evaluation approach is well-calibrated and unbiased within the experimental sample.
+This gives us confidence that the models correctly estimate treatment effects where treatment is randomized.

### v5

@@
 Abstract
-Women remain underrepresented in technology despite many support programs, partly due to limited evidence on which interventions work and how to allocate scarce program slots.
-We evaluate two approaches to facilitating technology-sector transitions for women in Poland: Challenges, a scalable and cheap-to-operate (\$15/person) online program developing job-relevant portfolios (designed and implemented during this research), and one-on-one mentoring.
-The programs address distinct labor market barriers: expanding professional networks and teaching their effective use (Mentoring) versus providing credible skill signals (Challenges).
-Randomizing oversubscribed admissions, we find both programs substantially increase technology employment twelve months post-completion, Mentoring by 15 percentage points (p.p.) and Challenges by 11 p.p.
-Exploiting treatment effect heterogeneity, we show that algorithmic targeting based on predicted treatment effects achieves outcomes 86% larger than random assignment at current capacity levels.
-While mentors successfully identify high-benefit candidates (outcomes 48% above random selection), algorithmic allocation using observable characteristics yields an additional 11% improvement.
-The results demonstrate that (i) relatively low-cost interventions can meaningfully address gender gaps in technology, and (ii) data-driven allocation rules can substantially enhance program effectiveness even when expert judgment performs well.
+We evaluate two interventions facilitating technology-sector transitions for women in Poland: Mentoring, focused on expanding professional networks, and Challenges, focused on building credible skill signals.
+Randomizing oversubscribed admissions, we find both programs substantially increase technology employment at twelve months—by 15 percentage points for Mentoring and 11 p.p.
+for Challenges.
+The distinct mechanisms through which the programs operate translate to heterogeneous treatment effects across geography, career stage, and baseline credentials.
+These differential effects create scope for improved allocation: algorithmic targeting across programs outperforms random assignment by 86% and experts’ selection into Mentoring by 11%.
 JEL: J24, J08, J16, C93
 1 Introduction
-An important concern for labor policymakers is identifying and implementing effective programs to support workers transitioning into high-growth sectors.
-The technology sector exemplifies this challenge: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide.111unitednations2021 estimates the share of women in the technology sector at approximately 30%.
+Facilitating worker transitions into high-growth sectors is a core challenge for labor market policy, particularly for underrepresented groups.
+The technology sector is a prime example: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide.
+[^1] United Nations (2021) estimates the share of women in the technology sector at approximately 30%.
 Many programs exist to support women in technology, including NGO initiatives (Women in Technology International, Girls Who Code), corporate programs (Women in Tech from Salesforce, The Women at Microsoft), and government efforts (UN’s Girls in ICT Day, European Commission’s Women in Digital Scoreboard).
-While women face multiple barriers entering technology, this paper examines two specific labor market frictions.
-[^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (reuben2014gender; moss-racusin2012science), work-family conflict and child penalties (goldin2014grand; kleven2019children; bertrand2010dynamics), and underrepresentation in STEM education pipelines (bayer2016diversity; carrell2010sex).
+While women face multiple barriers to entering technology, we focus on two labor market frictions that short-term interventions can directly address.
+[^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (Reuben et al., 2014; Moss-Racusin et al., 2012), work-family conflict and child penalties (Goldin, 2014; Kleven et al., 2019; Bertrand et al., 2010), and underrepresentation in STEM education pipelines (Bayer and Rouse, 2016; Carrell et al., 2010).
 We focus on network and information barriers because short-term labor market interventions can directly address these constraints, unlike structural factors requiring longer-term policy responses.
-First, women have weaker professional networks than men, limiting access to job opportunities (greguletz2019women; forret2004networking).
-Second, gender stereotypes lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (cheryan2013stereotypical; del2022more; murciano2022missing).
-Understanding which constraints bind, and for whom, is critical for designing effective policy.
+First, research across multiple settings has found that women have weaker professional networks than men, limiting access to job opportunities (Greguletz et al., 2019; Forret and Dougherty, 2004).
+Second, gender stereotypes have been shown to lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (Cheryan et al., 2013; Del Carpio and Guadalupe, 2022; Murciano-Goroff, 2022).
+Understanding which barriers pose the greatest impediment, and for whom, is critical for designing effective policy.
 If network barriers dominate, interventions should expand professional connections; if information frictions dominate, scalable credentialing programs may suffice.
-Alternatively, both frictions may bind simultaneously for different populations, necessitating multiple program types and optimal assignment rules.
-We provide experimental evidence distinguishing between these possibilities.
+Alternatively, different populations may face different primary barriers, necessitating multiple program types and targeted assignment rules.
+We find evidence that both network deficits and information frictions matter, but for different populations: network deficits primarily constrain women outside major urban centers and those with caregiving responsibilities, while information frictions primarily constrain early-career women with limited work history.
 We evaluate two interventions implemented in Poland by Dare IT, a social impact organization, each targeting one of these frictions.
 Mentoring pairs participants with senior female technology professionals for four months of one-on-one sessions.
-While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching effective relationship leverage.
+While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching participants how to effectively use these connections.
 Challenges is a scalable online program (\$15 per participant) we designed during this research to address skill signaling barriers.
-The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, mobile app designs or React applications, that participants reference in job applications.
+The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, such as mobile app designs or React applications.
+Participants later reference these projects in job applications.
 [^3] Program development began with structured interviews with hiring managers at twelve Polish technology firms, who uniformly cited the absence of credible practical experience signals in female applicants’ profiles.
@@
 The program has been replicated across multiple specializations (software testing, cloud computing, UI design) and delivered in Ukrainian for refugees.
-We exploit both programs’ substantial oversubscription to randomize access and measure employment outcomes via LinkedIn profiles over twelve months.
-[^4] Technology jobs include positions in technology firms (excluding finance, legal, HR, accounting roles) or technical positions elsewhere (software development, IT support, data analytics).
+Both programs were substantially oversubscribed, allowing us to randomize access and measure employment outcomes via LinkedIn profiles over twelve months.
+[^4] We define technology jobs as positions in technology firms (excluding finance, legal, HR, and accounting roles) or technical positions elsewhere (software development, IT support, data analytics).
 UX and front-end programming roles are classified as technology jobs regardless of firm type.
-Both programs generate large employment effects, but with markedly different patterns of treatment effect heterogeneity that reveal which constraints bind for different populations.
+Both programs generate large employment effects.
 Mentoring increases technology employment by 15 percentage points (SE 5.7, from a 46% control baseline); Challenges by 11 percentage points (SE 4.8, from 27%).
-Evidence supports distinct mechanisms.
-Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
-The programs benefit different populations, with heterogeneity patterns aligning with each mechanism.
+Consistent with distinct mechanisms, Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
+The programs benefit different populations.
 Professional experience reveals the starkest contrast: Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers, while Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced workers.
 This indicates that early-career women benefit from portfolio signals that compensate for limited work history, while experienced women benefit from network expansion that leverages their accumulated professional standing.
-Geography provides additional evidence: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12).
+Geographic heterogeneity supports the network mechanism: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12).
 This pattern is consistent with network barriers binding tightly outside major urban centers where local technology ecosystems are sparse, while online portfolio signals function equally regardless of location.
-Mentor characteristics also predict treatment effects: participants matched to mentors with managerial experience or over five years in technology show larger treatment effects, consistent with senior mentors possessing richer networks and better access to hiring decision-makers.
 These differential patterns demonstrate that both network barriers and information frictions constrain technology employment, but for distinct populations defined by career stage or geography.
 The differential treatment effects create an allocation problem when program capacity is limited.
-Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratios for Mentoring; 300 spots filled in four hours for Challenges), requiring selection among applicants.
-We evaluate counterfactual assignment policies using the methods of athey2021policy and zhou2023offline, estimating which applicants should receive which program to maximize aggregate outcomes.
-A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves outcomes 86% larger than random assignment.
-This gain reflects two sources of value: selecting individuals with highest treatment effects and matching individuals to their optimal program type.
-The matching component generates persistent value: even when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment.
-This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion, treatment effect heterogeneity enables welfare-improving assignment even when most applicants can be served.
-We further compare algorithmic targeting to expert human judgment.
-Mentoring relies on experienced female technology professionals to select mentees from pre-screened applicants based on pre-recorded short videos and the assessment of the application materials (cover letter, resume, and registration survey).
-This expert selection substantially outperforms random allocation, achieving outcomes 48% larger.
-Yet algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain preference) achieves outcomes 63% larger than random assignment, an additional 11% improvement over mentor judgment.
-The two approaches, by experts and by the algorithm, achieve high-degree of concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates.
+Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratio for Mentoring; all 300 spots filled in four hours for Challenges), requiring selection among applicants.
+We evaluate counterfactual assignment policies using the methods of Athey and Wager (2021) and Zhou et al.
+(2023), estimating which applicants should receive which program to maximize aggregate outcomes.
+A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves treatment effects 86% larger than randomly assigning from the pool of all applicants.
+This gain comes from two margins: selecting individuals with the highest treatment effects and matching individuals to their optimal program type.
+The matching component generates value even at scale: when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment.
+This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion.
+Treatment effect heterogeneity enables improved matching even when all applicants can be served.
+In Mentoring, experienced female technology professionals select mentees from a pre-screened pool based on application materials: cover letters, resumes, pre-recorded videos, and registration surveys.
+Mentors vary systematically in whom they select and what outcomes they achieve.
+Mentors with managerial experience or above-median career tenure select applicants with weaker baseline employment prospects and have stronger treatment effects: a pattern consistent with identifying "high-potential, high-barrier" candidates rather than those already likely to succeed.
+Less experienced mentors favor applicants with stronger baseline employability, selecting candidates who would likely find technology jobs regardless of program participation.
+If effective selection heuristics can be learned through experience, data-driven tools may offer a path to improving targeting among less experienced selectors—a possibility we examine next.
+We find that they can.
+Algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain expertise) achieves treatment effects 63% larger than random assignment and 11% than average mentor selections.
+The two approaches achieve substantial concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates.
 Despite mentors’ rich soft information, systematic patterns in treatment response remain that simple demographic targeting can exploit.
-The result extends recent work showing algorithms can complement expert decisions in domains from bail (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
+This result extends recent work showing algorithms can complement expert decisions in domains from bail (Kleinberg et al., 2018) to medical diagnosis (Mullainathan and Obermeyer, 2022), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
+Our focus on small, custom-designed programs limits external validity but enables deeper investigation of mechanisms.
+Mentoring programs are almost always small due to capacity constraints.
+[^5] Our 150 participants is comparable to Dennehy and Dasgupta (2017) and exceeds the 30–40 per cohort in Blau et al.
+(2010).
+Working with each mentor, we implemented a paired randomization design that allows us to measure both selection and treatment effects—observing not only program impacts but how mentor choices vary with mentor experience.
+We developed Challenges during this research and evaluate its first cohort in this paper; the program has since scaled to multiple technology domains (software testing, cloud computing, UI design) and languages of instruction, suggesting the model extends beyond our pilot context.
 The remainder of the paper proceeds as follows.
@@
 This paper contributes to the literature on treatment effect heterogeneity and optimal policy assignment.
-Building on recent advances in policy learning using machine learning methods (athey2017econometrics; kitagawa2018should; athey2021policy), we examine whether data-driven targeting can improve program allocation decisions in active labor market policy.
-Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (kitagawa2018should; bhattacharya2012inferring).
-Empirical applications of these methods have demonstrated value across diverse domains: customer retention (ascarza2018retention), marketing (hitsch2018heterogeneous), development programs (haushofer2022targeting), agriculture (kasy2021adaptive), and medical treatment (inoue2023machine).
+Building on recent advances in policy learning using machine learning methods (Athey and Imbens, 2017; Kitagawa and Tetenov, 2018; Athey and Wager, 2021), we show that data-driven targeting can substantially improve program allocation decisions in active labor market policy.
+Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (Kitagawa and Tetenov, 2018; Bhattacharya and Dupas, 2012).
+Empirical applications of these methods have demonstrated value across diverse domains: customer retention (Ascarza, 2018), marketing (Hitsch et al., 2024), development programs (Haushofer et al., 2025), agriculture (Kasy and Sautmann, 2021), and medical treatment (Inoue et al., 2023).
 We extend this literature to active labor market policy.
@@
 We further contribute to the growing literature examining whether algorithmic decision-making can complement or improve upon expert human judgment.
-Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), often by identifying systematic patterns that experts miss or cannot consistently apply.
-Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (pallais2014inefficient), we show that observable characteristics predict treatment effects systematically.
+Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (Kleinberg et al., 2018) to medical diagnosis (Mullainathan and Obermeyer, 2022), often by identifying systematic patterns that experts miss or cannot consistently apply.
+Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (Pallais, 2014), we show that observable characteristics predict treatment effects systematically.
 We demonstrate that data-driven targeting can improve upon expert human judgment.
 Mentor-based selection achieves treatment effects 48% larger than random allocation.
-Yet algorithmic targeting using only six observable characteristics achieves an additional 11% improvement.
+Yet algorithmic targeting achieves an additional 11% improvement.
 The 70% concordance between methods indicates complementarity: algorithms largely validate expert choices while identifying additional high-benefit candidates, demonstrating that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity.
 This paper contributes to the literature on mentoring effectiveness in improving career outcomes.
-Several influential papers focus on mentoring in academic contexts: dennehy2017female; resnjanskij2021can; gardiner2007show; ginther2020can; blau2010can show that mentorship increases sense of belonging among female engineering students, youth’ learning outcomes, and tenure rates, publications, and federal grants for women in research careers.
+Several influential papers focus on mentoring in academic contexts and show that mentorship increases sense of belonging among female engineering students (Dennehy and Dasgupta, 2017), youth’ learning outcomes (Resnjanskij et al., 2024), and tenure rates (Ginther et al., 2020), publications, and federal grants for women in research careers (Gardiner et al., 2007; Blau et al., 2010).
 [^6] The impact of mentoring on career outcomes has also been studied using observational data.
-ragins1999mentor survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair.
-kammeyer2008quantitative conduct a meta-analysis concluding that mentoring improves job and career satisfaction.
-Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (granovetter1973strength; ioannides2004job; pallais2016referential).
+Ragins and Cotton (1999) survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair.
+Kammeyer-Mueller and Judge (2008) conduct a meta-analysis concluding that mentoring improves job and career satisfaction.
+Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (Granovetter, 1973; Ioannides and Loury, 2004; Pallais and Sands, 2016).
+Athey et al.
+(2000) develop a formal theoretical model of the dynamics of mentoring and diversity.
 We provide randomized trial evidence on mentoring effectiveness in labor market transitions outside academic settings, demonstrating large effects on technology employment for women career-changers.
 Crucially, we provide experimental evidence that mentoring causally expands professional networks: Mentoring increases participants’ LinkedIn connections by 52.7 (SE 20.8), a 23% increase from baseline, while the non-network-based Challenges program shows no statistically significant effect.
-We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with observable characteristics of both mentors and mentees predicting treatment effects systematically.
+We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with more experienced mentors systematically selecting mentees with higher treatment effects.
+To our knowledge, we are the first to evaluate mentoring effectiveness using a paired randomization design within mentor-selected candidates.
+This approach offers three advantages over standard randomization from an applicant pool.
+First, it allows us to estimate causal treatment effects while preserving the program’s natural selection structure—mentors choose whom they want to work with, avoiding the forced assignments that standard randomization would impose and that mentors might resist.
+Second, by observing mentor choices, we can study how selection varies with mentor characteristics and whether experienced mentors develop effective selection heuristics.
+Third, because we observe outcomes for non-selected applicants alongside selected participants randomized to control, we can directly compare these groups, which is not possible in prior mentoring RCTs, which randomize only within pre-screened applicant pools.
 We contribute to the extensive literature evaluating active labor market training programs.
-card2018works conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion.
-[^6] See also heckman1999economics, card2010active, and crepon2016active for reviews of ALMP effectiveness.
-Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (sianesi2008differential; lechner2000microeconometric), though even intensive apprenticeships show mixed results (biewen2014effectiveness).
-fein2018bridging show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains.
-However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (abebe2021anonymity; bassi2022screening; carranza2020job; athey2024moocs).
-Building on the theoretical work of spence1978job, this literature demonstrates that credentials demonstrating capability can significantly boost earnings (tyler2000estimating; hadavand2018can) and generate market-level benefits when employers face uncertainty about worker quality (pallais2014inefficient).
-Recent evidence further suggests that training effectiveness may depend on complementary interventions: batista2022closing find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
+Card et al.
+(2018) conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion.
+[^7] See also Heckman et al.
+(1999), Card et al.
+(2010), and Crépon and van den Berg (2016) for reviews of ALMP effectiveness.
+Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (Sianesi, 2008; Lechner and Gerfin, 2000), though even intensive apprenticeships show mixed results (Biewen et al., 2014).
+Fein and Hamadyk (2018) show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains.
+However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (Abebe et al., 2021; Bassi and Nansamba, 2022; Carranza et al., 2022; Athey and Palikot, 2024).
+Building on the theoretical work of Spence (1978), this literature shows that credentials demonstrating capability can significantly boost earnings (Tyler et al., 2000; Hadavand et al., 2018) and generate market-level benefits when employers face uncertainty about worker quality (Pallais, 2014).
+Recent evidence further suggests that training effectiveness may depend on complementary interventions: Batista et al.
+(2022) find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
 We provide experimental evidence on training program effectiveness in technology sector transitions, a context involving higher-skilled, career-changing women rather than the unemployed workers typically studied.
-We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in card2018works’ meta-analysis.
+We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in Card et al.
+(2018)’ meta-analysis.
 We directly compare two program modalities: personalized one-on-one mentoring versus scalable online project-based training.
-This head-to-head comparison reveals complementarity rather than substitutability: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment.
+This head-to-head comparison reveals complementarity rather than substitutability at a policy-portfolio level: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment.
 Our subjects are women seeking technology roles who can utilize skill signaling to counter gender-based statistical discrimination.
 Finally, this paper adds to the literature on the gender gap in the technology sector.
-The persistent gender wage gap reflects both occupational segregation and within-occupation differences (juhn2017specialization).
-Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (murciano2022missing) and the impact of occupational stereotypes (cheryan2013stereotypical; del2022more).
-Strategies proposed to counter these barriers include increasing role model visibility (correll2016succeed) and shifting perceptions of technology roles.
+The persistent gender wage gap reflects both occupational segregation and within-occupation differences (Juhn and McCue, 2017).
+Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (Murciano-Goroff, 2022) and the impact of occupational stereotypes (Cheryan et al., 2013; Del Carpio and Guadalupe, 2022).
+Strategies proposed to counter these barriers include increasing role model visibility (Correll and Mackenzie, 2016) and shifting perceptions of technology roles.
 We contribute to this literature by demonstrating how mentoring and credentialing programs can enhance women’s success rates in obtaining technology jobs.
@@
 Labor market transitions face two fundamental frictions: imperfect information about worker abilities and search costs in matching workers to opportunities.
-Workers address these through skill signaling (spence1978job) and professional networks (granovetter1973strength; ioannides2004job; pallais2016referential).
+Workers can address these through skill signaling (Spence, 1978) and professional networks (Granovetter, 1973; Ioannides and Loury, 2004; Pallais and Sands, 2016).
 For women entering male-dominated technology sectors, both mechanisms face gender-specific barriers.
-First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (cheryan2013stereotypical; del2022more; murciano2022missing).
-Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (greguletz2019women; forret2004networking).
+First, occupational stereotypes bias employer perceptions of female candidates’ technical abilities, creating information frictions even for qualified women (Cheryan et al., 2013; Del Carpio and Guadalupe, 2022; Murciano-Goroff, 2022).
+Second, network structures systematically exclude women from informal professional connections that facilitate hiring, with women reporting smaller networks and facing both structural barriers and personal hesitation in networking behaviors that benefit men (Greguletz et al., 2019; Forret and Dougherty, 2004).
 Third, women in male-dominated fields face psychological barriers that may reduce job search effectiveness.
-Women systematically under-report their technical abilities on resumes compared to similarly skilled men (murciano2022missing), use less self-promoting language in professional contexts (lerchenmueller2019gender), and may experience reduced confidence when confronting occupational stereotypes (spencer1999stereotype).
-We study the relative importance of these barriers by evaluating two interventions with distinct mechanisms: a mentoring program that addresses network gaps and builds confidence through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
+Women systematically under-report their technical abilities on resumes compared to similarly skilled men (Murciano-Goroff, 2022), use less self-promoting language in professional contexts (Lerchenmueller et al., 2019), and may experience reduced confidence when confronting occupational stereotypes (Spencer et al., 1999).
+We study the relative importance of these barriers by evaluating two distinct interventions: a mentoring program that primarily focuses on network gaps and confidence building through connections to successful female tech professionals, and a scalable portfolio-building program designed to counter information frictions and stereotypes through credible skill signals.
 Context.
-Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (gus2022).
-Despite strong labor demand driving 9% annual wage growth, women held only 30% of technology positions and comprised just 1.1% of the female labor force, compared to 5.2% of men (EC2021; fluff2021).
-This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (unitednations2021).
+Poland’s technology sector employed 250,000 workers in 2021, with median monthly wages of USD 2,600, nearly triple the national median (GUS, 2022).
+Despite strong labor demand driving 9% annual wage growth, only 1.1% of women in the labor force held positions in the technology sector, compared to 5.2% of men (European Commission, 2021; Gawlowska-Bujok and Bujok, 2021).
+This persistent underrepresentation mirrors patterns across developed economies where women hold approximately 30% of tech jobs and persists despite numerous initiatives supporting women in technology (United Nations, 2021).
+As a high-income economy and a EU member state with gender gaps in technology employment comparable to Western European averages, Poland provides an appropriate setting for studying barriers facing women entering male-dominated sectors.
 3.1 Program Designs and Mechanisms
@@
 Both programs serve women with baseline technical skills who have not yet secured technology jobs.
-The programs’ contrasting designs allow us to contrast network effects from signaling effects.
-Mentoring: Network Expansion and Tacit Knowledge.
+Mentoring:
 Dare IT launched its Mentoring program in 2018 to support women transitioning into technology careers.
@@
 This persistent oversubscription reflects a fundamental scalability constraint: the program’s reliance on volunteer mentors limits expansion, as the pool of qualified female tech professionals willing and able to dedicate four months of weekly mentoring remains finite.
-Mentoring addresses gender-specific network barriers documented in prior research.
-Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (greguletz2019women).
-Female professionals may also hesitate to engage in networking behaviors that appear effective for men (forret2004networking).
-By connecting mentees to established female professionals, the program potentially overcomes both barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
-Our qualitative interviews with mentors reveal four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review.
+Mentoring aims to alleviate gender-specific network barriers documented in prior research.
+Women in technology report smaller professional networks and face structural exclusions from informal networking opportunities as compared to their male colleagues (Greguletz et al., 2019).
+Female professionals may also hesitate to engage in networking behaviors that appear effective for men (Forret and Dougherty, 2004).
+By connecting mentees to established female professionals, the program potentially overcomes these barriers and provides same-gender role models who can navigate gender-specific challenges in recruitment.
+Our qualitative interviews with mentors conducted prior to the experiment revealed four operational channels of the Mentoring program: (i.) direct network expansion through mentor connections and introductions to hiring managers, (ii.) transmission of tacit knowledge about navigating tech sector recruitment as a woman, including addressing potential bias in interviews, (iii.) self-efficacy reinforcement to build confidence, and (iv.) technical skill development through code review.
 Mentors consistently emphasize networking as the primary mechanism, actively facilitating introductions and teaching mentees how to leverage professional connections effectively.
-Challenges: Skill Signaling.
+Challenges:
 To overcome Mentoring’s capacity constraints, we designed Challenges as a scalable alternative targeting information frictions exacerbated by gender stereotypes.
@@
 Managers emphasized that portfolio projects demonstrating hands-on capability could substantially alter hiring decisions, but expressed skepticism about generic online certifications.
-These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (murciano2022missing) and stereotypes about women’s technical competence create additional screening burdens for female candidates (cheryan2013stereotypical).
+These perceptions align with documented gender bias in technical hiring: employers underestimate women’s programming abilities (Murciano-Goroff, 2022) and stereotypes about women’s technical competence create additional screening burdens for female candidates (Cheryan et al., 2013).
 A credible skill signal may be particularly valuable for women if it counters these negative stereotypes by providing concrete, verifiable evidence of technical proficiency.
@@
 If network barriers constrain technology employment for some women, Mentoring should be especially effective for participants with weak existing professional connections: women in smaller cities and towns where local technology networks are sparse, those without family or friends working in technology, and those with longer gaps in employment who may have lost previous professional contacts.
-The network mechanism also predicts effects on observable network measures: in particular, we should observe increases in LinkedIn connection counts among Mentoring participants.
+The network mechanism also predicts expansion of professional networks, which we measure through LinkedIn connection counts.
 Additionally, if experienced mentors’ networks generate value, treatment effects should vary with mentor characteristics.
@@
 We further examine effects on LinkedIn connection counts to directly test the network expansion mechanism.
+Before proceeding, we note an important limitation: the two interventions do not isolate single mechanisms.
+Mentoring may include technical feedback alongside network expansion, while completing Challenges may affect confidence in addition to generating skill signals.
+Our empirical strategy therefore cannot definitively attribute effects to specific channels.
+Instead, we interpret differential treatment effects and heterogeneity patterns as evidence about the relative importance of network-based versus signaling-based approaches, recognizing that each program operates through multiple reinforcing channels.
 4 Experimental Design and Data
 Mentoring.
-We evaluate the Autumn 2021 cohort, for which recruitment opened in October 2021 and the program launched in January 2022.
-As in previous cohorts, applications substantially exceeded the number of available mentors.
-Even after Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking technical skills at a junior professional level, the applicant-to-mentor ratio remained 4:1.
-This oversubscription allowed us to implement a randomized evaluation while preserving Dare IT’s mentor-mentee matching process.
-We modified the standard selection procedure: instead of each mentor selecting one mentee (the historical practice), we asked mentors to select their two preferred candidates from the pool of eligible applicants.
-We then randomized within each pair, assigning one applicant to treatment (participation in Mentoring) and one to control (no participation).
-This paired randomization design controls for mentor selection preferences while generating exogenous variation in program access.
-Of 160 mentors, 151 agreed to this modified selection process; 9 declined, because they did not want to spend effort selecting mentee who will not end up working with them.
-This yielded an initial sample of 302 applicants (151 treatment, 151 control) across 151 mentor pairs.
-However, 4 treatment-assigned participants secured technology jobs between randomization (November 2021) and program launch (January 2022).
-We excluded these individuals from analysis and allowed the initially-assigned control applicant to participate instead, maintaining program access for mentors who invested time in candidate selection.
-Our final experimental sample comprises 151 treatment participants and 147 control participants.
-Appendix G provides robustness checks: (i.) excluding the 4 mentor pairs without matched controls, (ii.) including the 4 pre-program employment cases with outcomes set to 1, and (iii.) including the 9 non-participating mentors under the assumption their hypothetical mentees did not find technology jobs.
-Results are substantively unchanged across specifications.
-In addition to experimental subjects, we collected outcome data on 468 eligible applicants whom mentors did not select.
-These applicants provide a benchmark for assessing mentor selection effects independent of program participation.
-As shown in Figure 2, selected-but-control participants have substantially higher employment success than non-selected applicants, highlighting the importance of randomization for isolating treatment effects from selection effects.
+We evaluate the cohort recruited in October 2021, with the program launching in January 2022.
+As in previous cohorts, applications substantially exceeded mentor availability.
+After Dare IT’s eligibility screening, which excluded applicants currently employed in technology and those lacking junior-level technical skills, the applicant-to-mentor ratio remained 4:1, enabling a randomized evaluation.
+We modified the standard selection procedure to generate exogenous variation in program access.
+Rather than each mentor selecting one mentee (the historical practice), we asked mentors to identify their two preferred candidates from eligible applicants.
+We then randomized within each pair in November 2021, assigning one applicant to treatment and one to control.
+This paired design controls for mentor selection preferences while ensuring balance on characteristics mentors value.
+Of 160 mentors, 151 agreed to this modified process; 9 declined because they did not want to invest effort selecting candidates who might not become their mentees.
+This yielded 302 applicants (151 treatment, 151 control) across 151 mentor pairs.
+Between randomization and program launch, 4 treatment-assigned participants secured technology jobs and withdrew.
+To maintain program access for mentors who had invested time in selection, these slots were filled by the corresponding control-group applicants, yielding 151 treated participants and 147 controls.
+We assess sensitivity to this replacement procedure in Appendix G, estimating effects under alternative sample definitions; results are substantively unchanged.
+We also collected outcome data on 468 eligible applicants whom mentors did not select.
+These non-selected applicants provide a benchmark for assessing selection: control participants whom mentors chose exhibit substantially higher employment rates than non-selected applicants with similar eligibility, confirming that mentor selection predicts success and underscoring the importance of within-pair randomization for identification.
 Challenges.
-We launched recruitment in December 2021, with the program scheduled to begin in February 2022.
+We launched recruitment in December 2021, with the program beginning in February 2022.
 To generate initial interest, we guaranteed admission to the first 100 applicants.
-Registration was capped at 500 applicants, but high demand led to 525 applications before we closed the portal.
-Applicants self-certified that they (i.) were not currently employed in the technology sector, (ii.) possessed English language proficiency at B1 level or higher, and (iii.) had working knowledge of either Figma (User Experience Design track) or React (Front-end Development track).
-Unlike Mentoring, we did not manually verify technical skills prior to admission due to scalability considerations.
-We exclude the first 100 guaranteed-admission applicants from our analysis, as they were not subject to randomization.
-Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access).
-Randomization employed stratification on four registration characteristics: track specialization (UX Design vs.
-Front-end Development), age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
-Between randomization (December 2021) and program launch (February 2022), 17 treatment-assigned applicants secured technology employment.
-Consistent with our approach in Mentoring, we exclude these individuals from our primary analysis.
-Our final analytic sample includes 183 treatment participants and 225 control participants.
-Appendix G demonstrates that results are robust to including these 17 early-employment cases with outcomes coded as 1.
-Among the 183 treatment participants, 41 (22%) did not submit any program assignments, suggesting passive non-compliance.
-However, these individuals retained access to all program materials, weekly webinars with industry experts, and the Dare IT Slack community.
-Following intention-to-treat principles, we include them in our analysis regardless of active participation level.
-We also collected outcome data on 160 applicants who submitted forms after enrollment closed.
-These late applicants were not randomized and do not contribute to our primary treatment effect estimates.
-However, they provide an additional comparison group for the off-policy evaluation in Section 6.
-Appendix B reports covariate balance tests across treatment and control groups.
-Of 9 baseline characteristics, only social science degree shows a statistically significant imbalance at the 5% level, consistent with chance given multiple comparisons.
-Our main specifications control for all observed variables to improve precision.
+Registration was capped at 500, but high demand produced 525 applications before we closed enrollment.
+Applicants self-certified that they (i) were not currently employed in the technology sector, (ii) possessed English proficiency at B1 level or higher, and (iii) had advanced proficiency in either Figma (UX Design track) or React (Front-end Development track).
+Unlike Mentoring, we did not manually verify technical skills prior to admission, reflecting the program’s design for scalability.
+We exclude the 100 guaranteed-admission applicants from our experimental analysis.
+Among the remaining 425 applicants, we randomly assigned 200 to treatment (program access) and 225 to control (no access), stratifying on track specialization, age category, residence type (Warsaw, other large city, or small town), and employment status (student, employed outside tech, or neither).
+Between randomization and program launch, 17 treatment-assigned applicants secured technology employment and chose not to participate.
+Unlike Mentoring, where the paired design required replacing such individuals, here we simply exclude them from the primary analysis.
+Our final analytic sample includes 183 treatment and 225 control participants.
+Appendix G demonstrates robustness to including these 17 cases with outcomes coded as 1.
+Among treatment participants, 41 (22%) submitted no assignments, though they retained access to program materials, weekly webinars with industry experts, and the Dare IT Slack community.
+Following intention-to-treat principles, we include all treatment-assigned participants regardless of engagement level; our estimates therefore capture the effect of program access rather than active participation.
+Under standard LATE assumptions, effects for active participants would be proportionally larger.
+We also collected outcome data on 160 applicants who registered after enrollment closed.
+These late applicants contribute to the off-policy evaluation in Section 6, which estimates treatment effects for populations beyond the experimental sample.
+Appendix B reports covariate balance; of 9 baseline characteristics, only social science degree shows imbalance at the 5% level, consistent with chance.
 4.1 Data
@@
 LinkedIn Data.
-Our primary outcome measure, pre-registered prior to data collection, is whether participants secured technology-sector employment following program completion.
-We collected employment data from participants’ LinkedIn profiles at two time points: four months post-program (May 2022 for Mentoring, June 2022 for Challenges) and twelve months post-program (May 2023 and June 2023, respectively).
+Our primary outcome measure is whether participants secured technology-sector employment following program completion.
+We collected employment data from participants’ LinkedIn profiles until twelve months post-program (May 2023 and June 2023, respectively).
 LinkedIn provides standardized job title metadata for each employment entry.
@@
   1. 1.
-    Any position in a technology company—firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms—excluding finance, regulatory, legal, accounting, and HR roles.
+    Any position in a technology company: firms in software development, testing, and sales; data analytics; IT services; digital marketing; or online platforms; excluding finance, regulatory, legal, accounting, and HR roles.
   2. 2.
@@
 Self-employment, freelancing, and volunteer positions were not coded as new employment.
-Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference (p=0.9), supporting their equivalence.
-Data collection and attrition.
-We obtained complete LinkedIn data for all Mentoring participants at both measurement points.
-For Challenges, we collected data on all participants at four months and 384 of 408 participants (94%) at twelve months.
+Appendix D validates this classification by comparing average salaries for categories (1) and (2) using Glassdoor data; we find no statistically significant difference between (1) and (2) (p=0.9).
+We obtained complete LinkedIn data for all Mentoring participants.
+For Challenges, we collected data on 384 of 408 subjects (94%).
 Missing data resulted from participants changing LinkedIn profile URLs or privacy settings between recruitment and follow-up waves.
-Appendix G provides bounds and estimates under extreme attrition assumptions; main results remain statistically significant under all scenarios.
+Appendix G provides bounds and estimates under extreme attrition assumptions; main estimates remain statistically significant under all scenarios.
 Outcome variables.
@@
 Managerial experience 151 0.52 0.50 0 1 145 0.51 0.50 0 1
-Career changer 151 5.86 3.79 2 35 145 6.02 3.79 2 35
-Years of tech experience 151 0.70 0.46 0 1 145 0.71 0.46 0 1
+Years of tech experience 151 5.86 3.79 2 35 145 6.02 3.79 2 35
+Career changer 151 0.70 0.46 0 1 145 0.71 0.46 0 1
 Mentoring: experimental groups
@@
 Tech job 166 0.38 0.49 0 1 218 0.27 0.44 0 1
-LinkedIn connections 151 277.45 173.47 0 500 147 223.20 185.60 0 500
+LinkedIn connections 166 221.00 180.74 0 500 218 232.40 186.32 0 500
 STEM 183 0.42 0.50 0 1 225 0.39 0.49 0 1
@@
 Small town 183 0.61 0.49 0 1 225 0.63 0.48 0 1
-Applicants both programs
+Applicants both programs (Non-experimental comparison group)
 Tech job 12 months - - - - - 628 0.28 0.45 0 1
@@
 Analytical Framework
-Let each Dare IT applicant be characterized by $(X_{i},Z_{i},W_{i}^{M},W_{i}^{C},S_{i},Y_{i})$, where $i$ is the applicant’s index $i=1,..,\mathcal{I}$, $X_{i}\in\mathcal{X}$ is a matrix of characteristics of applicants that are observed by mentors and recorded in data, $\zeta_{i}\in\mathcal{Z}$ are characteristic of applicants that are not observed in the data (though some might be observed by mentors prior to treatment assignment), and $S_{i}\in\mathcal{S}$ is a random variable indicating whether the applicant has been selected by a mentor.
-[^7] To select mentees, mentors received information from the application form; the important part of the application that mentors observed, which we did not, are short introductory videos recorded by the applicants.
-Finally, $W_{i}^{P}\in\left\{0,1\right\}$ is the treatment assignment into program $P$ and $Y_{i}\in\mathcal{Y}$ is the observed outcome.
-Let $Y_{i}(P)$ and $Y_{i}(0)$ be potential outcomes when applicant $i$ participates in program $P$ and when she does not; $P\in\left\{M,C\right\}$, where $M$ stands for Mentoring and $C$ for Challenges.
-We are primarily interested in estimating the magnitude of the average treatment effect, defined as $\tau^{P}\coloneqq E\left[Y_{i}(P)-Y_{i}(0)\right]$, and testing whether average treatment effects are higher than zero.
-We are also interested in comparing the treatment outcomes to one another, although our estimates of such contrasts are noisy due to the limited sample size in each experiment.
-Selected Sample in Mentoring.
-Dare IT asked mentors to choose two applicants they wanted to work with.
-Thus, selected applicants are not a random draw from the pool.
-Let $S_{i}$ be a random variable indicating whether a mentor selected the applicant $i$ or not; $S_{i}=1$ when the applicant has been selected by a mentor while $S_{i}=0$ otherwise.
-Only selected applicants were included in the randomization.
-Thus, we do not observe outcomes under treatment for the non-selected applicants.
-The average treatment effect of Mentoring that we estimate is defined as
-$\tau^{M}\coloneqq E\left[Y_{i}(M)-Y_{i}(0)|S_{i}=1\right].$ (1)
-Table 1 provides summary statistics of subjects in the experimental groups of the mentoring program and the non-selected applicants.
-There are notable differences between the mentoring control group (those selected by mentors ($S_{i}=1$) but not assigned to treatment ($W_{i}^{M}=0$) and the applicant group (where $S_{i}=0$ and thus, by construction, $W_{i}^{M}=0$), most notably in terms of the primary outcome of securing a tech job.
-In the Challenges program, participants were selected randomly from eligible applicants, where eligibility was determined based on observable characteristics.
-5.1 Average Treatment Effects
-In Figure 2, we display the proportions of participants employed in technology jobs over the span of thirteen months post-application to the program.
-This is illustrated across all experimental groups.
-[^8] We construct outcomes over time based on the new job’s start date.
-In some cases, we only observe the year the new job started rather than the specific month.
-For those observations, we assign December of that year as the starting month.
-across the experimental groups.
-In the left figure, we consider three main groups: participants of Mentoring, participants of Challenges, and all other applicants to Dare IT (we refer to this group as Out of Dare IT).
-We can see that all groups start at zero, with none of the applicants having a job in technology; over time, the share of subjects that have a tech job in each group increases, with the Mentoring group increasing right from the moment the program started, and achieving over 50% probability of having a job in tech after 8 months.
-In contrast, the Challenges group follows a similar trajectory to the Out of Dare IT group until the 8th month, which is approximately 3 months after the end of the program, but then the gap between those two groups increases to almost 20 percentage points.
-This differential timing, with Mentoring effects emerging during the program while Challenges effects manifest several months post-completion, suggests distinct mechanisms.
-These trajectories are consistent with the distinction between immediate network-based job matching versus the creation of portfolio signals and gradual employer response to them.
-The right panel decomposes selection from treatment effects in Mentoring.
-Applicants selected by mentors but randomized to control achieve substantially higher technology employment (43% at month thirteen) than non-selected applicants (19%), confirming that mentors identify promising candidates.
-Furthermore, the treatment-control gap among selected applicants is smaller than the selection effect itself, underscoring the importance of randomization for isolating program impacts from mentor selection.
+Let each applicant $i=1,\ldots,N$ be characterized by observable characteristics $X_{i}\in\mathcal{X}$, unobservable characteristics $Z_{i}\in\mathcal{Z}$, treatment assignment $W_{i}^{P}\in\{0,1\}$ for program $P\in\{M,C\}$, and observed outcome $Y_{i}\in\mathcal{Y}$.
+Let $Y_{i}(P)$ denote the potential outcome under participation in program $P$, where $M$ denotes Mentoring and $C$ denotes Challenges, and let $Y_{i}(0)$ denote the potential outcome absent any program participation.
+For a subpopulation $\mathcal{G}\subseteq\{1,\ldots,N\}$, we define the average treatment effect as
+$\tau^{P}_{\mathcal{G}}\coloneqq\mathbb{E}\left[Y_{i}(P)-Y_{i}(0)\mid i\in\mathcal{G}\right].$ (1)
+Selection into Treatment.
+The two programs employ different selection mechanisms, which affects the interpretation of our estimates.
+In the Challenges program, participants were randomly selected from all eligible applicants, where eligibility was determined by observable characteristics.
+Let $\mathcal{G}^{C}$ denote this eligible population.
+Random selection implies that our estimate of $\tau^{C}_{\mathcal{G}^{C}}$ identifies the average treatment effect for the eligible population.
+In the Mentoring program, applicants were first screened by mentors who selected preferred mentees based on application materials, including short introductory videos, some of which we do not observe in the data.
+Let $\mathcal{G}^{S}\subseteq\{1,\ldots,N\}$ denote the subpopulation of applicants selected by mentors; only selected applicants entered the randomization, and thus we do not observe treated outcomes for non-selected applicants.
+This selection process implies that our estimand for the mentoring program is:
+$\tau^{M}_{\mathcal{G}^{S}}\coloneqq\mathbb{E}\left[Y_{i}(M)-Y_{i}(0)\mid i\in\mathcal{G}^{S}\right].$ (2)
+This quantity represents the average treatment effect for applicants whom mentors chose to work with, a population that may differ systematically from the broader applicant pool on both observable and unobservable dimensions.
+To the extent that mentors select applicants with higher potential returns to mentoring, $\tau^{M}(\mathcal{G}^{S})$ may exceed the average treatment effect that would obtain under random assignment from the full applicant pool.
+5.1 Treatment Effects
+Figure 2 displays technology employment rates over fourteen months following program application, separately for each program.
+[^8] These are simple unweighted averages.
+We drop from these plots subjects who reported only a year in which they started a new job.
+Appendix C presents the plots with these subjects included; the only group with noticeable difference are non-selected applicants to Mentoring.
+In the analysis, we imput month 14 for cases when the month is missing.
+The left panel presents results for the Mentoring program, comparing three groups: treated applicants (selected by mentors and assigned to receive mentoring), control applicants (selected by mentors but assigned to the control group), and non-selected applicants (not chosen by any mentor and therefore excluded from randomization).
+All groups begin at zero as no applicants held technology jobs at baseline.
+Employment rates rise over time across all groups, but at markedly different rates.
+The treatment group exhibits rapid gains beginning immediately after program start, reaching 52% employment by month eight.
+The control group follows a similar but attenuated trajectory, stabilizing at 43%.
+Non-selected applicants show substantially lower employment throughout, reaching only 19% by the end of the observation period.
+Two comparisons merit emphasis.
+First, the gap between selected-control and non-selected applicants (43% versus 19% at month thirteen) confirms that mentors identify promising candidates, whether through observable signals or information revealed in application videos.
+Second, the treatment effect among selected applicants (52% minus 43% = 9 percentage points) is considerably smaller than this selection effect (24 percentage points), underscoring the importance of randomization for isolating program impacts from mentor selection.
+The right panel presents results for the Challenges program, comparing three groups: treated applicants (randomly assigned to participate in the Challenges program), control applicants (randomly assigned to the control group), and late applicants (who applied after randomization closed and therefore received no program exposure).
+The control group and late applicants exhibit nearly identical employment trajectories, with both reaching approximately 20% by month thirteen.
+This rate mirrors that of non-selected applicants in the Mentoring program, suggesting a common counterfactual employment rate for women seeking to enter technology absent intervention.
+The treatment group, by contrast, achieves 39% employment by month fourteen, nearly double the control rate.
+The timing of employment gains differs notably across programs.
+Mentoring participants begin finding technology jobs during the program itself, consistent with mentors actively facilitating job placement through referrals or direct assistance.
+Challenges participants, by contrast, show employment gains concentrated after program completion, consistent with the program’s focus on building portable credentials that participants subsequently deploy in job search.
 Figure 2: Share of Applicants with a Tech Job per Group Over Time.
@@
 The month zero is the month of application to Dare IT.
-In the left figure, we group subjects according to the experimental groups: the treatment group in Mentoring (in green), the treatment group in Challenges (in red), and Out of Dare IT (in blue), where the last group consists of the applicants in control groups, non-selected applicants to Mentoring, and late applicants to Challenges.
-In the right figure, we group applicants into those selected and those not selected by mentors.
-"Selected and treated" group (in green) shows the outcomes of the treatment group in Mentoring, "Selected but not treated" group (in blue) is the control group of Mentoring "Not selected" group (in red) includes applicants to Mentoring that were not selected by mentors, both experimental groups of Challenges, and late applicants to Challenges.
-Table 2 presents estimated average treatment effects.
-We estimate the averate treatment effects using the difference-in-means estimator (columns 1-3, 5-7) and using Cox proportional hazard models that exploit monthly outcome variation (columns 4, 8).
-Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns in Figure 2.
-Both programs generate substantial, statistically significant increases in technology employment.
-Mentoring increases the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group baseline of 45.6%.
-Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the 26.6% control baseline.
-Both estimates use difference-in-means; results are nearly identical using Cox proportional hazard models that exploit monthly variation in job start dates (14.5pp for Mentoring, 11.2pp for Challenges).
-Table 2: Average Treatment Effects of Mentoring and Challenges.
-Difference-in-means Prop. hazard Difference-in-means Prop. hazard
+In the left figure, we present outcomes of three Mentoring groups: the treatment group (in blue), the control group (in green), and the applicants who were not selected by mentors (in red).
+In the right figure, we group focus on the Challenges program groups.
+The average outcome in the treated group (in blue), the control group (in green) and the late applicants to the program who were not included in the randomization and the main analysis (in red).
+Table 2 presents estimated average treatment effects using two approaches: difference-in-means estimation (columns 1–3 and 5–7) and Cox proportional hazard models (columns 4 and 8).
+We report effects on three outcomes: any new job, new technology job, and new non-technology job.
+Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns visible in Figure 2; Appendix I estimates treatment effects using causal survival forests, the method employed in the counterfactual policy analysis of Section 6; Appendix H reweighs the estimates for Challenges by stratum used for randomization.
+Both programs generate substantial and statistically significant increases in technology employment.
+Mentoring raises the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group mean of 45.6%.
+Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the control mean of 26.6%.
+Cox proportional hazard estimates, which exploit the full monthly panel structure, yield nearly identical results: 14.5 percentage points for Mentoring and 11.2 percentage points for Challenges.
+Table 2: Average Treatment Effects on Employment Outcomes
 Mentoring Challenges
-New job Tech job Non-tech job Tech job Cox New job Tech job Non-tech job Tech job Cox
+Difference-in-means Cox PH Difference-in-means Cox PH
+New job Tech job Non-tech job Tech job New job Tech job Non-tech job Tech job
 ATE 0.078 0.153*** -0.076** 0.145*** 0.072 0.113** -0.041 0.112**
-SE (0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
-ATE % baseline 13.576 33.676** -65.641** 31.718*** 14.458 42.646** -17.600 42.232**
-SE (9.893) (14.783) (26.720) (11.992) (10.268) (18.131) (17.978) (19.301)
-No. obs. treated 151 151 151 2265 166 166 166 2656
-No. obs. control 147 147 147 2205 218 218 218 3488
-Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology, and a new job not in technology.
-Columns 1-3 and 5-7 present estimates using the difference-in-means estimator and columns 4 and 8 using a proportional hazard model with a full set of covariates.
-The first row shows estimates of the average treatment effect in the difference in the probability of reporting employment at month 15.
-The third row presents the estimates from the first row divided by the mean outcome in the control group.
-Standard errors are presented below the average treatment effect estimates.
-The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator.
+(0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
+ATE / Control mean (%) 13.6 33.7** -65.6** 31.7*** 14.5 42.6** -17.6 42.2**
+(9.9) (14.8) (26.7) (12.0) (10.3) (18.1) (18.0) (19.3)
+Treated 151 151 151 2,265 166 166 166 2,656
+Control 147 147 147 2,205 218 218 218 3,488
+Notes: Average treatment effects on employment outcomes measured at month 12.
+Columns 1–3 and 5–7 report difference-in-means estimates; columns 4 and 8 report Cox proportional hazard estimates with the full covariate set.
+Cox models use person-month observations.
+“ATE / Control mean” expresses the treatment effect as a percentage of the control group mean.
+Standard errors in parentheses.
 *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
 The differential patterns support distinct mechanisms: Mentoring increases both participants’ capability to access technology jobs (through networks and tacit knowledge) and their willingness to persist in pursuing them (for example, through confidence-building), while Challenges primarily improves application success conditional on applying.
-This finding provides evidence consistent with the self-efficacy channel hypothesized in Section 3: mentors help women maintain focus on technology careers.
 In Appendix E, we show the results based on the outcomes survey.
-We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group.
+Response rates were modest, which is why we rely on LinkedIn data for our primary analysis.
+Among respondents, both treatment groups report a higher number of job offers and higher salaries than the control group.
 The differences between treatment and control are greater in the case of Mentoring than Challenges.
-5.2 Impact of Program Participation on the Number of LinkedIn Connections
+5.2 Impact of Program Participation on LinkedIn Connections
 We directly test the network mechanism by examining whether program participation increases professional network size, measured by LinkedIn connection counts.
-This analysis distinguishes network expansion from other channels: if Mentoring works solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges.
-Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions or employer contact, should not.
+This analysis distinguishes network expansion from other channels: if Mentoring operates solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges.
+Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions, should not.
 The data support the network mechanism.
-Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group baseline of 225, a 23% increase.
-In contrast, Challenges produces a statistically insignificant reduction of 11.4 connections (SE 17.8) from a baseline of 232.
-The differential treatment effect of 64 connections between programs (SE 28) provides direct evidence that Mentoring expands professional networks while Challenges does not.
-To understand where these effects appear in the connection distribution, Table 3 presents quantile regression estimates.
+Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group mean of 225, a 23% increase.
+Challenges, by contrast, produces a statistically insignificant change of 11.4-11.4 connections (SE 17.8) from a control mean of 232.
+The differential effect between programs is 64 connections (SE 28), providing direct evidence that Mentoring expands professional networks while Challenges does not.
+Table 3 presents quantile regression estimates to examine where these effects appear in the connection distribution.
 Mentoring generates large, statistically significant increases at the 25th percentile (62 connections, SE 27) and median (92 connections, SE 35), but no detectable effect at the 75th percentile.
-This pattern indicates the program particularly benefits participants with initially sparse networks.
-In contrast, Challenges shows no significant effects at any quantile, confirming this program does not operate through networking channels.
-Table 3: Treatment Effects on LinkedIn Professional Networks.
-Quantile Treatment Effect Mentoring Treatment Effect Challenges Difference: Mentoring and Challenges
-0.25 62** 5 57*
+The program thus particularly benefits participants with initially sparse networks.
+Challenges shows no significant effects at any quantile, confirming that this program does not operate through network channels.
+Table 3: Treatment Effects on LinkedIn Professional Networks
+Mentoring Challenges Difference
+25th percentile 62** 5 57*
 (27.2) (13.4) (30.3)
-0.5 92*** 5 87*
+Median 92*** 5 87*
 (35.4) (29.5) (46.1)
-0.75 5 -8 13
+75th percentile 5 -8 13
 (47.0) (50.7) (69.1)
-Note: The table provides the quantile regression results for the Challenges and Mentoring programs.
-The treatment effects of the experimental groups on LinkedIn connections are shown at the 25th, 50th, and 75th quantiles.
-The difference column captures the difference in the treatment effects between the Mentoring and Challenges programs.
-Estimates are obtained by regressing the treatment indicator on the number of connections.
-Standard errors are in parentheses.
+Notes: Quantile regression estimates of treatment effects on LinkedIn connection counts.
+The difference column reports Mentoring minus Challenges treatment effects.
+Standard errors in parentheses.
 *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
 The magnitude of network expansion is economically meaningful.
-A 92-connection increase at the median represents substantial growth from the control baseline of 225, comparable to the network size differences between mid-career and senior professionals in our sample.
-Moreover, the increase occurs despite LinkedIn’s 500-connection display cap, which likely attenuates our estimates: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, suggesting the true effect exceeds our measured 52.7-connection average.
-This network expansion cannot be solely attributed to new employment.
-Both Mentoring and Challenges participants secure technology jobs at high rates, yet only Mentoring participants expand their networks.
-Had job acquisition been the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable increases in both programs.
-The absence of connection growth among Challenges participants, despite their 38% technology employment rate, indicates that Mentoring’s network effects operate independently of (and plausibly prior to) job acquisition.
-This timing is consistent with mentors actively facilitating introductions during the program, as reported in our qualitative interviews, rather than networks expanding passively after job placement.
+A 92-connection increase at the median represents substantial growth, comparable to network size differences between mid-career and senior professionals in our sample.
+Moreover, the measured effects likely understate true network expansion due to LinkedIn’s 500-connection display cap: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, attenuating our estimates toward zero.
+This network expansion is not driven by new employment.
+Both programs increase technology employment at substantial rates, yet only Mentoring participants expand their networks.
+If job acquisition were the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable network increases in both programs.
+The small and statistically insignificant estimate for Challenges, despite a 38% technology employment rate among participants, indicates that Mentoring’s network effects operate independently of job acquisition.
+This pattern is consistent with mentors actively facilitating introductions during the program, as reported in qualitative interviews, rather than networks expanding passively after job placement.
 5.3 Heterogeneous Treatment Effects
-The hypothesized contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program.
-We exploit treatment effect heterogeneity to provide evidence consistent with these mechanisms: if Mentoring operates primarily through network expansion (as suggested by LinkedIn connections analysis), effects should concentrate among participants with weak existing networks; if Challenges works through skill signaling, effects should be largest for those with weaker pre-treatment credentials.
-We estimate heterogeneous treatment effects using Cox proportional hazard models that estimate treatment effects in subgroups defined by the pre-treatment characteristics of interest.
-We control for the full covariate set.
-Table 4 presents heterogeneous treatment effects across three participant characteristics that map to credential strength and network access.
-These patterns provide evidence on which populations each program serves most effectively and tests the distinct mechanisms we hypothesize.
+The contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program.
+If Mentoring operates primarily through network expansion, as suggested by the LinkedIn analysis, effects should concentrate among participants with weak existing networks.
+If Challenges works through skill signaling, effects should be largest for participants with weaker observable credentials.
+We estimate heterogeneous treatment effects using Cox proportional hazard models with the full covariate set, estimating separate models for subgroups defined by pre-treatment characteristics.
+[^9] The characteristics we examine are motivated by our theoretical framework.
+However, one might be concerned that these subgroup analyses reflect specification search.
+Section 6 addresses this concern using data-driven methods: we estimate optimal treatment assignment policies via causal survival forests with sample splitting, allowing the algorithm to identify predictive heterogeneity without researcher input on which variables matter.
+The resulting policies generate substantial gains over random assignment, confirming that policy-relevant heterogeneity exists in the data.
+Table 4 presents results across three participant characteristics that map to credential strength and network access.
 The clearest differential pattern emerges from professional experience.
-Challenges generates 28 percentage points larger treatment effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced participants (SE 11).
+Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern, with 20 percentage points larger effects for experienced participants (SE 11).
 The cross-program differential is 49 percentage points (SE 11).
-This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal their technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records.
-Educational background reveals similar patterns.
-Recall that all participants were pre-screened for technical competency, so heterogeneity by educational background likely reflects differences in credential signals and professional networks, and to a lesser degree underlying ability.
+This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records.
+Experienced workers, by contrast, have credentials but may lack networks in technology; mentoring addresses this constraint.
+Educational background reveals an asymmetric pattern.
 Among Mentoring participants, those without STEM degrees show 19 percentage points larger effects than those with STEM backgrounds (SE 11).
-This pattern is consistent with the network access: non-STEM participants followed educational paths outside technology-adjacent fields, limiting opportunities to build connections with peers entering the sector.
-In contrast, Challenges shows no STEM heterogeneity (difference: 0.3pp, SE 11).
-This null result demonstrates that portfolio signals are equally valuable regardless of educational credentials—employers evaluate demonstrated capability rather than inferring skill from degrees.
+A STEM degree might bundle both network access, through connections formed during technical education, and credential signaling value; we cannot cleanly separate these channels.
+However, the fact that Challenges shows no educational heterogeneity (difference: 0.3pp, SE 12) is informative.
+One explanation is that STEM education, as measured here, is too broad to function as a meaningful signal for specific technology roles: a degree in physics or electrical engineering may not credibly signal competence as a front-end developer.
+If so, the Challenges portfolio provides equally valuable skill demonstration for STEM and non-STEM participants alike.
 Geography provides additional evidence on network mechanisms.
-Mentoring generates 26 percentage point treatment effects for participants from small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, p=0.11).
-This pattern aligns with network barriers binding more tightly outside major urban centers, where local technology ecosystems are sparse and professional connections to the sector are limited.
+Mentoring generates 26 percentage point effects for participants currently residing in small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, $p=0.11$).
+Network barriers bind more tightly outside major urban centers, where local technology ecosystems are sparse.
 Mentor networks based in Warsaw and other major cities provide greater marginal value when participants lack local access to technology professionals.
-Challenges shows no significant geographic heterogeneity (difference: -2pp, SE 1.2), consistent with online portfolio signals functioning equally regardless of location, employers evaluate portfolios based on technical content rather than applicant geography.
-Table 4: Heterogeneity in Treatment Effects: Mentoring vs. Challenges.
+Challenges shows no significant geographic heterogeneity (difference: 2-2pp, SE 12), consistent with online portfolio signals functioning independently of applicant location.
+Table 4: Heterogeneous Treatment Effects: Mentoring vs. Challenges
 Mentoring Challenges
-Baseline CATE Baseline CATE Diff. Mentoring – Challenges
+Control mean CATE Control mean CATE Difference
 STEM degree
-Yes 0.538 (0.057) 0.038 (0.079) 0.283 (0.059) 0.12 (0.099) -0.081 (0.082)
-No 0.362 (0.058) 0.228 (0.07) 0.259 (0.035) 0.116 (0.057) 0.111 (0.068)
-Difference subgroups 0.176 (0.081) -0.189 (0.106) 0.024 (0.068) 0.003 (0.114) -0.193 (0.106)
-Length of professional experience
-More than 5 years 0.412 (0.049) 0.202 (0.066) 0.31 (0.039) -0.003 (0.065) 0.205 (0.063)
-Less than 5 years 0.556 (0.075) 0 (0.087) 0.19 (0.042) 0.281 (0.075) -0.281 (0.086)
-Difference subgroups -0.144 (0.089) 0.203 (0.109) 0.120 (0.057) -0.284 (0.099) 0.486 (0.106)
-Small Town
-Small town 0.429 (0.085) 0.26 (0.101) 0.256 (0.067) 0.149 (0.13) 0.111 (0.108)
-Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166 (0.089) -0.133 (0.093)
-Difference subgroups -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
-Note: Estimates of the average treatment effects conditioned on subgroups in Mentoring and Challenges.
-The outcome variable is Tech job, and estimates are obtained using Cox proportional hazards models per subgroup with a full set of covariates.
-The first and third columns show the average outcome per subgroup within the control groups.
-The second and fourth columns show the difference between the treatment and control groups in each subgroup.
-The last column shows the difference in the treatment effects between Mentoring and Challenges.
-Standard errors are in parentheses.
-Additional Characteristics from Mentoring.
-The Mentoring registration survey was more extensive than the Challenges one, allowing us to observe more characteristics.
-We also obtain the characteristics of mentors from their registration survey.
-In Table 5, we present additional results documenting heterogeneity in treatment effects along these dimensions.
-Table 5: Heterogeneity in Treatment Effects: Mentoring.
-Characteristic variable value is 0 Characteristic variable value is 1
-Baseline CATE Baseline CATE Diff. Subgroups
+Yes 0.538 (0.057) 0.038 (0.079) 0.283 (0.059) 0.120 (0.099) -0.081 (0.082)
+No 0.362 (0.058) 0.228*** (0.070) 0.259 (0.035) 0.116** (0.057) 0.111 (0.068)
+Difference 0.176** (0.081) -0.189* (0.106) 0.024 (0.068) 0.003 (0.114) -0.193* (0.106)
+Professional experience
+More than 5 years 0.412 (0.049) 0.202*** (0.066) 0.310 (0.039) -0.003 (0.065) 0.205*** (0.063)
+Less than 5 years 0.556 (0.075) 0.000 (0.087) 0.190 (0.042) 0.281*** (0.075) -0.281*** (0.086)
+Difference -0.144 (0.089) 0.203* (0.109) 0.120** (0.057) -0.284*** (0.099) 0.486*** (0.106)
+Location
+Small town 0.429 (0.085) 0.260** (0.101) 0.256 (0.067) 0.149 (0.130) 0.111 (0.108)
+Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166* (0.089) -0.133 (0.093)
+Difference -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
+Notes: Heterogeneous treatment effects on technology employment by participant characteristics.
+Estimates from Cox proportional hazard models with the full covariate set, estimated separately by subgroup.
+“Control mean” is the employment rate in the control group for each subgroup.
+“CATE” is the average treatment effect conditional on being in a subgroup.
+“Difference” column reports Mentoring CATE minus Challenges CATE.
+Standard errors in parentheses.
+*** $p<0.01$, ** $p<0.05$, * $p<0.1$.
+Additional Heterogeneity in Mentoring.
+The Mentoring registration survey collected more extensive information than the Challenges survey, and we observe mentor characteristics.
+Table 5 presents heterogeneity along these additional dimensions.
+Mentor characteristics strongly predict treatment effects.
+Mentors with managerial experience generate 26 percentage points larger effects than those without (SE 10).
+Similarly, mentors with over five years of technology sector experience produce 26 percentage points larger effects than less experienced mentors (SE 11).
+These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements directly.
+Interestingly, mentor selection patterns reveal that more experienced mentors choose less promising candidates.
+Control group employment among participants selected by mentors with managerial experience is 38%, compared to 55% for those matched to non-managerial mentors.
+Yet treated participants of experienced mentors achieve substantially higher treatment effects (26pp versus 0pp).
+This pattern suggests that experienced mentors either deliberately select candidates who need more assistance, recognizing their capacity to help difficult cases, or value candidate attributes that differ from baseline employment probability.
+First-time mentors are as effective as repeat mentors (difference: 6pp, SE 11).
+This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks and industry standing rather than mentoring skill developed through practice.
+Table 5: Heterogeneous Treatment Effects in Mentoring: Additional Characteristics
+Yes No
+Control mean CATE Control mean CATE Difference
 Mentee characteristics
-Mothers 0.237 (0.07) 0.348 (0.097) 0.532 (0.048) 0.071 (0.062) 0.277 (0.115)
-Applied before 0.5 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143 (0.06) -0.004 (0.124)
+Mother 0.237 (0.070) 0.348*** (0.097) 0.532 (0.048) 0.071 (0.062) 0.277** (0.115)
+Applied before 0.500 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143** (0.060) -0.004 (0.124)
 Mentor characteristics
-Managerial exp. 0.378 (0.056) 0.259 (0.073) 0.549 (0.059) 0.002 (0.074) 0.257 (0.104)
-Long exp. in tech 0.39 (0.063) 0.277 (0.085) 0.512 (0.054) 0.021 (0.066) 0.257 (0.107)
-1st time mentor 0.444 (0.062) 0.155 (0.08) 0.476 (0.055) 0.1 (0.073) 0.056 (0.108)
-Note: Estimates of the average treatment effects conditioned on subgroups in the Mentoring program.
-The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazards models, which include the treatment indicator and the set of controls.
-We estimate the model per subgroup.
-"Long exp.
-in tech" takes the value of 1 if the mentor has worked for over 5 years in the technology sector (the median value).
-Columns 2 and 3 describe the values for subgroups when the subgroup indicator takes the value 0, while columns 4 and 5 describe the values when the indicator takes the value 1.
-The last column shows the difference in the effectiveness of Mentoring between the two subgroups.
-Columns 1 and 3 are the average outcomes in the control groups, and columns 2 and 4 are the estimates of the average treatment effect.
-Standard errors are in parentheses.
-Mentors with managerial experience generate 26 percentage point larger treatment effects than those without (SE 10).
-Similarly, mentors with over five years of technology sector experience produce 26 percentage point larger effects (SE 11).
-These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements.
-Interestingly, mentor selection patterns reveal that experienced, managerial mentors choose less promising candidates.
-Control group participants matched to managerial mentors have 38% baseline technology employment compared to 55% for those matched to non-managerial mentors.
-Yet treated participants of managerial mentors achieve higher outcomes (64% vs.
-64%), because the treatment effect is substantially larger.
-This suggests that experienced mentors either (i.) deliberately select candidates who need more help, recognizing their own capacity to assist difficult cases, or (ii.) value different candidate attributes than baseline employment probability.
-Notably, first-time mentors are as effective as repeat mentors (difference: 6pp, SE 11).
-This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks, technical knowledge, and industry standing—not from mentoring skill developed through practice.
-A senior engineer with extensive connections generates value on her first mentoring experience; repeated mentoring does not improve outcomes.
+Managerial experience 0.378 (0.056) 0.259*** (0.073) 0.549 (0.059) 0.002 (0.074) 0.257** (0.104)
+Long experience in tech 0.390 (0.063) 0.277*** (0.085) 0.512 (0.054) 0.021 (0.066) 0.257** (0.107)
+First-time mentor 0.444 (0.062) 0.155* (0.080) 0.476 (0.055) 0.100 (0.073) 0.056 (0.108)
+Notes: Heterogeneous treatment effects in the Mentoring program by mentee and mentor characteristics.
+Estimates from Cox proportional hazard models with the full covariate set, estimated separately by subgroup.
+“Long experience in tech” indicates mentor has worked over 5 years in the technology sector (median value).
+Standard errors in parentheses.
+*** $p<0.01$, ** $p<0.05$, * $p<0.1$.
 Mothers and Work-Family Barriers.
-The largest heterogeneity we observe is for mothers.
-Among non-mothers, Mentoring generates 7 percentage point treatment effects (SE 6); among mothers, the effect is 35 percentage points (SE 10), a difference of 28 percentage points (SE 12).
-Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 50% for non-mothers.
-Mentoring completely eliminates this gap: treated mothers achieve 59% employment, surpassing non-mothers.
+The largest heterogeneity we observe is by parental status.
+Among non-mothers, Mentoring generates a 7 percentage point treatment effect (SE 6); among mothers, the effect is 35 percentage points (SE 10)—a difference of 28 percentage points (SE 12).
+Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 53% for non-mothers.
+Mentoring eliminates this gap: 59% of treated mothers secure technology employment.
 This large effect for mothers is consistent with Mentoring addressing multiple barriers simultaneously.
-First, mothers may have weaker networks if career interruptions weakened professional connections.
-Second, work-family conflict concerns may reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful mentors in tech may counter this.
+Mothers may have weaker professional networks if career interruptions reduced opportunities to maintain connections.
+Work-family conflict concerns may also reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful women in technology may counter these concerns.
 This finding contributes to understanding gender gaps in technology beyond representation.
 Even among women who have acquired technical skills and are actively seeking technology employment, motherhood creates additional barriers.
-Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that confidence and network deficits also play important roles, and relatively light-touch mentoring interventions can substantially mitigate these barriers.
-The contrasting heterogeneity across programs aligns with our theoretical framework.
-Network barriers might bind most tightly for geographically isolated women and mothers, the groups showing the largest Mentoring effects.
-Information frictions bind most for early-career women with limited work history, those showing the largest Challenges effects.
-The programs’ differential impacts across populations indicate that successful policies must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly.
-Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in which constraints bind most tightly, with implications for optimal program assignment that we explore in Section 6.
+Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that network deficits and confidence also play important roles, and mentoring interventions can substantially mitigate these barriers.
+The contrasting heterogeneity patterns across programs align with our theoretical framework.
+Expanding professional networks creates the greatest opportunity for geographically isolated women and mothers, the groups showing largest Mentoring effects.
+Building credible skill signals offers the greatest opportunity for early-career women with limited work history, those showing largest Challenges effects.
+These differential impacts indicate that effective policy must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly.
+Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in where the greatest opportunities for improvement lie.
+We explore implications for optimal program assignment in Section 6.
 6 Counterfactual Program Assignment Policies
-The number of spots in each program is limited.
-The availability of mentors restricts the capacity of Mentoring.
-In Challenges, we limited the available spots to 300 to accommodate potential operational complexities and logistical problems.
-Considering the substantial treatment effects of both programs, relaxing these capacity constraints will increase Dare IT’s impact.
-However, capacity constraints are likely to remain a long-term issue for the Mentoring program and may persist, albeit to a lesser extent, for Challenges.
-Consequently, Dare IT will continue to face the question of which applicants to accept for each program.
-Given the heterogeneous treatment effects, admission policies that favor applicants with high predicted benefits could enhance the impact of Dare IT programs.
-This section aims to quantify these potential gains.
-We perform three analyses: (i.) we estimate gains from the introduction of a program admission policy that, given capacity levels in each program, assigns applicants to programs, maximizing the sum of predicted treatment effects of programs’ participants; (ii.) we evaluate the gains from relaxing capacity constraints with the targeting policy in place; (iii.) we compare the overall impact of Mentoring when mentors select the most promising candidates to a case where they select to following alternative strategies.
-An assignment policy is a mapping from the characteristics of the applicants $X\in\mathcal{X}$ and the capacity levels of the programs $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT.
-$\mathcal{Q}$ is the grid of possible capacity levels under consideration.
-Policy estimation and evaluation are performed separately in train and test sets.
-Across all analyses, we obtain train and test sets by randomly splitting data at the applicant level with equal probability ($\mathcal{I}_{train}$ and $\mathcal{I}_{test}$ are resulting sets of applicants in the train and test data sets, respectively).
-The objective of policy estimation is to obtain a mapping from applicant characteristics and programs’ capacity levels to program assignments.
-We use the train set to obtain the estimate of the policy $\pi$, which we denote $\hat{\pi}$.
-Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping $\hat{\pi}$ to the test set, where for each $i$ in the test set and $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of the applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$.
-Second, we construct the AIPW estimate of the “policy value,” that is, the outcome value when treatments are assigned according to $\hat{\pi}(X_{i},Q)\in\mathcal{P}$.
-We do that using cross-fitting at the applicant level in the test sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds; specifically, we use 100 repetitions with 5 folds and consider the median outcome across repetitions in (chernozhukov2017double).
-Throughout this section, we combine the two experimental data sets with data on non-selected applicants to Mentoring and late applicants to Challenges.
-Thus, we have 997 applicants who did not participate in any program within our sample, resulting in a total sample of 1314 applicants and 19,000 applicant-month observations.
-For both policy estimation and evaluation, we use survival forests to estimate outcome models and regression forests for the propensity models.
-The propensity model adjusts for differences between participants and non-participants determined by selection to Mentoring and differences in the populations of applicants to Mentoring and to Challenges.
-Appendix H provides details of the models used in policy estimation.
-Appendix I, presents details of the models used in the policy evaluation, the cross-fitting procedure, and validates the approach by comparing the off-policy estimates with the randomized controlled trial estimates.
-Our procedure recovers the experimental estimates.
+Program capacity is limited.
+Mentor availability restricts Mentoring capacity, while we capped Challenges at 300 participants to manage operational complexity.
+Given the substantial treatment effects documented above, relaxing these constraints would increase Dare IT’s impact.
+However, capacity constraints are likely to persist, particularly for Mentoring.
+Dare IT therefore faces an ongoing allocation problem: which applicants should be admitted to each program?
+The heterogeneous treatment effects documented in Section 5.3 suggest that admission policies favoring applicants with high predicted treatment effects could enhance program impact.
+This section quantifies potential gains from such targeting, characterizes which applicant profiles benefit most from each program, and examines whether the patterns align with the theoretical mechanisms identified above.
+We perform three analyses: (i) we estimate gains from an assignment policy that allocates applicants to programs to maximize the sum of predicted treatment effects, given capacity constraints; (ii) we evaluate gains from relaxing capacity constraints with this targeting policy in place; and (iii) we compare overall Mentoring impact under mentor-based selection to alternative assignment strategies.
+Policy framework.
+An assignment policy maps applicant characteristics $X\in\mathcal{X}$ and program capacities $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\{M,C,O\}$ denotes Mentoring, Challenges, and Out of Dare IT, respectively.
+$\mathcal{Q}$ is the grid of capacity levels under consideration.
+We perform policy estimation and evaluation in separate samples.
+We randomly split the data at the applicant level with equal probability into training and test sets ($\mathcal{I}_{\text{train}}$ and $\mathcal{I}_{\text{test}}$).
+Using the training set, we estimate the policy $\hat{\pi}$.
+To evaluate the policy, we apply $\hat{\pi}$ to the test set, obtaining counterfactual assignments $\hat{\pi}(X_{i},Q)\in\mathcal{P}$ for each applicant $i$ and capacity level $Q$.
+We then construct augmented inverse propensity weighted (AIPW) estimates of the policy value—the expected outcome when treatments are assigned according to $\hat{\pi}$.
+We implement this using cross-fitting at the applicant level in the test sample: we estimate outcome and propensity models and obtain AIPW estimates in held-out folds, using 100 repetitions with 5 folds and reporting the median across repetitions (Chernozhukov et al., 2017).
+Sample and estimation.
+We combine the two experimental samples with non-selected applicants to Mentoring and late applicants to Challenges, yielding 1,314 total applicants (682 in experimental samples, 632 non-participants) and 19,000 applicant-month observations.
+We use survival forests to estimate outcome models and regression forests for propensity models.
+The propensity model adjusts for selection into Mentoring and for differences between the Mentoring and Challenges applicant populations.
+Appendix J details the estimation procedure; Appendix K describes the evaluation models and cross-fitting procedure, and validates the approach by comparing off-policy estimates to experimental estimates.
 Outcome measure.
-In all the preceding analyses, we measured treatment effects as changes in the probability of securing technology employment.
-For policy evaluation, we adopt an alternative metric that captures additional dimensions of program impact: the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period.
-The RMST metric offers two advantages for comparing policies.
-First, it incorporates both whether participants find jobs and how quickly they do so, a policy that helps participants find jobs faster, even if it doesn’t change the ultimate employment probability, generates economic value.
-Second, RMST estimates aggregate more naturally across heterogeneous populations and policy scenarios, facilitating comparisons of the assignment rules.
-We report RMST gains in months: negative values indicate participants find technology employment sooner relative to non-participants.
-An RMST gain of -3 months means treated participants find technology jobs an average of three months earlier.
-We use survival forests to estimate average treatment effects (athey2019generalized; cui2023estimating).
+The preceding analyses measured treatment effects as changes in the probability of technology employment.
+For policy evaluation, we adopt the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period.
+RMST incorporates both whether participants find jobs and how quickly they do so: a policy that accelerates job-finding generates economic value even if it does not change ultimate employment probability.
+We report RMST effects in months, with negative values indicating faster job acquisition.
+An RMST effect of 3-3 months means treated participants find technology jobs three months earlier on average than non-participants.
+We estimate treatment effects using causal survival forests (Athey et al., 2019; Cui et al., 2023).
+Appendix I presents average treatment effect estimates in the experimental samples using this method.
 6.1 Targeted Assignment Policies
-Let $z_{ip}\in\left\{0,1\right\}$ be a variable that indicates whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments that maximize the sum of treatment effects given by Equation 2:
-$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\thickspace\forall p\text{ \& }\sum_{p\in\mathcal{P}}z_{ip}=1\thickspace\forall i,$ (2)
-where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$.
-The first constraint ensures that capacity limits are not violated, and the second is that each applicant is assigned to one allocation.
-A policy that induces allocation $z_{ip}^{*}$ is what we call the targeted policy.
-$\tau_{i}^{p}$’s are estimated using the AIPW estimator trained on the train set.
-In Algorithm 1 in Appendix H, we provide further details of the estimation of the targeted policy.
-The outcome of interest, policy value, is the participants’ average treatment effect given program allocation induced by the policy $\pi$.
-Specifically, we use the AIPW estimator to estimate the average difference in outcomes between being selected to participate based on the assignment policy and not participating in any Dare IT program.
-This emphasis on impact, rather than average outcomes of participants, aligns with Dare IT’s mission to create the greatest benefit for their participants.
+We formulate optimal program assignment as an integer linear program.
+Let $z_{ip}\in\{0,1\}$ indicate whether applicant $i$ is assigned to program $p\in\mathcal{P}$, where $\mathcal{P}=\{M,C,O\}$ denotes Mentoring, Challenges, and no program (Out), respectively.
+The optimal assignment solves:
+$\max_{z_{ip}}\sum_{i=1}^{N}\sum_{p\in\mathcal{P}}z_{ip}\tau_{i}^{p}\quad\text{s.t.}\quad\sum_{i=1}^{N}z_{ip}\leq Q^{p}\;\forall p,\quad\sum_{p\in\mathcal{P}}z_{ip}=1\;\forall i,$ (3)
+where $\tau_{i}^{p}=\mathbb{E}[Y_{i}(p)-Y_{i}(O)]$ is applicant $i$’s treatment effect from program $p$ relative to no program, and $Q^{p}$ is the capacity of program $p$ (with $Q^{O}=\infty$).
+The first constraint ensures capacity limits are respected; the second ensures each applicant receives exactly one assignment.
+We call the policy inducing the solution $z_{ip}^{*}$ the targeted policy.
+Individual treatment effects $\tau_{i}^{p}$ are estimated using AIPW with models trained on the training set.
+Algorithm 1 in Appendix J provides implementation details.
+To evaluate policies, we estimate the policy value: the average treatment effect achieved when applicants are assigned according to policy $\pi$ rather than receiving no program.
+This focus on program impact—the causal effect of participation—rather than raw employment rates among participants aligns with Dare IT’s goal of maximizing benefit to participants.
+Below, we compare targeting based on predicted treatment effects to alternative approaches: selecting applicants with the lowest baseline employment probability or those with the highest baseline probability.
 6.2 Value of Targeted Policies
 Given limited program capacity, which applicants should be assigned to which program to maximize overall impact?
-We estimate the targeted policy at three capacity levels: the 15/15/70 policy approximates existing capacity (assigning 15% of all applicants to Mentoring and 15% to Challenges); the 33/33/34 policy increases capacity to 33% for each program; and the 50/50/0 policy allows each program to admit 50% of applicants, ensuring a program slot for every applicant.
-We compare these policies against random allocation as a benchmark (random policy allocates all applicants at random to either Mentoring or Challenges).
-In Section 6.3, we examine how mentor-based selection compares to the algorithm-driven policy in greater detail.
-The policies vary in both capacity levels and whether participants are systematically targeted.
-To understand which applicant types are assigned to each program under data-driven targeting, Figure 3 presents standardized differences in baseline characteristics between optimally assigned individuals and those not assigned to treatment.
-We implement the 15/15/70 targeted allocation policy and compare covariate distributions of each assigned group to the not-assigned group.
-Circles represent comparisons for Mentoring; triangles represent Challenges.
-Dashed vertical lines mark the $\pm$0.1 threshold for imbalance.
-Figure 3: Characteristics of Optimally Assigned Groups
-Notes: Standardized differences in baseline characteristics between optimally assigned individuals and those not assigned under the 15/15/70 allocation policy.
-Circles show Mentoring vs Not Assigned; triangles show Challenges vs Not Assigned.
-The solid vertical line at zero indicates perfect balance; dashed lines mark the $\pm$0.1 threshold.
-Standardized differences are calculated as the difference in means divided by the pooled standard deviation.
-Positive values indicate the optimally assigned group has higher values on that characteristic.
-The targeted allocation exhibits clear differentiation in targeting patterns across programs.
-Applicants selected for Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old.
-Notably, although the algorithm does not use information about prior mentor selection, previously mentor-selected applicants are disproportionately targeted by the algorithm—suggesting the model captures similar signals to those used by mentors.
-For Challenges, the algorithm shows more balanced selection across characteristics.
-Assigned applicants are moderately more likely to have been previously selected by mentors but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience compared to those not selected.
-Main result.
-Table 6 shows gains from data-driven targeting, with effect sizes varying by program and capacity constraint.
-Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, which corresponds to an 86% improvement over random allocation’s 2.33-2.33 months.
-These gains exhibit a strong capacity-selectivity tradeoff: as capacity expands from 15% to 50%, per-person treatment effects decline (from 4.34-4.34 to 2.86-2.86 months overall), though even universal coverage (50/50/0) achieves effects 23% larger than random allocation.
-Thus, we find that there are gains from targeting across programs.
-The targeting gains differ across programs.
-Mentoring exhibits the largest effects under selective policies (5.55-5.55 months at 15% capacity versus 2.31-2.31 under random), while Challenges shows more modest differentiation (3.13-3.13 versus 2.35-2.35).
-Notably, even at 50% capacity where the algorithm must accept marginal applicants with lower predicted treatment effects, Mentoring maintains treatment effects 54% larger than random allocation, demonstrating that observable characteristics contain substantial predictive power for treatment response beyond what random assignment captures.
-Table 6: Targeted Allocation Policies: Treatment Effects by Program
-Capacity Allocation Mentoring Challenges Overall
-(Ment. / Chal. / None) Effect Effect Effect
+We estimate targeted policies at three capacity levels, where capacity is expressed as the percentage of applicants assigned to Mentoring/Challenges/Out (no program): 15/15/70 approximates existing capacity; 33/33/34 expands both programs moderately; and 50/50/0 ensures every applicant receives a program slot.
+As a benchmark, we compare these targeted policies against random assignment, which allocates applicants to Mentoring or Challenges with equal probability.
+We evaluate the random policy at full capacity (50/50/0) because lower capacity levels yield the same expected treatment effects but noisier estimates due to fewer treated subjects.
+To understand which applicant types the algorithm assigns to each program, Figure 3 presents standardized differences in baseline characteristics between optimally assigned applicants and those assigned to no program under the 15/15/70 policy.
+Circles represent Mentoring; triangles represent Challenges.
+Dashed vertical lines mark the $\pm 0.1$ threshold conventionally used to indicate meaningful imbalance.
+Figure 3: Characteristics of Groups Under Targeted Policy
+Notes: Standardized differences in baseline characteristics between applicants assigned to each program and those assigned to no program under the 15/15/70 targeted policy.
+Circles show Mentoring versus Out; triangles show Challenges versus Out.
+Dashed lines mark the $\pm 0.1$ threshold.
+Standardized differences are calculated as the difference in means divided by the pooled standard deviation; positive values indicate the assigned group has higher values on that characteristic.
+Selected by Mentor is not used in the policy assignment, presented here for comparison.
+The algorithm exhibits clear differentiation in targeting across programs.
+Applicants assigned to Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old.
+These patterns align with the heterogeneous treatment effects documented in Section 5.3, where Mentoring showed larger effects for participants outside Warsaw and those without STEM backgrounds.
+Although the algorithm does not observe prior mentor selection status, applicants whom mentors selected in the experiment are disproportionately assigned to Mentoring, suggesting the algorithm captures similar signals to those mentors used when choosing mentees.
+For Challenges, targeting is more balanced across characteristics.
+Assigned applicants are moderately more likely to have been mentor-selected but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience.
+This pattern is consistent with Challenges benefiting early-career women who lack credentials to signal their skills, as documented in the HTE analysis.
+Main results.
+Table 6 presents treatment effects under data-driven targeting at varying capacity levels, estimated over the pooled population of experimental participants from both programs and all other eligible applicants (non-selected Mentoring applicants and late Challenges applicants).
+Recall that negative values indicate faster job finding (months saved).
+Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, an 86% improvement over random allocation’s 2.33-2.33 months.
+These gains exhibit a capacity-selectivity tradeoff: as capacity expands from 15% to 50% per program, average treatment effects decline from 4.34-4.34 to 2.86-2.86 months as the algorithm must include applicants with lower predicted benefits.
+Even at universal coverage (50/50/0), however, targeted assignment achieves effects 23% larger than random allocation.
+[^10] Appendix L presents an alternative specification using repeated cross-fitting following Fava (2025), which uses all observations for both policy estimation and evaluation rather than separate training and test samples.
+Point estimates are nearly identical (e.g., 4.60-4.60 versus 4.34-4.34 months for the 15/15/70 policy), while standard errors are smaller due to efficiency gains from the larger effective sample size.
+The qualitative conclusions are unchanged.
+Targeting gains differ substantially across programs.
+Mentoring exhibits the largest effects under selective policies: 5.55-5.55 months at 15% capacity versus 2.31-2.31 under random assignment.
+Challenges shows more modest gains from targeting (3.13-3.13 versus 2.35-2.35 months, a 33% improvement).
+Even at 50% capacity, targeted allocation achieves Mentoring treatment effects of -3.55 months compared to -2.31 months under random allocation (54% larger), demonstrating that observable characteristics contain substantial predictive power for treatment response.
+Table 6: Treatment Effects Under Targeted Allocation Policies
+Mentoring Challenges Overall
 Panel A: Benchmark
@@
 50% / 50% / 0% -3.55 (0.04) -2.17 (0.03) -2.86 (0.02)
-Notes: Treatment effects measured as change in months to employment (negative values = faster job finding).
+Notes: Treatment effects measured as change in restricted mean survival time (months); negative values indicate faster job finding.
+Capacity allocations specify the share of applicants assigned to Mentoring/Challenges/Out.
+Random allocation assigns each applicant to either program with equal probability.
 Standard errors in parentheses.
-Each row shows a capacity-constrained allocation policy specifying the share of applicants assigned to Mentoring, Challenges, or neither program.
-Random Allocation assigns each program to a random 50% of applicants.
-Table 7 examines treatment effects within each assignment group under the 15/15/70 targeted policy.
-An assignment group consists of all applicants allocated to a specific program (Mentoring, Challenges, or neither) under this policy.
-For each group, we estimate counterfactual treatment effects under alternative program assignments to assess whether the targeted allocation correctly matches individuals to their highest-benefit program.
-The results reveal effective targeting.
-The Mentoring assignment group receives large benefits from their targeted allocation (5.55-5.55 months, SE = 0.086), substantially exceeding what they would gain from Challenges (3.12-3.12 months, SE = 0.062).
-The difference of 2.43 months (SE = 0.106) is statistically significant.
-The Challenges assignment group also benefits most from their assigned program (3.13-3.13 months, SE = 0.069) compared to Mentoring (1.87-1.87 months, SE = 0.050); a difference of 1.26 months (SE = 0.085).
-This pattern demonstrates benefits from targeting across both programs rather than only selecting to a program.
-The Out of Dare IT group (70% of applicants receiving no program assignment) exhibits substantially smaller treatment effects under either counterfactual assignment: 1.72-1.72 months (SE = 0.024) for Mentoring and 2.02-2.02 months (SE = 0.022) for Challenges.
-While this group would experience gains from program participation, their predicted effects are smaller than those optimally assigned to treatment.
-Table 7: Gains From Assignment Policy per Assignment Group
-Mentoring group Challenges group Out of Dare IT group
-Share 0.15 0.15 0.70
-ATE under assignment to:
-Targeted -5.552 (0.086) -3.132 (0.069)
-Mentoring -5.552 (0.086) -1.872 (0.050) -1.720 (0.024)
-Challenges -3.122 (0.062) -3.132 (0.069) -2.017 (0.022)
-Difference in ATE between assignments:
-Mentoring - Challenges -2.431 (0.106) 1.259 (0.085) 0.297 (0.032)
-Note: ATEs are cross-fitted AIPW estimates from causal survival forests (RMST changes up to the evaluation horizon).
-Negative values indicate earlier job finding (months sooner) relative to out of Dare IT programs.
-Standard errors in parentheses are GRF mean-of-variance plug-in SEs aggregated within arm.
+Table 7 examines whether targeting correctly matches applicants to their highest-benefit program.
+We estimate counterfactual treatment effects for each assignment group under the 15/15/70 policy.
+In other words, we estimate, what would applicants assigned to Mentoring gain if instead assigned to Challenges, and vice versa?
+The results confirm effective assignment.
+Applicants assigned to Mentoring gain 5.55-5.55 months from their assignment but would gain only 3.12-3.12 months from Challenges—a difference of 2.43 months (SE 0.11).
+Applicants assigned to Challenges similarly benefit more from their assigned program (3.13-3.13 months) than from Mentoring (1.87-1.87 months), a difference of 1.26 months (SE 0.09).
+The algorithm thus identifies not only who benefits most from treatment, but which treatment benefits each applicant most.
+Applicants assigned to no program (70% of the pool) exhibit smaller predicted effects under either counterfactual: 1.72-1.72 months for Mentoring and 2.02-2.02 months for Challenges.
+While these applicants would benefit from participation, their predicted gains are smaller than those of optimally assigned applicants.
+Interestingly, this group would benefit relatively more from Challenges than Mentoring, suggesting that among lower-priority applicants, the lower-cost program delivers comparable or greater value.
+Table 7: Counterfactual Treatment Effects by Assignment Group
+Assigned to Mentoring Assigned to Challenges Assigned to Out
+Share of applicants 0.15 0.15 0.70
+Treatment effect if assigned to:
+Mentoring -5.55 (0.09) -1.87 (0.05) -1.72 (0.02)
+Challenges -3.12 (0.06) -3.13 (0.07) -2.02 (0.02)
+Difference:
+Mentoring - Challenges -2.43 (0.11) 1.26 (0.09) 0.30 (0.03)
+Notes: Treatment effects measured as change in restricted mean survival time (months); negative values indicate faster job finding.
+Each column represents applicants assigned to that program under the 15/15/70 targeted policy.
+Rows show estimated treatment effects if that group were counterfactually assigned to each program.
+Standard errors in parentheses.
 Policy value across capacity levels.
-When comparing the average treatment effects under targeted assignments and the random assignment, we find that there are subsets of applicants who benefit substantially from targeting.
-However, when the capacity of the two programs increases, the additional participants who get in have lower treatment effects, as suggested by the comparison of targeted policies in Table 6.
-Thus, there are diminishing returns to targeting additional slots.
-To identify the level of capacity that realizes most of the available gains from targeting, we estimate the outcomes from targeted assignment under a sequence of capacity values and compare the resulting average probability of getting a job in technology with this outcome under random assignments to Mentoring and to Challenges.
-Figure 4: Policy Value Across Capacity Levels: Targeted vs Random Allocation
-Notes: Average treatment effects under targeted and random allocation policies as capacity varies from 5% to 50% per program.
-The solid curve shows targeted allocation that assigns individuals with highest predicted treatment effects to each program; the dashed line shows random assignment of 50% to each program.
-Treatment effects measured as change in months to employment (negative = faster job finding).
-Shaded region indicates 95% confidence interval for targeted policy.
-Policy value evaluated using cross-fitted AIPW estimates on test set.
-At each capacity level, targeted allocation maximizes expected policy value by solving the assignment problem in Equation 2.
-Figure 4 presents the capacity-selectivity tradeoff by plotting average treatment effects as program capacity varies from 5% to 50% per program.
+The preceding results examine discrete capacity levels.
+To characterize the full capacity-selectivity tradeoff, Figure 4 plots average treatment effects as program capacity varies continuously from 5% to 50% per program.
 The solid curve shows targeted allocation; the dashed horizontal line represents random assignment.
-Several patterns emerge.
-First, targeted allocation consistently outperforms random allocation across all capacity levels, with gains largest when capacity is most constrained.
-At 15% capacity per program, the targeted policy achieves treatment effects of 4.3-4.3 months compared to random allocation’s 2.3-2.3 months, a 2-month improvement.
-Second, the targeted policy exhibits sharply diminishing returns as capacity expands: per-person effects decline from approximately 4.5-4.5 months at 5% capacity to 2.9-2.9 months at 50% capacity.
-This reflects the algorithm’s sequential assignment of highest-benefit individuals first, with marginal participants exhibiting progressively smaller treatment effects.
-Third, even at universal coverage (50% per program, accommodating all applicants), targeted allocation maintains a 0.6-month advantage over random assignment—demonstrating that complementarity between individual characteristics and program type generates value beyond pure capacity expansion.
+Figure 4: Policy Value Across Capacity Levels
+Notes: Average treatment effects (RMST, in months) under targeted allocation as capacity varies from 5% to 50% per program.
+Shaded region indicates 95% confidence interval.
+Policy value evaluated using cross-fitted AIPW estimates on the test set.
+At each capacity level, targeted allocation solves the assignment problem in Equation 3.
+Dashed line shows the average treatment effect under random assignment with equal probability across programs.
+Three patterns emerge.
+First, targeted allocation outperforms random assignment at every capacity level, with the largest gains when capacity is most constrained.
+At 5% capacity, targeting achieves effects of approximately 4.5-4.5 months compared to random allocation’s 2.3-2.3 months.
+Second, returns to targeting diminish as capacity expands: per-person effects decline steadily as the algorithm exhausts high-benefit applicants and must include those with smaller predicted gains.
+Third, and most notably, targeted allocation maintains a meaningful advantage even at universal coverage (50% per program).
+The 0.6-month improvement at full capacity reflects gains from matching applicants to their best-suited program, not just from selecting high-benefit individuals.
+This complementarity between applicant characteristics and program type generates value independent of capacity constraints.
 6.3 Counterfactual Mentors’ Selection Policies
-Applicants selected by mentors and randomly assigned to the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2).
-This suggests that mentors identify promising applicants who are likely to secure a tech job regardless of their participation in Mentoring.
-Table 8 compares characteristics of selected and non-selected applicants (panel A), and selected by different groups of mentors (panels B and C).
-Baseline is the share of control group applicants that found a job.
+Now, we examine selection within the Mentoring program: how do mentors choose mentees, and could alternative selection strategies improve program impact?
+Understanding selection matters for two reasons.
+First, econometrically, mentor selection determines the population for which our treatment effects are identified.
+Our estimates of $\tau^{M}(\mathcal{G}^{S})$ apply to the mentor-selected population, which may differ from the broader applicant pool.
+Second, mentoring occurs outside experimental settings, and mentor preferences may not align with social efficiency: if mentors favor applicants who would succeed regardless of intervention, or if less experienced mentors lack the judgment to identify high-potential candidates, there is scope for improved targeting.
+Mentors do demonstrably identify promising applicants: as shown in Figure 2, applicants selected by mentors but randomly assigned to the control group achieved substantially higher employment rates than non-selected applicants.
+But whether they identify applicants with high treatment effects—rather than high baseline employability—is an empirical question.
+Table 8 characterizes these selection patterns by comparing selected versus non-selected applicants (Panel A) and by comparing applicants chosen by different types of mentors (Panels B and C).
 Table 8: Differences in Applicant Characteristics: Selection and Mentor Experience
@@
 below the median years of experience.
-Panel A of Table 8 reveals systematic differences between applicants selected by mentors and those not selected.
-Selected applicants exhibit substantially stronger baseline employment prospects (0.456 vs 0.286), despite having similar years of professional experience (7.1 vs 7.5 years).
-Selected applicants are more likely to be over 30 years old (0.487 vs 0.147) and to hold STEM degrees (0.510 vs 0.182), while being considerably less likely to come from small towns (0.245 vs 0.476).
-They are also less likely to have social science backgrounds (0.134 vs 0.207).
-Panel B compares mentors with versus without managerial experience.
-Mentors with managerial experience select applicants with weaker baseline employment prospects (0.378 vs 0.549) but who are substantially older (0.588 vs 0.378) and have more professional experience (7.6 vs 6.5 years).
-Panel C shows that mentors with above-median total career experience exhibit similar but more pronounced patterns.
-They select applicants with weaker baselines (0.390 vs 0.512) who are older (0.590 vs 0.419), have substantially more professional experience (8.2 vs 6.3 years), and are more likely to hold a STEM degree (0.564 vs 0.480).
-The consistency across Panels B and C suggests a common selection philosophy among more experienced mentors: they look beyond current employability to identify "high-potential, high-barrier".
-Less experienced mentors appear to select more conventionally, favoring applicants already demonstrating stronger baseline employability.
-We use our off-policy tools to examine alternative selection strategies that mentors could adopt and their potential implications for the mentoring program’s effectiveness.
-Specifically, we consider four scenarios: (i) Highest CATE—targeting applicants predicted to have the highest treatment effects, (ii) Most Promising—targeting applicants expected to have the highest outcomes without participating in Mentoring, (iii) Least Promising—selecting applicants expected to have the lowest outcomes if they do not participate, and (iv) Highest Chances if Treated—selecting participants with the highest predicted outcomes if they participate in Mentoring.
-Additionally, we include two benchmarks: mentors selecting at random (Random Choice) and the observed allocation (Observed).
-In each of the four policies, we assume that 15% of applicants participate in Mentoring, which is an approximate share of participants to applicants observed by DareIT in recent years.
-Table 9 examines the extent to which different allocation policies select the same individuals Each cell shows the percentage of individuals in the row policy who also appear in the column policy.
-First, we consider the relationship between mentor-based selection (Current) and the Highest CATE algorithmic policy.
-Among current mentor-selected participants, 70.4% would also be assigned to Mentoring under the Highest CATE algorithm, demonstrating substantial concordance between expert judgment and data-driven targeting (note that we do not use the information on whether a candidate was selected or not in the policy estimation).
-Current mentor selections align most closely with Highest CATE (70.4%), followed by Most Promising (29.6%), with little overlap with Least Promising (18.3%) or Highest Chances (19.7%).
-This pattern suggests mentors implicitly optimize for treatment effects rather than baseline characteristics or final outcomes, though with some cream-skimming tendency.
-The 29.6% overlap with Most Promising indicates approximately one-third of mentor selections favor strong baseline candidates, while the predominant 70% overlap with Highest CATE confirms that mentors primarily target based on perceived program responsiveness.
-The low overlap with Least Promising (18.3%) suggests mentors, whether by design or implicit bias, rarely select individuals with the weakest baseline labor market attachment, even when such individuals might benefit from treatment.
-Table 9: Overlap Between Allocation Policies
-Policy H. CATE Most Prom. Least Prom. H. Chances Current
-Highest CATE 100.0
+Panel A reveals that mentors select applicants with substantially stronger baseline employment prospects: 46% of selected applicants in the control group find technology jobs, compared to 29% of non-selected applicants.
+Selected applicants are more likely to be over 30 (49% vs.
+15%), hold STEM degrees (51% vs.
+18%), and come from urban areas outside small towns (25% vs.
+48% from small towns).
+This selection pattern presents a puzzle: Section 5.3 documented larger treatment effects for applicants from small towns and those without STEM degrees.
+Mentors appear to select on baseline employability rather than on treatment effect magnitude.
+Panels B and C suggest that experienced mentors employ a different selection philosophy.
+Mentors with managerial experience select applicants with weaker baseline employment prospects (38% vs.
+55%) but who are older (59% vs.
+38% over 30) and have more professional experience (7.6 vs.
+6.5 years).
+Panel C shows similar patterns for mentors with above-median career experience.
+These experienced mentors appear to identify applicants with high potential who face significant barriers; perhaps recognizing that such applicants offer greater scope for mentoring impact.
+Less experienced mentors select more conventionally, favoring applicants with stronger baseline employability.
+Alternative selection strategies.
+We use our policy evaluation framework to examine how alternative selection strategies would affect Mentoring effectiveness.
+We consider four algorithmic approaches, each selecting 15% of applicants (approximating Dare IT’s typical capacity):
+  • Highest Treatment Effect: select applicants with the largest predicted treatment effects $\hat{\tau}_{i}^{M}$
+  • Most Promising: select applicants with the highest predicted outcomes absent treatment, $\hat{Y}_{i}(0)$
+  • Least Promising: select applicants with the lowest predicted outcomes absent treatment
+  • Best if Treated: select applicants with the highest predicted outcomes under treatment, $\hat{Y}_{i}(M)$
+We benchmark these against random selection and the current mentor-based allocation.
+Overlap between selection strategies.
+Table 9 examines the extent to which different selection strategies choose the same individuals.
+Each cell shows the percentage of individuals selected under the row policy who also appear in the column policy.
+Importantly, the algorithmic policies do not use information on whether an applicant was actually selected by mentors; any overlap reflects convergence on similar applicant characteristics through independent criteria.
+Table 9: Overlap Between Selection Strategies (%)
+Highest TE Most Prom. Least Prom. Best if Treated Current
+Highest Treatment Effect 100.0
 Most Promising 28.5 100.0
 Least Promising 18.2 0.0 100.0
-Highest Chances 27.9 26.7 31.5 100.0
-Current 70.4 29.6 18.3 19.7 100.0
-Notes: Each cell shows the percentage of individuals in the row policy that also appear in the column policy.
-Only lower triangle shown to avoid redundancy; upper triangle omitted.
-Sample sizes in parentheses after policy names.
-Diagonal entries are 100% by construction (bold).
-H.
-CATE = Highest CATE; H.
-Chances = Highest Chances; Most/Least Prom.
-= Most/Least Promising.
-Table 10 compares treatment effects across six allocation policies for the Mentoring program.
-The results demonstrate that targeting based on predicted treatment effects substantially outperforms alternative strategies.
-The Highest CATE policy achieves treatment effects of 3.59-3.59 months (SE = 0.033), the largest magnitude across all policies.
-Random allocation serves as the baseline benchmark, generating treatment effects of 2.20-2.20 months (SE = 0.040).
-Current mentor-based selection performs remarkably well, achieving 3.25-3.25 months (SE = 0.100), a 48% improvement over random allocation.
-This validates that mentors successfully identify high-benefit participants through their evaluative selection process.
-Table 10: Allocation Policy Comparison: Mentoring Program
-Policy Baseline Time Treatment Effect
+Best if Treated 27.9 26.7 31.5 100.0
+Current (Mentor) 70.4 29.6 18.3 19.7 100.0
+Notes: Each cell shows the percentage of row-policy selections that also appear in the column policy.
+Lower triangle only; upper triangle symmetric.
+All policies select 15% of applicants.
+Under random selection, expected overlap would be 15%.
+The current mentor-based selection exhibits substantial overlap with the Highest Treatment Effect policy: 70% of mentor-selected applicants would also be selected by the treatment-effect-maximizing algorithm.
+This overlap far exceeds the 15% expected under random selection, indicating that mentors, through whatever criteria they apply, tend to choose applicants whom the algorithm also identifies as high-benefit.
+The overlap with Most Promising is considerably lower (30%), and overlap with Least Promising is lower still (18%), suggesting mentors do not primarily select based on baseline employability in either direction.
+These patterns are consistent with two interpretations.
+First, mentors may implicitly target treatment effects, recognizing which applicants would benefit most from mentoring.
+Second, the characteristics that predict high treatment effects (such as facing network barriers) may correlate with other attributes mentors find compelling in application materials.
+We cannot distinguish these explanations, but the high concordance between mentor judgment and algorithmic targeting suggests that scaling the Mentoring program with algorithmic selection would likely preserve much of the effectiveness achieved under mentor-based selection.
+Comparing selection strategies.
+Table 10 compares treatment effects across six selection strategies for the Mentoring program.
+The Highest Treatment Effect policy achieves the largest impact: 3.59-3.59 months (SE 0.03), meaning selected participants find technology jobs 3.6 months faster than they would without the program.
+Random selection serves as a benchmark, generating effects of 2.20-2.20 months (SE 0.04).
+The Current mentor-based selection performs well, achieving 3.25-3.25 months (SE 0.10), a 48% improvement over random allocation.
+This validates that mentors successfully identify high-benefit applicants through their selection process.
+Table 10: Treatment Effects Under Alternative Selection Strategies: Mentoring
+Strategy Baseline Treatment Effect
 (months) (months)
-Current 11.09 (0.131) -3.25 (0.100)
+Current (Mentor) 11.09 (0.13) -3.25 (0.10)
 Random 11.36 (0.06) -2.20 (0.04)
-Data-driven policies:
-Highest CATE 11.08 (0.086) -3.59 (0.033)
+Algorithmic strategies:
+Highest Treatment Effect 11.08 (0.09) -3.59 (0.03)
 Most Promising 9.71 (0.04) -2.27 (0.08)
 Least Promising 13.44 (0.09) -1.94 (0.07)
-Highest Chances 11.77 (0.152) -2.14 (0.079)
-Notes: Comparison of allocation policies for the Mentoring program.
-Baseline Time is the predicted months to employment without treatment (lower is better); Treatment Effect is the predicted change in months from treatment (more negative = faster employment).
+Best if Treated 11.77 (0.15) -2.14 (0.08)
+Notes: Comparison of selection strategies for Mentoring, each selecting 15% of applicants.
+“Baseline” is predicted months to employment without treatment (lower = faster job finding absent intervention).
+“Treatment Effect” is the change in months from treatment (negative = faster job finding due to program).
 Standard errors in parentheses.
-Current reflects mentor-selected participants.
-Random is random assignment benchmark.
-Highest CATE maximizes treatment effects; Most/Least Promising target best/worst baseline outcomes; Highest Chances minimizes post-treatment time to employment.
-Notably, the Highest CATE algorithmic policy achieves an additional 10.4% improvement over mentor selection, increasing treatment effects from 3.25-3.25 to 3.59-3.59 months.
-We obtain the 0.34-month gain using only six readily-observable characteristics (education, location, age, experience) while mentors access rich soft information from interviews and holistic assessment.
-The overlap analysis (Table 9) reveals that 70.4% of mentor-selected participants would also be assigned under Highest CATE, demonstrating substantial concordance between expert judgment and data-driven targeting.
-The finding that simple observable characteristics can meaningfully improve upon expert selection—even when experts demonstrate strong performance—underscores the value of systematic data-driven approaches as complements to human judgment.
-Policies optimizing objectives other than treatment effects yield substantially weaker results.
-Most Promising, which targets applicants with the strongest baseline prospects (9.71 months to employment), achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation despite successfully cream-skimming the highest-potential candidates.
-This demonstrates that strong baseline prospects do not predict strong treatment responsiveness.
-Conversely, Least Promising targets individuals with weakest baselines (13.44 months) but generates the smallest treatment effects (1.94-1.94 months), revealing an unfortunate equity-efficiency tradeoff: those who need help most benefit least from the program.
-Highest Chances, which optimizes final employment outcomes, achieves 2.14-2.14 months, marginally worse than random assignment.
-These comparisons underscore that maximizing program impact requires explicitly targeting predicted treatment effects rather than baseline characteristics or final outcomes, as treatment response heterogeneity does not align with baseline heterogeneity.
+The Highest Treatment Effect policy achieves a 10% improvement over mentor selection (3.59-3.59 versus 3.25-3.25 months).
+Notably, this algorithmic policy uses only observable application characteristics, while mentors access richer information including video introductions and holistic assessment.
+The substantial overlap between mentor selections and algorithmic targeting (70%, from Table 9) suggests that mentors and the algorithm identify similar applicant profiles through different information sets.
+Data-driven selection could complement mentor judgment by providing a systematic check or by scaling selection when mentor capacity is limited.
+Policies targeting objectives other than treatment effects perform substantially worse.
+Most Promising selects applicants with the strongest baseline prospects—those predicted to find jobs in 9.7 months even without intervention—yet achieves treatment effects of only 2.27-2.27 months, barely exceeding random allocation.
+Strong baseline prospects do not predict strong treatment response.
+Conversely, Least Promising targets applicants with the weakest baselines (13.4 months to employment without intervention) but generates the smallest treatment effects (1.94-1.94 months).
+This pattern reveals a tension between equity and efficiency: applicants with the greatest baseline disadvantage benefit less from the program than those with moderate baselines.
+However, even the Least Promising group experiences meaningful gains: nearly two months of accelerated job finding, suggesting that equity-focused selection remains beneficial in absolute terms, if not maximal in efficiency terms.
+Best if Treated, which selects applicants predicted to have the best outcomes under treatment rather than the largest treatment effects, performs similarly to random selection (2.14-2.14 versus 2.20-2.20 months).
+This underscores a key distinction: applicants who perform best after treatment are not the same as applicants who benefit most from treatment.
+Program impact depends on the counterfactual, what would have happened absent intervention, not just realized outcomes.
 7 Conclusion
@@
 However, effectively designing such policies requires principled evaluation of various training programs in terms of their overall impact and their heterogeneity.
-In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector.
-This way we identify a class of interventions that can effectively narrow the gender gap in technology.
-We show that a mentoring program and an online training program called Challenges aimed at developing labor market signals increase the probability of having a job in the technology sector by over 30% twelve months after the end of the program.
-We estimate that Mentoring increases technology employment from 46% to 61%, while Challenges increases it from 27% to 38%, with treatment effects persisting over the full observation period.
-Furthermore, the online training program developed during this research project is cheap to operate, at approximately \$15 per person, and suitable for scaling up, suggesting that it can have an impact extending beyond the current setting.
-Our findings reveal strong patterns of heterogeneity in treatment effects.
-The benefits derived from the two programs depend on a number of participants’ characteristics including the educational background, length of professional experience, type of residence, and age.
-For example, we find that older participants that come from small towns have particularly high treatment effects in both programs, especially in Mentoring.
-Furthermore, certain groups benefit distinctly from one program over the other.
-Specifically, participants with short professional experience benefit from Challenges, while those with over five years of experience from Mentoring.
-These differential patterns provide evidence on the underlying mechanisms: Mentoring expands professional networks (increasing LinkedIn connections by 52.7, a 23% increase), while Challenges provides credible skill signals.
-Notably, network expansion occurs only in Mentoring, and heterogeneity patterns align with each program’s mechanism: those with weaker networks benefit most from Mentoring, while those with weaker credentials benefit most from Challenges.
-For the foreseeable future, the demand for the two programs, particularly for Mentoring will exceed the number of available spots, and Dare IT will grapple with the question of whom to admit.
-We propose an admission policy that maximizes Dare IT’s impact, by prioritizing applicants with the highest predicted treatment effects.
-Consistent with the high degree of heterogeneity in the effectiveness of the two programs we find that such targeted admission policies can substantially increase Dare IT’s effectiveness.
-Specifically, we estimate that optimal targeting achieves outcomes 86% larger than random assignment at current capacity levels (30% of applicants).
-Notably, the value of targeting persists even as capacity expands: at 50% capacity per program, targeted allocation maintains outcomes 23% larger than random assignment.
-This demonstrates that targeting generates value both through selecting who participates and through matching individuals to their highest-benefit program.
-The complementarity between individual characteristics and program type creates allocation gains beyond simple capacity expansion.
-A noteworthy aspect of Mentoring is that mentors have the autonomy to choose their mentees.
-To facilitate randomized access for our experiment, we asked mentors to shortlist two candidates and we randomized within these pairs.
-Comparing selected-but-control applicants to non-selected applicants reveals that mentors successfully identify promising candidates: mentor-selected participants achieve outcomes 48% larger than random allocation.
-However, mentors appear to prioritize candidates with strong baseline prospects, those likely to secure technology employment even without the program, rather than those who would benefit most from participation.
-We evaluate an alternative algorithmic allocation policy that selects participants based on predicted treatment effects using only six readily-observable characteristics (education, location, age, experience).
-Despite mentors’ rich soft information from interviews and holistic assessment, algorithmic targeting achieves outcomes 63% larger than random allocation, representing an additional 11% improvement over mentor judgment.
-This finding is particularly striking given that algorithmic and mentor-based selections exhibit 70% concordance: the algorithm largely validates expert choices while identifying additional high-benefit candidates.
-The results demonstrate that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity, and that simple demographic targeting can complement human judgment to enhance program effectiveness.
-This study comes with certain limitations worth noting.
-First, the limited participant size in both programs led to wide confidence intervals in several estimates, leaving room for uncertainty regarding the exact impacts of Mentoring and Challenges.
-Secondly, we could not pinpoint the exact mechanisms driving the outcomes for both programs.
-Although our findings suggest that Mentoring likely enhances participants’ professional networks and teaches how to utilize them, and Challenges supports participants in the development of a credible signal of their skills, a more in-depth exploration is required to understand the specific mechanisms through which the training provided by both programs impacts employability.
-Such an analysis could also help scale up Mentoring by identifying the crucial aspects of the program and directing Dare IT resources to them.
-Lastly, the participants from the first edition of Challenges may not be representative of those in later editions.
-Early adopters of such programs might experience greater benefits.
-To validate the program’s continued efficacy, subsequent evaluations are essential.
+In this paper, we provide evidence of the high effectiveness of two programs supporting women transitioning into the technology sector, identifying a class of interventions that can effectively narrow the gender gap in technology.
+Both programs generate large, statistically significant effects on technology employment.
+Mentoring increases technology employment by 15 percentage points (from 46% to 61%), while Challenges increases it by 11 percentage points (from 27% to 38%), with effects persisting throughout our observation period.
+The Challenges program operates at approximately \$15 per participant, implying a cost of roughly \$150 per additional technology job—a remarkably low figure relative to traditional training programs.
+The two programs address distinct barriers to labor market entry.
+Mentoring expands professional networks: participants gain 53 LinkedIn connections on average, a 23% increase, while Challenges participants show no network growth.
+Challenges instead provides credible skill signals through portfolio projects that demonstrate technical competence to prospective employers.
+These distinct mechanisms generate predictable heterogeneity in treatment effects.
+Expanding professional networks creates the greatest opportunity for women from small towns, those without STEM degrees, and mothers.
+Building credible skill signals offers the greatest opportunity for early-career women with limited work history.
+The 28 percentage point larger effect for mothers in Mentoring is particularly striking, suggesting that mentoring can substantially reduce motherhood penalties in technology employment.
+These heterogeneous effects have implications for program design.
+We develop an assignment policy that allocates applicants to their highest-benefit program based on predicted treatment effects.
+At current capacity levels, targeted assignment achieves effects 86% larger than random allocation.
+Even at expanded capacity where every applicant receives a program slot, targeting generates 23% larger effects than random assignment by matching individuals to their best-suited program.
+This demonstrates that the value of targeting comes not only from selecting high-benefit individuals but also from exploiting complementarities between participant characteristics and program mechanisms.
+We also examine mentor-based selection within the Mentoring program.
+Mentors successfully identify promising candidates: their selections achieve effects 48% larger than random allocation.
+Algorithmic selection based on observable characteristics achieves a further 10% improvement.
+The 70% overlap between mentor and algorithmic selections suggests both approaches identify similar applicant profiles.
+Data-driven tools could complement mentor judgment, particularly for scaling selection when mentor capacity is limited.
+Several limitations merit acknowledgment.
+First, moderate sample sizes yield wide confidence intervals for some estimates, particularly for subgroup analyses.
+Second, while our evidence is consistent with network and signaling mechanisms, we cannot definitively isolate the causal channels through which each program operates.
+Understanding these mechanisms more precisely could inform efforts to scale Mentoring by identifying its essential components.
+Third, for Challenges, our estimates come from the first program cohort; subsequent cohorts may differ in composition or in treatment response as the programs mature and selection patterns evolve.
+Yet studying small, custom-designed programs enabled the access and stakeholder cooperation necessary for our research design, including the novel paired randomization within mentor-selected candidates that allows us to separately identify selection and treatment effects, a contribution absent from prior mentoring RCTs.
+More broadly, having two randomized interventions targeting distinct mechanisms within the same population enables comparisons rarely possible in program evaluation.
+We observe not only employment outcomes but also professional network expansion, allowing us to directly test whether mentoring operates through networks.
+We examine mentor selection behavior—how mentors choose, whether experienced mentors select differently, and how mentor choices compare to algorithmic efficiency.
+These analyses were possible only because our design preserves mentor discretion while enabling randomization.
+Our findings demonstrate that women’s underrepresentation in technology reflects multiple distinct barriers—network access, credentialing, confidence—that require different interventions.
+Effective policy should provide multiple pathways tailored to different populations rather than assuming a single intervention serves all.
+The substantial heterogeneity we document, and the gains from exploiting it through targeted assignment, suggest that personalized approaches to workforce development may substantially outperform one-size-fits-all programs.
 Appendix
@@
 In Figure 6 we show the balance of covariates in the Challenges experiment: each dot corresponds to the absolute difference of means in treatment and control divided by the square root of the sum of variances.
-The only variable that has a statistically significant differnece is social science.
+The only variable with a standardized absolute mean difference above the 0.1 threshold is is social science.
 Figure 6: Covariate Balance Challenges.
@@
 Points correspond to the absolute difference in means between treatment and control divided by the square root of the sum of variances in treatment and control.
-Appendix C Kaplan Meier Survival Curves
+Appendix C Monthly Outcomes Plots
+Monthly Employment Outcomes.
+Figure 7 presents proportions of subjects across programs and treatment groups who reported a job in the technology sector with the inclusion of subjects who reported a job without a month in which they started.
+When a subject reported a job without a month, but only with the year in which they started we assign the last month in the data.
+Figure 7: Share of Applicants with a Tech Job per Group Over Time.
+Note: Share of subjects with a tech job per group across time.
+The month zero is the month of application to Dare IT.
+In the left figure, we present outcomes of three Mentoring groups: the treatment group (in blue), the control group (in green), and the applicants who were not selected by mentors (in red).
+In the right figure, we group focus on the Challenges program groups.
+The average outcome in the treated group (in blue), the control group (in green) and the late applicants to the program who were not included in the randomization and the main analysis (in red).
+We notice that there is an increase in the shares that are employed in the last month and this is particularly pronounced for non-selected Mentoring applicants.
+Kaplan Meier Survival Curves.
 We present Kaplan Meier estimates of the survival function $S(t)$, defined as the probability of not yet obtaining a technology job by month $t$.
@@
 We use data from outcomes surveys to provide additional evidence of the effectiveness of the two programs that give insights into the mechanisms through which participation boosts outcomes.
-There were three surveys: two main ones carried out approximately four months after the end of each program, and a shorter survey right after the end of the mentoring program from which we obtained information on salaries in jobs found during or right after the mentoring program.
+There were three surveys: one main survey for each program, both administered about four months after completion, and a shorter survey right after the end of the Mentoring from which we obtained information on salaries in jobs found during or right after the mentoring program.
 As discussed in Section 4, results based on the outcomes survey suffer from a selection problem.
@@
 Orange bars indicate the treatment group, and the green ones are the control group.
-The right panel is based on the survey carried out right after the end of the program, while the left panel is from the survey conducted four months after the end of the program.
+The left panel is based on the survey carried out right after the end of the program, while the right panel is from the survey conducted four months after the end of the program.
 Figure 13: Salary ranges: Mentoring.
@@
 The length of a bar represent the share of respondents declaring a salary within the range.
-The right panel survey was conducted right after the end of the program; the left panel survey was conducted four months after the program.
+The left panel survey was conducted right after the end of the program; the right panel survey was conducted four months after the program.
 Evidence presented in Figure 13 suggests that subjects in the treated group of the mentoring experiment that found a new job are more likely to have a salary in a high salary range.
@@
 The first three columns show results for the Mentoring model, and the last three for Challenges.
-City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is also a discrete variable with levels at: no education mentioned, non-traditional credential, college degree, masters, and postgraduate.
+City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is a discrete variable indicating the highest level of education with levels: no education mentioned, non-traditional credential, college degree, masters, and postgraduate.
 Prof experience is the number of years between the start date of the first employment of program participants and 2023.
@@
 First, we consider several cases in which we change which subjects are included in the analysis.
-Second, we estimate AT bounds following lee2021bounding.
+Second, we estimate AT bounds following Lee and Weidner (2021).
 Third, we estimate the average treatment effects at 4 months after the programs.
@@
 We consider four specifications for Mentoring.
-First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where both candidates were assigned to treatment due to control participants securing early employment.
-Second, restricting to the 126 perfectly matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3).
+First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where the mentees who were initially assigned to participate found jobs before the program start and were replaced by the control group.
+Second, restricting to the 147 matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3).
 Third, including the 4 participants who found jobs before program launch—coding their outcomes as 1—increases the estimated effect to 16.4 percentage points.
@@
    Main specification 151 147 0.609 0.456 0.153***
-   Balanced pairs only 126 126 0.603 0.484 0.119*
+   Matched pairs only 147 147 0.603 0.484 0.119*
    Include pre-program employment 155 147 0.619 0.456 0.164***
@@
 Our twelve-month outcome measurement experienced modest attrition in the Challenges program: 6% had missing LinkedIn data due to profile URL changes or privacy setting modifications.
-We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of lee2009training.
+We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of Lee (2009).
 We consider the most conservative case, in which subjects that found jobs before the start of the program are not considered; thus, we focus only on the attrition during the program, due to missing LinkedIn data.
@@
 The upper bound uses all observed treatment outcomes compared to all observed control outcomes.
+Trimmed mean is calculated over the remaining observations.
 We estimate Lee bounds of [0.065, 0.114].
@@
 Any attrition-induced bias is therefore eliminated at this time horizon.
+Furthermore, comparing the estimates at 4 months with the main results shows persistence of the estimates over time.
 Table 13 presents the results.
@@
 Importantly, these four-month estimates use the full experimental sample without any missing observations, confirming that our main findings are not an artifact of selective attrition.
-The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts card2018works.
+The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts Card et al.
+(2018).
 If anything, the point estimates suggest persistent or growing effects, particularly for Challenges, though the conclusions are limited to 12 months from the start of each program.
-Appendix H Policy estimation algorithms
+Appendix H Stratified Estimation for Challenges
+The Challenges experiment employed stratified randomization on track specialization (UX Design vs.
+Front-end Development), age category (over/under 30), and residence type (Warsaw, other large city, or small town).
+To account for this design, we re-estimate treatment effects using stratum-weighted estimators that properly reflect the randomization procedure.
+Table 14 reports these stratum-reweighted estimates.
+For difference-in-means, we estimate treatment effects within each stratum and aggregate using stratum-size weights.
+For Cox proportional hazard models, we allow separate baseline hazards.
+Point estimates are similar to those in Table 2, with modestly smaller standard errors reflecting the efficiency gains from accounting for stratification.
+Table 14: Average Treatment Effects on Employment Outcomes (Stratum-Reweighted)
+Challenges (Stratum-Reweighted)
+Difference-in-means Cox PH
+New job Tech job Non-tech job Tech job
+ATE 0.080 0.120** -0.040 0.111**
+(0.051) (0.048) (0.042) (0.049)
+ATE / Control mean (%) 16.0 45.2** -17.2 41.8**
+(10.2) (18.2) (17.8) (18.4)
+Treated 166 166 166 2,656
+Control 218 218 218 3,488
+Notes: Average treatment effects on employment outcomes using stratum-reweighted estimators.
+Columns 1–3 report stratum-weighted difference-in-means estimates; column 4 reports stratified Cox proportional hazard estimates.
+Standard errors in parentheses.
+*** $p<0.01$, ** $p<0.05$, * $p<0.1$.
+Appendix I Average Treatment Effect Estimates Using Survival Forest
+This section estimates the average treatment effects of Mentoring and Challenges, on employment outcomes using causal survival forests.
+We use experimental samples only.
+The outcome of interest is the time until participants secured a new job, measured over a 13-month follow-up period.
+This approach estimates the difference in survival probabilities between treated and control groups at a specified time horizon.
+For each program, we fit a causal survival forest with 2,000 trees and compute the average treatment effect at 12 months post-intervention.
+Results are presented in Table 15.
+Table 15: Average Treatment Effects of Mentoring and Challenges at 12 Months.
+Mentoring Challenges
+   ATE -5.180 -2.895
+   SE (0.291) (0.244)
+   N Treatment 151 183
+   N Control 147 225
+Note: Estimates of the average treatment effect of Mentoring and Challenges at 12 months.
+Estimates obtained using causal survival forests.
+Estimates based on experimental samples with no control variables.
+The results indicate that the Mentoring program reduce the time it takes to find a job by 5.18 months,, while the Challenges program reduced this time by 2.9 months.
+Both estimates are statistically significant.
+Figure 15 displays the average treatment effects of the Mentoring and Challenges programs on employment outcomes over time, measured from 3 to 13 months after program start.
+Both programs show negative treatment effects that grow larger in magnitude over time, indicating that participants in the treatment groups experience accelerated job finding compared to control groups.
+At 3 months, both programs show treatment effects near zero, suggesting minimal immediate impact.
+However, as time progresses, the effects become increasingly negative, meaning treated participants find jobs faster (have lower survival times).
+By month 13, the Mentoring program shows a treatment effect of approximately -6 months, while the Challenges program shows an effect of about -3.5 months.
+Figure 15: Treatment Effect Estimates from Survival Forest.
+Note: Note: Figure shows average treatment effects on employment probability at different time horizons following program completion.
+Treatment effects estimated using causal survival forests with 2,000 trees.
+Shaded regions represent 95% confidence intervals.
+Horizontal line at zero indicates no treatment effect.
+Appendix J Policy estimation algorithms
 Algorithm 1 describes the estimation and evaluation of optimal policy.
@@
 Considering the dataset of all applicants to both programs, randomly split the dataset into train and test with equal probability at the applicant level, denote the resulting sets as $I^{train}=\left\{1,..,\mathcal{I}^{train}\right\}$ and $I^{test}=\left\{1,..,\mathcal{I}^{test}\right\}$.
-2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
+2. In the training dataset, estimate outcome model $\mu_{train}$ and propensity model $e_{train}$.
 3.
-In the test set, construct predicted treatment effects using predictions from model $\mu_{train}$.
-Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p)_{train}-\hat{Y}(O)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p)_{train}$ is the predicted outcome under program $p$ constructed using $\mu_{train}$ and $\hat{Y}(O)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months.
+In the test set, construct predicted treatment effects using predictions from models $\mu_{train}$ and $e_{train}$.
+Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p=P)_{train}-\hat{Y}(p=OutofDareIT)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p=P)_{train}$ is the predicted outcome under program $p\in$ (Mentoring, Challenges, Out of Dare IT) and $\hat{Y}(p=OutofDareIT)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months.
 Compute mean treatment effects per user as: $\tau^{Op,train}_{i}=1/T\times\sum_{t=1}^{15}\tau^{Op,train}_{i,t}$,
@@
 We solve the following constrained optimization problem:
-$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{i=p}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
+$\max_{z_{ip}}\sum_{i\in I^{test}}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i\in I^{test}}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{p=1}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
 The first constraint ensures that the capacity constraints are not violated.
@@
 Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$.
-See Appendix LABEL:cross_fit_appendix for details of the cross-fitting procedure.
 Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
@@
 return $(\mathcal{A}_{X,Q}^{*},\hat{V}_{X,Q}^{*},\sigma_{X,Q}^{*},\hat{Y}_{1,k}(a^{*}),...,\hat{Y}_{\mathcal{I}^{test},k}(a^{*}))$
-H.1 Assignment Models
+J.1 Assignment Models
 We pool data from Mentoring experimental participants (N=298), Challenges experimental participants (N=408), non-selected Mentoring applicants (N=468), and late Challenges applicants (N=160), yielding 1,334 total applicants.
@@
 Using the training set, we estimate predicted treatment effects for each program relative to non-participation.
-We fit two causal survival forests (athey2019generalized; cui2023estimating): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
+We fit two causal survival forests (Athey et al., 2019; Cui et al., 2023): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
 Each causal survival forest is estimated with the following parameters: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months.
 Covariates include: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization.
-These settings follow standard recommendations in the generalized random forests (grf) package (tibshirani2024grf).
+These settings follow standard recommendations in the generalized random forests (grf) package (Tibshirani et al., 2024).
 The causal survival forest estimates conditional average treatment effects on the restricted mean survival time (RMST) scale—the average time until technology employment over our observation horizon.
@@
 nothing) for each test-set applicant.
-Appendix I Evaluation Models
+Appendix K Evaluation Models
 While assignment models determine which applicants receive which program under various policies, evaluation models estimate the causal effects of these assignments.
@@
 Model specification.
-For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (athey2019generalized; cui2023estimating).
+For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (Athey et al., 2019; Cui et al., 2023).
 Within each cross-fitting fold (detailed below), we fit two causal survival forests: one comparing Mentoring participants to applicants who did not participate in any program, and another comparing Challenges participants to non-participants.
@@
 The covariate set is identical to that used in assignment models: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization.
-Model fitting follows the grf package implementation (tibshirani2024grf).
+Model fitting follows the grf package implementation (Tibshirani et al., 2024).
 The causal survival forest produces augmented inverse propensity weighted (AIPW) scores $\psi_{i}^{M}$ and $\psi_{i}^{C}$ for each test-set individual, representing individual-level treatment effect estimates from Mentoring and Challenges relative to non-participation, respectively.
-The forest automatically constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a survival forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates.
+The forest constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a regression probability forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates.
 These components are combined via the AIPW formula, yielding estimates that remain consistent if either the propensity or outcome model is correctly specified.
@@
 After completing all folds in repetition $m$, each individual has predicted treatment effects $\psi_{i}^{M,m}$ and $\psi_{i}^{C,m}$ estimated on data that did not include that individual.
-We then aggregate across the 100 repetitions: following chernozhukov2017double, we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$.
+We then aggregate across the 100 repetitions: we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$.
 Variance estimates are similarly aggregated by taking the median across repetitions.
+We follow the method described in Chernozhukov et al.
+(2017).
 Policy value estimation.
 Given an allocation policy $\pi$ that assigns each test-set individual to Mentoring, Challenges, or neither, we estimate the policy value as:
-$\hat{V}(\pi)=\frac{1}{n_{M}}\sum_{i:\pi(i)=M}\psi_{i}^{M}+\frac{1}{n_{C}}\sum_{i:\pi(i)=C}\psi_{i}^{C}$
+$\hat{V}(\pi)=\frac{1}{n_{M}+n_{C}}\left(\sum_{i:p_{i}=M}\psi_{i}^{M}+\sum_{i:p_{i}=C}\psi_{i}^{C}\right),$
 where $n_{M}$ and $n_{C}$ are the number of individuals assigned to Mentoring and Challenges, respectively.
@@
 The validity of our policy comparisons depends on the evaluation models accurately estimating treatment effects, not on the assignment models perfectly ranking applicants.
-I.1 Model Validation: Recovering Experimental Effects
+K.1 Model Validation: Recovering Experimental Effects
 To validate our evaluation model specification, we assess whether cross-fitted AIPW scores correctly recover experimental average treatment effects.
@@
 Deviations statistically indistinguishable from zero in both programs.
-For the Mentoring program (Panel a), AIPW predictions center tightly around the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03).
-The distribution is approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE.
+For the Mentoring program (Panel a), AIPW predictions center slightly above the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03).
+The difference between RCT and the mean deviation is statistically insignificant.
+The distribution appears approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE.
 The modest spread in the distribution reflects both the treatment effect heterogeneity among participants and sampling variation in the predictions.
-For the Challenges program (Panel b), AIPW predictions show similarly strong performance, with a mean deviation of $0.06$ months (SE = 0.04).
-While the distribution exhibits slightly more dispersion than Mentoring—consistent with Challenges’ larger experimental sample and potentially greater heterogeneity—the mean deviation remains statistically indistinguishable from zero.
-The slight positive mean suggests modest underestimation, but the bias is economically negligible (less than 2% of the experimental ATE).
+For the Challenges program (Panel b), AIPW predictions show similarly good performance, with a mean deviation of $0.06$ months (SE = 0.04).
+While the distribution exhibits slightly more dispersion than Mentoring, the mean deviation remains statistically indistinguishable from the RCT mean.
 The cross-fitted causal survival forests accurately recover experimental treatment effects, demonstrating that our evaluation approach is well-calibrated and unbiased within the experimental sample.
 This gives us confidence that the models correctly estimate treatment effects where treatment is randomized.
+Appendix L Alternative Policy Evaluation Using Cross-Fitting
+Our main off-policy evaluation in Section 6 employs a sample-splitting approach in which policy estimation and evaluation occur in separate datasets.
+This appendix presents an alternative approach using repeated cross-fitting following Fava (2025).
+The procedure proceeds as follows.
+For each repetition $m\in\{1,\ldots,M\}$ with $M=100$, we randomly partition the full sample of applicants into $K=5$ folds, stratified by experimental group, using a different random seed.
+Within each repetition, we use $K$-fold cross-fitting: for each fold $k\in\{1,\ldots,K\}$, we train causal survival forests on the remaining K1K-1 folds, obtain out-of-sample predicted treatment effects $\hat{\tau}^{M}_{i}$ and $\hat{\tau}^{C}_{i}$ for observations in fold $k$, solve the constrained optimization problem in Equation 3 to determine optimal assignments, and compute AIPW scores under these assignments.
+After completing all folds within a repetition, we aggregate out-of-fold AIPW scores to obtain policy value estimates and standard errors computed as the sample standard deviation of individual scores divided by $\sqrt{n}$.
+Following Fava (2025) and Chernozhukov et al.
+(2018), we aggregate across the $M$ repetitions by taking the median of point estimates and the median of confidence interval bounds.
+This approach requires a uniform convergence condition on the estimated treatment effects:
+$\sup_{x\in\mathcal{X}}\left|\hat{\tau}^{p}(x)-\tau^{p}(x)\right|=o_{p}(1)\quad\text{for }p\in\{M,C\},$ (4)
+where $\hat{\tau}^{p}(x)$ denotes the estimated conditional average treatment effect for program $p$ at covariate value $x$ and $\tau^{p}(x)$ is the true effect.
+This assumption is stronger than the pointwise consistency required by the sample-splitting approach in our main specification but enables valid inference when the same data are used for both policy learning and evaluation.
+Table 16: Treatment Effects Under Targeted Allocation Policies: Repeated Cross-Fitting
+Mentoring Challenges Overall
+Panel A: Benchmark
+Random allocation -2.22 (0.05) -2.54 (0.02) -2.38 (0.03)
+Panel B: Targeted allocation policies
+15% / 15% / 70% -5.26 (0.02) -3.94 (0.03) -4.60 (0.04)
+33% / 33% / 34% -4.40 (0.05) -3.07 (0.04) -3.73 (0.04)
+50% / 50% / 0% -3.41 (0.06) -2.55 (0.04) -2.98 (0.04)
+Notes: Treatment effects measured as change in restricted mean survival time (months); negative values indicate faster job finding.
+Capacity allocations specify the share of applicants assigned to Mentoring/Challenges/Out.
+Random allocation assigns each applicant to either program with equal probability.
+Results are median estimates across 100 repetitions of cross-fitting with different random fold assignments.
+Standard errors in parentheses.
+Table 16 presents median treatment effects across the 100 repetitions.
+Point estimates are similar to those in our main specification (Table 6): the overall treatment effect under the most selective policy (15/15/70) is 4.60-4.60 months compared to 4.34-4.34 months in the main analysis, while effects under expanded capacity (50/50/0) are 2.98-2.98 versus 2.86-2.86 months.
+The qualitative conclusions are unchanged: targeted allocation substantially outperforms random assignment, with gains from targeting diminishing as capacity expands.
+The principal difference between specifications lies in the estimated standard errors.
+Cross-fitting yields smaller standard errors, particularly for selective policies where few applicants are assigned to each program.
+At 15% capacity, for example, the standard error on the Mentoring treatment effect is 0.02 compared to 0.09 in the main specification.
+This difference reflects efficiency gains from using the full sample for both estimation and evaluation rather than reserving half for each task.
+Additionally, the median aggregation across 100 repetitions improves reproducibility: different researchers using the same data but different random seeds will obtain identical results.
+These smaller standard errors, however, rely on the uniform convergence assumption holding in finite samples; the sample-splitting approach provides more conservative inference under weaker conditions.
+The concordance of point estimates across both approaches strengthens confidence in our conclusions regarding the value of targeted program assignment.